The Cerebus Companion

Jerry Sweet

All rights reserved. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989 by Jerry N. Sweet.

Artwork copyright \begin{matrix}c \\ ◯ \end{matrix} 1989 by Dave Sim.

The Cerebus Companion is published with the permission of Dave Sim, President of Aardvark-Vanaheim, and author of Cerebus. However, the contents of the Cerebus Companion have not been reviewed for accuracy by Mr. Sim. Any errors or omissions are the responsibility of the author of the Companion. Although the Companion contains reference material that the author believes to be accurate, the Companion should not necessarily be considered as the authoritative source of information concerning Cerebus. After all, there is always Cerebus itself for that. Enjoy!

“Cerebus wants you to run that by him again...slowly.”

— Cerebus the Aardvark

Contents

1 Introduction 1

2 The World of Cerebus 1

3 Cerebus the Aardvark 4

4 Cerebus the Comic Book 6 4.1 The Artists and Publishers . . . . . . . . . . . . . . . . . . . . 6 4.2 The Stories and the Art . . . . . . . . . . . . . . . . . . . . . 7 4.3 Cerebus On the Bookshelf . . . . . . . . . . . . . . . . . . . . 8 4.4 Cerebus in the Archives . . . . . . . . . . . . . . . . . . . . . 9

4.4.1 Swords of Cerebus . . . . . . . . . . . . . . . . . . . . 9 4.4.2 AV in 3-D . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.4.3 Cerebus Jam . . . . . . . . . . . . . . . . . . . . . . . 11 4.4.4 Cross-Overs . . . . . . . . . . . . . . . . . . . . . . . . 12 4.4.5 Epic Illustrated . . . . . . . . . . . . . . . . . . . . . . 13 4.4.6 Benefits, Covers, and Interviews . . . . . . . . . . . . . 14 4.4.7 Portfolios . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.4.8 Miscellanea . . . . . . . . . . . . . . . . . . . . . . . . 16

5 Early Cerebus Stories 17 5.1 Recurring Characters . . . . . . . . . . . . . . . . . . . . . . . 17 5.2 Summaries of the Early Cerebus Stories . . . . . . . . . . . . . 22

6 High Society 38 6.1 Cast of High Society . . . . . . . . . . . . . . . . . . . . . . . 40 6.2 Summary of High Society . . . . . . . . . . . . . . . . . . . . 45

7 Exodus 58

8 Church & State 59 8.1 Cast of Church & State . . . . . . . . . . . . . . . . . . . . . . 61 8.2 Summary of Church & State . . . . . . . . . . . . . . . . . . . 72

8.2.1 Volume 1 . . . . . . . . . . . . . . . . . . . . . . . . . 72 ii CONTENTS

8.2.2 Volume 2 . . . . . . . . . . . . . . . . . . . . . . . . . 83 9 Square One 93 10 Jaka’s Story 94

10.1 Cast of Jaka’s Story . . . . . . . . . . . . . . . . . . . . . . . 94 10.2 Summary of Jaka’s Story . . . . . . . . . . . . . . . . . . . . . 95

10.2.1 In the Present . . . . . . . . . . . . . . . . . . . . . . . 95 10.2.2 Jaka’s Childhood . . . . . . . . . . . . . . . . . . . . . 98

11 Further Reading 99 12 Aardvark-Vanaheim Ordering Information 100 13 Acknowledgements 101 14 The Cerebus Glossary 103 15 Index 117

1 Introduction

How can an aardvark become a prime minister or a pope? How can a story have humor and adventure, depth and allegory, pathos and romance—with an aardvark as its central character? It all depends on the kind of good storytelling that Cerebus portions out just about every month. Cerebus is a black and white comic book series featuring a nasty barbarian aardvark named Cerebus, a quasi-medieval fantasy setting, and humor of many sorts: dark, light, anachronistic, slapstick, subtle, visual, parodic.

The Cerebus Companion is a reference work that may be used to supplement readings of Cerebus. The Cerebus Companion is a bibliography, an episode guide, and a glossary. Longtime readers of Cerebus may find here new insights into what they have already seen; new readers may find themselves better able to follow some of the references to previous occurrences. As always in outline works, there is no way that reading the Cerebus Companion can take the place of reading Cerebus. Instead, curious readers may use the Companion to determine whether Cerebus might be to their liking, or to help pinpoint obscure throw-away references. Readers are encouraged to see for themselves how well Cerebus is put together. After all, Cerebus isn’t “just” a story, it’s artwork, and that’s at least half the fun.

2 The World of Cerebus

Somewhere on Earth, on a European-like continent called Estarcion, it is the early 15th century. Estarcion is not quite Europe, for the geography of the continent, as shown in the map on the next page, must be forgiven at least a missing Italian peninsula, as well as a few other minor subtractions or additions. The time is not the 15th century anno Domini, for the predominant religions do not include any known today, and there are curiously anachronistic spins to the vocabularies used by the inhabitants, as well as to their dress. This time and place is a relative high point among the cycles of civilizations that have come and gone on Earth, but, if one is to believe one of the semi-omniscient characters, it is six thousand years in the past.

2 2 THE WORLD OF CEREBUS

Aardvarkian Age map

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

The map does not show definite political boundaries, but in the western part of the continent are Boreala and the Sepran Empire; in the east are the Hsifan Khanate and the Tcapmin Kingdoms. Separating east and west are the Red Marches. In the south is a region containing several city-states, including Palnu, New Sepra, and Iest, in which much of Cerebus takes place. The weather in Iest, one of the more northerly city-states, is wretched; it drizzles, rains, and snows often, contributing to a wonderfully gloomy mood. The weather gets warmer as one goes south towards the city-state of Eshnosopur.

Regional accents abound, rendered as Italian (like Chico Marx’s) for New Seprans, U.S. Southwestern (like Yosemite Sam’s) for those from Onliu, Cockney accents for those living in Lower Iest, and, ostensibly, broad upper-class English (like Alistair Cooke’s) for those from Greater Iest. French is used sometimes to indicate higher education or sophistication, but does not yet have a specific regional origin.

The technology of Estarcion includes elaborate architecture, indoor water pumps, seagoing ships, printing, paper money, crossbows, and cannons (although cannons are very new, and unknown to most). There is magic and there are sorcerers. Appearances are made by an elf, by a giant, animated, stone statue, and by other extranormal phenomena.

The danger of using either magic or technology with ill-defined parameters is that readers might begin to wonder why, for example, a magic amulet cannot be used simply to wish away a bad situation. Fortunately, in Cerebus, readers need not seriously fear a deus ex machina at every climax. New technological story elements either can be taken for granted or are have enough attention directed to them that they are not serious surprises when they are used. It cannot be said that all new magical story elements are introduced as carefully, but magic is intrinsically unpredictable, and magic is traditionally part of the attraction of fantasy stories. Although the nature of magic in Cerebus is not well defined, it tends to be limited to illusion or to small morphological changes. Of course, if one can swallow a sentient aardvark, the rest is easy. There is “big magic” in Cerebus, but it tends to be foreshadowed heavily. So far, there is no need for readers to worry about death-wish-at-a-distance; proximate sword arms are still needed for that sort of thing.

3 Cerebus the Aardvark

The title character, Cerebus, is an anthropomorphic aardvark—a “funny animal,” standing upright, and possessing hands. He is three feet tall, with round, flat feet, and stubby legs. He is entirely covered with short, grey fur that smells terrible when it is wet; this fact is used humorously sometimes. He has a thick forelock of hair in between pointed, expressive ears. He has a long, flat-ended snout with nostrils like single quote marks in front. His travelling clothes are an open vest and a necklace of three or more featureless medallions.

Born ca. 1385, Cerebus is twenty-eight at the time of the first Cerebus novel, High Society. He has an education in reading, writing, and to a small extent, in sorcerous and meditative arts. He speaks with a Northern accent, manifested as a tendency to refer to himself by his own name instead of using the first person, and to say “aye” or “nay” instead of “yes” or “no.” Cerebus likes ale, apricot brandy, and whiskey, the more the better. He has a nasty, hair-trigger temper and brusque manners. He glowers almost constantly. He can beat up most anyone, and often does. Cerebus has been seen to use his tail and snout as an auxiliary fists when his arms are otherwise occupied. Women think that Cerebus is cute, and they like to stroke his fur (when it’s dry). That Cerebus is an aardvark does not prevent him from having carnal knowledge of human women, although he had evidently been slipped a drug when first introduced to the practice, and presumably would not have acquired the taste without it.

Cerebus is not a friendly creature. Worse, Cerebus’s ignorance of certain social graces and his lack of sensitivity have a tendency to destroy the few friendships he is offered. Cerebus is therefore perhaps the least sympathetic fantasy character since Thomas Covenant. Nevertheless, Cerebus is not without sympathetic qualities. As one character puts it, Cerebus has the demeanor of an orphaned puppy and has a vast need for a mother substitute. It is easy to feel sorry for Cerebus, somehow, even after he has done something cruel and is suffering the consequences.

Cerebus is a sword-wielding northern barbarian, a mercenary, a thief, and a somewhat unwilling politician. Cerebus has been known to espouse repub-Young Cerebus

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Cerebus the Aardvark

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

licanism (albeit somewhat cynically), but when push comes to shove, he is somewhere to the right of Attila the Hun. Cerebus has just enough charisma, ambition, and blundering luck to find himself from time to time in positions of leadership over men. Unfortunately, he is rudderless enough and callow enough to be manipulated thoroughly by those he perceives as being better informed than he is. He is reasonably competent in matters military, but is out of his depth when he must mix in politics or diplomacy. Cerebus is lucky enough to have wonderful opportunities drop into his lap, but he is usually not quite smart enough, sophisticated enough, or patient enough to maintain a permanent grasp on the good things those opportunities bring him.

With very few exceptions, all the other major characters in Cerebus are shown as human, and the story takes place in human societies. What does it mean that Cerebus is an aardvark in a world of humans? It is easy to forget this fact in spite of his appearance, but there are frequent reminders—if not from those who see him merely as a short, or deformed, or young human, then from those who rudely address him simply as “aardvark.” Attempts to attach some great literary significance to Mr. The Aardvark don’t work very well, not because Cerebus is not told in a traditional literary medium, but because Cerebus is not purely a metaphor; Cerebus’s personality is too strong for that, and Sim has bigger allegories to fry than, for example, exploring the “animal nature of man.” However, time and again it is hinted that aardvarks are important, perhaps divine, creatures in the historical cycles of Estarcion. The early stories do not invite one to take these hints seriously, but the hints become difficult to ignore later on.

4 Cerebus the Comic Book

4.1 The Artists and Publishers

Cerebus is written and drawn by Dave Sim, with the assistance of a fellow whose nom-de-nib is Gerhard. Cerebus was introduced in 1977. Dave Sim did all, or at least much, of the work until about 1984, when he started sharing the workload and the credit with Gerhard. In general, Gerhard now does inking and the densely detailed background art; Sim handles everything else.

Denise (Deni) Loubert was Dave Sim’s helper and publisher of Cerebus until the end of 1984; she was also Sim’s wife until 1983. Ms. Loubert now runs Renegade Press in Long Beach, California. Dave Sim’s then brother-in-law, Michael Loubert, occasionally provided sociological and geographical advice in the early days of Cerebus. He invented many, if not most, of the place names, and designed the “Aardvarkian Age” map.

Ms. Loubert claimed responsibility for the name of the series, which resulted from a misspelling of Cerberus, the three-headed dog guarding the gate of Hades in Greek mythology. The name stuck. According to a commentary by Dave Sim, Cerebus the aardvark came about as a logo for Ms. Loubert’s fanzine publishing enterprise, which was given the name Aardvark-Vanaheim from suggestions by Ms. Loubert’s brother and sister.

4.2 The Stories and the Art

Unlike most of the later part of the series, the first few Cerebus issues are selfcontained stories. The stories spoof the swords-and-sorcery fantasy genre, making for a light comedy/fantasy/adventure mix, not unlike Fritz Leiber’s stories about Fafhrd and the Grey Mouser (Swords Against Wizardry, The Swords of Lankhmar, and others). The contrasts both in storytelling and in artwork from the early Cerebus stories to the later ones are interesting; the series has acquired a great deal more professional polish as it has progressed. This is not to say that the early stories are poorly done; the stories are certainly amusing, and are essential in setting up events and in introducing characters that appear in the later, longer works.

At the time of this writing, the Cerebus series includes two complete novels, High Society and Church & State, and one novel in progress, Jaka’s Story. The Companion summarizes the first two novels and part of the third.

Apart from the artwork, much of the appeal of Cerebus is in the storytelling and the dialogue. Sim takes his time to tell his stories in a level of detail uncommon in comics. Information about bureaucratic setups and socioeconomic conditions is slipped into conversations in bits and pieces. The true meanings of conversations may not be revealed immediately. References to places, persons, and historical occurrences give a sense of a much larger whole than the localities on which individual episodes may focus.

The approach may tend to appeal more to an audience that likes French cinema than one that seeks constant action in its choice of graphic media, although even the later Cerebus stories are not without liberal doses of adventure. There have been cries from some quarters to revive the simpler flavor of the earlier barbarian adventures, but it is probably safe to say that most of the current readers of Cerebus like things the way they are now. The stories mix comedy, adventure, pathos, romance (of a sort), and political and religious satire. Cerebus is not exactly a soap opera, at least not in the same sense that Omaha the Cat Dancer might be considered a soap opera, but Cerebus does have a soap opera’s sense of an ongoing story, an ongoing life.

The artwork of Cerebus is largely uncluttered, unlike some black and white comics that make the eye search hard for the important details. Uncluttered it may be, but the backgrounds, including buildings and room furnishings, are carefully detailed, more often than not indicating that some thought goes into them. In both foregrounds and backgrounds, attention is actually paid to light sources, shadowing, diffusion, reflection, and the optical effects of water on glass. Repetition is used interestingly: a page of six panels may feature a character entirely motionless throughout while action or conversation takes place off-camera, or may just show the character’s thoughts. Little arabesques adorn the spaces around the panels when it is appropriate to contribute to a sense of an elegant environment.

4.3 Cerebus On the Bookshelf

Cerebus is presently available in several forms. There is a bi-weekly edition of Cerebus that started December 2, 1988, reprinting the series from the beginning and including one-page (non-Cerebus) backup features. Then there is the regular monthly edition, which is up to issue 120 at the time of this writing. Dave Sim has steadfastly maintained that Cerebus is a 300-issue limited series.

Original back issues of the comic book, of course, are available from shops specializing in comic books. There are many reasons to buy the back issues, not the least of which is that Cerebus, during Deni Loubert’s incumbency, often featured interesting non-Cerebus backup features running under the banner A Unique Story. And the occasional Cerebus backup features do appear. The “Note from the President” and the “Aardvark Comment” sections of each issue may also interest those willing to sift for behind-the-scenes tidbits about the creative process and occasionally cogent analyses from fans.

For those primarily interested in obtaining the story in large chunks, there are the Cerebus “phone book” collections available directly from the publisher, Aardvark-Vanaheim:

Title  Issues Covered
Cerebus^{∗}  1–25
High Society  26–50
Church & State volume 1 52–80
Church & State volume 2 81–111

∗ The first Cerebus volume’s working title was Cerebus the Barbarian; it may be used to disambiguate references to the phone book and the series.

For ordering information, see section 12 on page 100.

Note that issue 51, Exodus, is not included in the phone books. The reason for this is discussed in section 7 on page 58.

4.4 Cerebus in the Archives

This section summarizes all “non-mainstream” or archival Cerebus material that has come to the attention of the author.

4.4.1 Swords of Cerebus

Cerebus issues #1 to #25 were repackaged in six smaller collections called Swords of Cerebus, printed between 1981 and 1984. Some of the prints of Swords volume 6 accidentally omitted the story from issue #25 (This Woman, This Thing), and were accompanied by supplements. In addition to the main stories originally printed in the Cerebus comics, each Swords book contains commentaries by Dave Sim and additional Cerebus features that never appeared in the monthly Cerebus comics. The commentaries are being reprinted in the bi-weekly reissues of Cerebus. The commentaries discuss Sim’s inspirations, drawing techniques, and sometimes his personal life.

Those outtake stories that do fit squarely into the main continuity of Cerebus are summarized in section 5 at appropriate points. The outtake stories that stand alone and that do not quite fit into the main continuity of Cerebus are summarized below.

• Diamondback. [Swords vol. 1] Cerebus plays Diamondback in a tavern, defeating all opponents, including Leopold, the famous Gambling Priest. The cards, the hands, and Cerebus’s favorite variation of the game, Karet, are explained.

• Demonhorn. [Swords vol. 2] This story was originally published in Nucleus #1. It is told in verse. M’isly, a god of some sort, plays a small jest on Cerebus from on high. As Cerebus tries to cross the mountain Demonhorn on his way to Iest, M’isly teleports Cerebus a few hundred yards backward each time he approaches the summit. Cerebus stops and starts to dig. When M’isly’s curious eye appears nearby to get a closer look at what Cerebus is doing, Cerebus stabs the eye and goes on his way without further incident.

• The Morning After. [Swords vol. 2] Cerebus has a pounding hangover and does all sorts of interesting things to get rid of it, including paying to have his head pressed in a blacksmith’s vise. Then he sits in front of the tavern waiting for it to open again.

• Cerebus Dreams. [Swords vol. 5] This is a little story done in cooperation with Barry Windsor-Smith. Cerebus, sleeping away a morning in a hayloft, dreams that gems are falling all around him. A rooster pecks at some of the gems and swallows a few. A cone of light follows Cerebus as he tries to gather the gems. A jester appears, playing on a fiddle, and knocks Cerebus over. The jester takes bows in the cone of light, accepting applause from an unseen audience. Cerebus awakens, but lies on his side and scratches as a rooster wanders about nearby.

• A Night on the Town. [Swords vol. 6] This story is told primarily in pantomime, using icon balloons. Cerebus drunkenly staggers through the snow with a bottle of alcohol. He trips over a corpse lying face-down in the snow. Not realizing that the man is dead, he tries to revive the man with a drink from the bottle, then stands the corpse upright and walks away with it. A constable stops Cerebus and tries to tell him that his drinking buddy is dead. Cerebus refuses to believe it, and asks the corpse to prove it to the constable. The constable threatens Cerebus with gaol, and Cerebus flees, leaving the constable entangled with the corpse. The frustrated constable finds himself unable to catch Cerebus, and walks away, leaving the corpse in the snow. Cerebus emerges from his hiding place and proceeds to waltz his dead drinking buddy down the street again.

4.4.2 AV in 3-D

In December, 1984, Aardvark-Vanaheim published AV in 3-D. It contains six short features, all rendered with 3-D effects by Ray Zone. Work by various artists then publishing through Aardvark-Vanaheim appears in the comic. Undisturbed copies of the comic include two pairs of 3-D glasses. The fivepage Cerebus story, “Cerebus Dreams II,” details one of Cerebus’s dreams, foreshadowing an event in Church & State called the Final Ascension.

The non-Cerebus stories in the comic feature Valentino’s normalman, Bob Burden’s Flaming Carrot, Max Collins’s and Terry Beatty’s Ms. Tree, Bill Messner-Loebs’s Journey, and Arn Saba’s Neil the Horse.

4.4.3 Cerebus Jam

Aardvark-Vanaheim published another non-mainstream work in April, 1985, a black and white comic called Cerebus Jam containing four cooperative short works:

• “The Defense of Fort Columbia” with Scott and Bo Hampton. This has cartoon versions of Scott and Bo arguing about the best path through a forest as Cerebus looks on with a sort of disgusted patience.

• “The First Invention of Armour: 1404” with Murphy Anderson. Cerebus and a fellow named Bear (later to appear in Church & State), between mercenary assignments, meet Hagbard Spleen, inventor of armor. Hagbard, armored from head to foot, defends the virtues of his armor to Cerebus and Bear, but falls into a river and apparently drowns.

• “Squinteye the Sailor” with Terry Austin. Nine-year-old Cerebus meets a grizzled, peglegged sailor who smokes a corncob pipe, carries a Jeep-like creature on his shoulder, and is unbelievably strong.

• “Cerebus versus the Spirit” with Will Eisner. Cerebus dukes it out with the Spirit.

A Cerebus Jam #2 has yet to arrive (and may not), but there was a Jam story, done in cooperation with Colleen Doran (A Distant Soil), that appeared in Cerebus #91:

• “The Applicant.” An extremely annoyed young woman slams into a tavern where Cerebus is having an otherwise calm evening enjoying his ale. She grabs Cerebus by his vest and roughs him up a bit, all the while disgustedly reciting the text of an ad for a housekeeper, in the interview for which she clearly was sexually harrassed. She also terrorizes the barkeep. Cerebus’s friend Bear enters, obliviously chucks her under the chin, and tries a pickup line, whereupon he gets kicked where it hurts the most.

4.4.4 Cross-Overs

Cerebus also appeared in several “cross-over” works, including, but not necessarily limited to, the following:

• Teenage Mutant Ninja Turtles #8; reprinted in color in the third TMNJ graphic novel from First Publishing. TMNJ is by Kevin Eastman and Peter Laird. This story features the Turtles travelling back in time, losing their means to get back to their own time, and falling in with Cerebus to recover their time-travelling device from an evil wizard.

• Flaming Carrot Comics #18, by Bob Burden, published by Dark Horse Comics in 1988. Black and white. Cerebus appears in exactly one panel as a cut-in commentator. The Flaming Carrot makes a reciprocal appearance in Cerebus #104.

• Miami Mice #4, by Mark Bodē and Larry Todd, published by Rip Off Press in 1987. Black and white. Cerebus makes an ungracious cameo appearance in a bar visited by the Mice and also by the Teenage Mutant Ninja Turtles.

A few other comics have been known to make oblique Cerebus references. That is only fair, since Cerebus frequently contains in-jokes about other comics.

An internal cross-over story, “Elfguest,” appeared in Cerebus issue #52. The story is really more of an affectionate parody than a cross-over. Two elf characters similar to Cutter and Skywise from WaRP Graphics’ Elfquest series mistake Cerebus for an elf and invite him to visit their dimension. Their visit is short, however, and the elves return through the dimensional warp. Cerebus, who was sleeping off a good drunk in an alley, resolves never again to mix ale and apricot brandy.

4.4.5 Epic Illustrated

Four short works featuring “Young Cerebus,” colored by Gerhard, were published in Epic Illustrated, Marvel’s now defunct science fiction and fantasy anthology magazine. All of the Young Cerebus stories are done in pantomime. The stories are:

• “His First Fifth;” October, 1984 (vol. 1, #26). Nine pages plus one page of introduction. Although the story is done primarily in pantomime, icon balloons (instead of speech balloons) are used. Six-year-old Cerebus, mopping the floor of a tavern, chances on a bottle of liquor left on the table top of a booth. He swipes the bottle and goes to the cellar to try it out. Caught by a young boy who threatens to tell the owner, Cerebus instead convinces the boy to try some of the liquor. The boy gets a bit tipsy. Cerebus asks about the owner of the bottle, a large man with an eyepatch. As the boy offers some derogatory opinions about the man, the two of them discover him standing over them, listening. The man picks up the boy and stuffs him into the furnace; he then turns his attention to Cerebus, who fearfully denies any connection with the bottle and rushes off to finish the mopping.

• “A Friendly Reminder;” February, 1985 (vol. 1, #28). Four pages. This work features eighteen-year-old Cerebus as a tax collector who smashes his way into a hovel and truncheons a man. It is not a story so much as a very violent gag: “Next year file early.”

• “Selling Insurance;” June, 1985 (vol. 1, #30). Two pages. Eight-year-old Cerebus races down a cobblestone street, pulling a large bag by a rope. The bag bounces and tumbles heavily behind him. He slams the bag down in front a small group of children and with a menacing glower, holds out his hat. The bag opens to reveal a bruised child, who, along with the rest of his peers, throws copper bits into Cerebus’s hat.

• “The Girl Next Door;” June, 1985 (vol. 1, #30). Three pages. Fiveyear-old Cerebus, hiding, shyly watches a blond girl his age as she goes out to pick berries. He gather flowers for her, then thinks better of it; he beans her with a rock and runs away, completely in love.

The last Cerebus-related material published in Epic Illustrated was “The Ages of Cerebus;” October, 1985 (vol. 1, #32). Three pages. Five portfolio portraits of Cerebus at various ages, with brief commentary.

4.4.6 Benefits, Covers, and Interviews

Another “Young Cerebus” vignette, “Breaking Up Is Hard To Do,” appeared in Anything Goes! #3, published by the Comics Journal in October, 1986.

It is three pages long, with prose and three color illustrations. The color is Tom Luth’s work; he also colors Sergio Aragonés’s Groo the Wander. Like “A Friendly Reminder,” BUIHTD is short and ultra-violent (blessedly off-camera), having to do with the departure of a girlfriend named Michelle. The cover of the comic, featuring a disturbed-looking Cerebus, is by Neal Adams and Cory Adams.

Cerebus also graced the cover of True North, which, like Anything Goes, is a legal defense benefit comic. A benefit poster of True North’s cover is available as of August, 1989. A legal benefit portfolio called “F.O.O.G.” (Friends of Old Steve Gerber) contained one Cerebus drawing by Dave Sim. Amazing Heroes Preview Special #4, published by Fantagraphics Books, has a cover by Dave Sim and Sergio Aragonés featuring Cerebus and Groo.

There have been a number of interviews of Dave Sim in various magazines and in at least one documentary (“Masters of Comic Book Art”). The Comics Journal issues #82 and #83 contain a two-part interview with Dave Sim, along with some cover artwork. The Comics Journal #130 features another interview with Dave Sim. Back issues of the Comics Journal may be obtained from its publisher, Fantagraphics Books. The Comic Informer #8 also contained an interview and some cover artwork by Dave Sim.

4.4.7 Portfolios

Sim and Gerhard have made individual pages and portfolio prints available in limited quantities in the past. In 1983, there was a forty-five plate color portfolio called The Animated Cerebus containing three stories and an introduction. The portfolio was produced in contemplation of an animated Cerebus film. The stories are told entirely in pantomime. They are:

• “A Well-Equipped Bar” (16 plates). Cerebus sits in a tavern, nursing a mug of ale. He gazes contemplatively at an apparently abandoned brandy glass sitting at the corner of the bar. He looks around, and finding no one watching, he snatches it and drinks the contents. Leaving the empty glass on the bartop, he sits back down and suffers agony trying to keep the liquid down. A man with something in his cheek approaches the glass and spits a stream into it. Cerebus realizes that he has just swallowed the contents of a makeshift spitoon.

• “Add One Mummified Bat” (12 plates). A young Cerebus watches Magus Doran conjuring over a roiling cauldron. Doran, distracted as he consults his book of magic, reaches for a mummified bat. He accidentally grabs Cerebus by the ears and drops Cerebus into the cauldron. A brown, demonic-looking smoke-blob with red eyes mushrooms out of the cauldron and spits out Cerebus. The blob looks angrily at Doran, then toasts the wizard with a stream of flame. Cerebus tiptoes away as the blob sinks back into the cauldron, leaving Doran with a frizzled face.

• “His First Sword” (17 plates). A large, bald, blacksmith hammers a sword into its final shape, then tosses it to Cerebus. Cerebus makes experimental dueling motions with a happy look, then decides to test more serious combat grips. Hefting the sword backward, he accidentally stabs the blacksmith. The blacksmith falls by his anvil. Cerebus is frozen with mortification. He looks around contritely, then drops a few coins into the apparently dead blacksmith’s outstretched fingers, then walks away, examining his sword with a satisfied expression.

In addition to the “unlimited edition” of The Animated Cerebus, there was a 100-copy limited edition with one extra plate.

Another portfolio printing was made available in 1985 called “The First Fifth” (not to be confused with “His First Fifth”), consisting of six prints limited to one-hundred copies each, thirty of which were color. There was also a 1,750-copy print called “The Six Deadly Sins.”

4.4.8 Miscellanea

There once were several reputedly well done episodes of a Cerebus radio show done by the Shockwave public radio group in Minneapolis. Tapes of the show exist, but are not commercially available. An adaptation of Cerebus #6 (The Secret!) by Kara Dalkey was performed by the same group at Minneapolis’s Minnecon in 1983.

There have been T-shirts, holographic medallions, stuffed toys, posters, Diamondback card decks, and other Cerebus items advertised in the pages of Cerebus.

There was once a Cerebus fan club run by Aardvark-Vanaheim. The newsletters featured Cerebus drawings on the covers.

5 Early Cerebus Stories

A few of the summaries given in this section are of stories found only in Swords of Cerebus, and are included in this section where they fit into the continuity.

The issue number to which a particular story summary corresponds is given in the margin.

5.1 Recurring Characters

The first recurring characters other than Cerebus are introduced starting with the third issue of Cerebus. What follows is not a comprehensive cast list, because the stories are so self-contained. However, it does list those major characters who have important appearances later in Cerebus.

Several of the characters parody other comic book and fantasy characters. Cerebus himself initially parodied Robert E. Howard’s Conan the Barbarian (as drawn by Barry Smith and written by Roy Thomas). Almost all of the parodic characters, rather than existing for the sole purpose of parody, also have distinct personalities. They do not simply make cameo appearances and vanish; they have importance to the stories in their own rights.

The characters are, in order of appearance:

Henrot, a magician. His powers are reputed to be mighty, but they are evidently not adequate to avenge insults to the honor of his daughter, Red Sophia. As a result, Henrot had occasion once to hire Cerebus for some wet work. In spite of this penchant for employing thugs, Henrot interacts with the mundane world somewhat detachedly or perhaps bemusedly, in a sort of amused, knowing way. Henrot dresses in the stereotypical wizard fashions: peaked hat, flowing robes, moon and stars decorations, full beard.

Henrot’s name is an anagram of Thorne, as in Frank Thorne, one of the artists on Red Sonja. Henrot also resembles Thorne, who cultivates a wizardly image.

Red Sophia, a sword-wielding, brawling barbarienne who parodies Red Sonja (“She-Devil with a Sword”). Because Red Sophia always wears a plate-mail bikini, she is the subject of several jokes about scars. Her personality seems to alternate between Bimbo and Competent Swordmistress. Sophia is a bit fond of B&D. She is sexually attracted to Cerebus, the only man or creature, or perhaps one of the few, to defeat her in single combat. Cerebus, being an aardvark, exhibits no interest in her considerable, but human, physical charms the first two times she appears in Cerebus’s life.

Bran Mak Mufin, leader of a loin cloth wearing, militant, undergrounddwelling race called the Pigts. Bran is a thin, rugged fellow with long blond hair. He wears two necklaces, presumably as badges of office; one necklace strings together what appear to be animal teeth, and the other necklace holds oval, multicolored medallions. Bran believes truly in the divinity of Cerebus because of some ancient Pigt legends and an ancient aardvark statue hidden in the Pigt warrens.

The Pigts live beneath the Red Marches, an area rich in quicksand and black flies, but which is nevertheless a popular area of attempted conquest by adjoining states. In turn, the Pigts live to topple adjoining empires.

Bran is a parody of a Robert E. Howard character, Bran Mak Morn, King of the Picts of Caledonia.^{1}

Jaka, a blond, beautiful, tavern dancer. She takes a bit of a fancy to Cerebus and his adventurous life.

1The Picts were not purely figments of Howard’s imagination; they were a people inhabiting ancient Britain, and they gave the Roman conquerors a lot of trouble.

Elrod, King of Melvinbone, who parodies Michael Moorcock’s Elric of Melnibone, as drawn by Barry Smith (later Barry Windsor-Smith) in Conan #14 and #15. Elrod, in one episode, also satirizes Deadman. Elrod is a tall albino with a conical hat. He talks like Foghorn Leghorn. Elrod thinks Cerebus is a kid in a bunny suit. Elrod isn’t overly bright. He usually shows up, loudly announces himself, and in the process causes trouble for Cerebus precisely when Cerebus is trying to be inconspicuous.

The original Cockroach

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Captain Cockroach and Bunky

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

The Cockroach, a less than sane Lower Feldan merchant who bounces about the rooftops of Beduin, victimizing anyone whom he thinks are the Red Claw cultists who killed his parents. Cerebus calls the Cockroach “Cootie” or “Bug.” The Cockroach wears tights, a hood with antennae, and a crenulated masque. The Cockroach makes a hissing noise when he’s annoyed or about to go into action. He has a fixation about helping orphans.

The Cockroach makes appearances throughout Cerebus, mutating into different parodic characters as the series progresses. More than once, the Cockroach satirizes Batman, but also takes off on Superman, Captain America, Wolverine, the Moon Knight, and the Hulk, to name the most recognizeable targets.

Necross the Mad, also known as Necross the hahaha Mad, an evil sorcerer. He wears a hood with a skull symbol painted on one side. The name says everything else.

Lord Julius, essentially Groucho Marx^{2} in the part of Rufus T. Firefly. Lord Julius is the Grandlord of Palnu. His technique for keeping himself in power is to use his considerable wit to confuse everyone.

Suenteus Po, the 182-year-old founder of Illusionism. He spends a great deal of time meditating in the Seventh Sphere of consciousness with the aid of certain mind-altering substances. For someone who spends so much time out of the “real” world, he has a good understanding of secret societies (he should, as the founder of one), and a fair grasp of realpolitik.

Adam Weisshaupt, a Lower Feldan aristrocratic schemer, fairly ingenious, and quick to see opportunities. He is also a major snob. See section 11 on page 99 for some background on the origin of the Weisshaupt character.

Professor Charles X. Claremont, a mild-mannered, but evil, sorcerer who conjures up Apocalypse Beasts. He appears initially disguised as 5.1 Recurring Characters 21

2Groucho was born Julius Henry Marx.

The Merchant Cockroach

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Elrod

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Madame DuFort, a strict, elderly schoolmistress for aristocratic young women.

Professor Claremont’s name is taken from Chris Claremont, who writes the X-Men series, and from one of the X-Men characters, Dr. Charles Xavier, the wheelchair-bound mentor of the X-Men.

Theresa, one of Madame DuFort’s students. She is bold and assertive, although trained in the social graces. She speaks with a French accent.

Prof. Charles X. Claremont

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

5.2 Summaries of the Early Cerebus Stories

1 The Flame Jewel. This story introduces Cerebus bouncing into town on horseback and, inside the local tavern, gaining immediate respect for his ability to defend himself. There he makes an arrangement with two other thieves to steal a jewel from a wizard. After doing battle with the wizard’s illusionary forces, Cerebus slays the wizard and gives the jewel to the other thieves, accepting a bag of gold in payment. The jewel, without the wizard’s magic, turns into a walnut. Cerebus rides off, leaving the thieves behind.

Captive in Boreala. This issue parodies a Roy Thomas/Barry Smith Conan 2 story from Savage Tales #1 and Conan #16. Cerebus is taken prisoner by muscular Borealan marauders. After proving his ability and ruthlessness in single combat, he is taken into the ranks. Three days later, marching through snow, the group encounters a stronger force of sorcerously animated warriors with glowing white eyes. During combat, Cerebus slides over the edge of a cliff and falls through a snow drift into an underground chamber. Cerebus identifies the chamber from his memory of an ancient book, and goes in search of the valuables he knows will be there. When he finds and carries away the glowing Eye of Terim, a succubus, seen in its true, horrific form, swirls out of the Eye and makes an unsuccessful attempt to steal Cerebus’s soul.^{3} When the succubus flees, Cerebus finds the Eye of Terim to have turned into a tarnished iron sphere. Making his way out of the chamber, he finds skeletal remains of the ensorceled warriors and Borealan corpses. Cerebus, broke and hungry, heads toward the coastline.

Song of Red Sophia. The title of this story is borrowed from the story “Song 3 of Red Sonja,” which appeared in Conan #24. Henrot, a magician, hires Cerebus for a bit of assassination and torture. Cerebus dislikes such work, but he needs the money to escape Borealan press gangs. Loaded with supplies and accompanied by Red Sophia, Henrot’s daughter, he goes in search of Henrot’s intended victim, a man who has besmirched Red Sophia’s honor. On the way, Sophia challenges Cerebus to a swordfight, which Cerebus quickly wins without harming Sophia. Sophia makes a surprising announcement:

sophia: I have sworn a vow that any man who defeats me in battle shall possess me...I am yours to command! Let me do your bidding, oh short grey master...my body is yours! cerebus: Do my bidding, eh? Cerebus is most pleased! You can carry the supplies!

3This is one of two explicit attempts to locate Cerebus’s soul. Either he doesn’t have one or it is well protected.

That night, there is a nasty incident in a tavern that begins when a lout challenges Cerebus for Sophia’s attentions. Sophia exacerbates the situation by exaggerating Cerebus’s prowess, but in the end proves herself capable of defending her own honor. The next day, Sophia pesters Cerebus by offering him favors he doesn’t want. Finally, Sophia and Cerebus encounter Tanes Feras, the intended victim. Feras turns out to be a somewhat effete admirer of Sophia, and as Cerebus prepares to torture Feras with hot iron, Feras faints. Cerebus decides instead that the best torture for Feras, while also serving to rid himself of Sophia, is to make Sophia marry Feras. Cerebus returns to Henrot, assures Henrot that Feras is suffering, and takes his leave with his payment.

4 Death’s Dark Tread. Cerebus comes across a gem in a back alley in Serrea. Unbeknownst to Cerebus, Death is watching, and desires the evil power of the gem. Cerebus, finishing his apricot brandy, takes the gem and almost immediately runs into a monster, which he slays. Death, who evidently must work through third parties, looks about for someone who can wrest the gem from Cerebus and from whom the gem in turn may be wrested. Later that day, Cerebus, posing as a merchant and trying to stay inconspicuous, is set upon by Elrod, a loud-mouthed fellow who postures belligerently and won’t leave Cerebus alone. Elrod picks a fight with some soldiers and gets himself and Cerebus thrown into a dungeon, but Cerebus escapes. Cerebus decides that the gem is a source of bad luck, and throws it into a well, much to the despair of Death.

5 The Idol. Cerebus, having taken employment as a mercenary, walks alone on patrol in the Red Marches. He seeks shelter from the constant rain in a tree hollow. He is found by some Pigt warriors out hunting. The Pigts are astonished at Cerebus’s appearance, and invite him to return to warm shelter with them. On the way, a wretched odor is the subject of much comment until the source of the odor is discovered to be Cerebus’s wet fur. In the underground warrens of the Pigts, Cerebus is taken to the Pigt leader, Bran Mak Mufin. Bran, also taken aback by Cerebus’s appearance, nevertheless tries to make Cerebus comfortable, while exhibiting a keen interest in Cerebus’s species and age. Bran offers Cerebus a Pigtian traditional drink of fermented goat’s blood. He tells Cerebus that the Pigts, although small in number (try fifty warriors), live to topple the surrounding empires. Cerebus decides that Bran is a few bricks shy of a full load, but accepts Bran’s offer of shelter. Later that night, Cerebus sneaks into a Pigt rally at which he sees an aardvark-like statue that is identified as the image of an ancient Pigt god-king. Bran incites the Pigts to war in honor of the restored god-king, Cerebus. After the rally, when the Pigts have gone, Cerebus destroys the statue, and makes his way back aboveground, heading for Iest.

The Secret! In an alley in Iest, a dying man speaks his last words to Cerebus. 6 This is witnessed at a distance by E’lass, a sleazoid, and his companion, Turg the Unduly-Tall, a dunce, who were hot on the trail of the dying man. Turg, attempting to grab Cerebus, is instead beaten up. E’lass, apologizing for Turg, offers to buy Cerebus a drink. Once the three are seated in a tavern, E’lass drugs Cerebus’s ale to make him suggestible. When E’lass experimentally points out Jaka, the tall, blond, tavern dancer, Cerebus is love-struck. Cerebus approaches Jaka and offers to kill a yak for her supper. Jaka thinks Cerebus is very cute, and kisses him, then retires to her dressing room. E’lass finds that he has lost control of Cerebus, and while Cerebus is distracted by a bar brawl, E’lass threatens Jaka to make her get Cerebus’s secret, the location of treasure hidden in a temple. Jaka invites Cerebus to her room.

cerebus: I’d love to lick apricot brandy out of your navel...

jaka: Look, Cerebus, Jaka has to tell you something.

cerebus: If Cerebus had a navel would you lick apricot brandy out of it?

Jaka tells Cerebus that E’lass threatened to kill her. Cerebus immediately leaves to beat up E’lass. When he returns to Jaka, Cerebus offers to take Jaka with him to find the treasure and share his adventures. Jaka accepts. Cerebus goes to gather supplies, but then the drug wears off. When Jaka appears, dressed for the journey, Cerebus does not recognize her, and rejects her. Jaka sadly gives him a gift of apricot brandy and promises to wait until Cerebus remembers her again.

Black Sun Rising!^{4} Cerebus rides through snow-covered mountain passes to 7 the Temple of the Black Sun, where awaits a treasure guarded by a death worshipping cult. Elrod appears, wearing snow shoes, and loudly announces himself ready for the Festival of the Black Sun, which Elrod believes is a party, until Cerebus points out a rotting head on a stake nearby. Cerebus and a reluctant Elrod invade the fortress. While Cerebus searches, Elrod wanders off and runs into a man who is wearing an aardvark-like suit. The man in the suit, Mit, is a short, put-upon priest, who has made himself the suit so that he will be mistaken for the Nameless God of the Black Sun. Elrod and Mit promptly raise a ruckus that gets themselves and Cerebus into trouble.

4The cover art for this issue was by Frank Thorne, the only Cerebus comic book cover to date not done by Dave Sim and/or Gerhard.

In trying to escape from the aroused priests of the Black Sun Cult, Cerebus falls into a pit and is grabbed by an enormous spider-beast, a semisupernatural creature. Like the succubus in Captive in Boreala, the spiderbeast probes for Cerebus’s soul. Cerebus is saved when the incantations of the priests cause an earthquake and explosion. The earthquake drops the spider-beast into the pit, while the explosion destroys the fortress-temple and hurls Cerebus out and into the snow. Elrod and Mit, unscathed, head south. Cerebus, picking himself up out of the snow, unhappily wanders away.

8 Day of the Earth-Pig! Cerebus, delirious from an infected wound inflicted by the Black Sun cult’s spider-beast, is captured by Conniptin soldiers after a fierce struggle. The Conniptin prince, a wastrel addicted to a nasally ingested narcotic, wants Cerebus made into a bathrobe. Cerebus is saved by the commander. The commander and his captain contrive a plot to make Cerebus a hero: they send Cerebus out to kill a Hsifan border patrol, which Cerebus does, although he is still in the grip of febrile hallucinations. When Cerebus regains consciousness, the commander explains that the troops won’t fight for the prince, but they would fight for a heroic leader, a potential that the commander sees in Cerebus. After being threatened with disembowelment, Cerebus agrees to take the job. Cerebus then addresses the troops in a fine autocratic rant. That night he escapes, but thinks better of it and returns.

9 Swords Against Imesh. Cerebus leads the Conniptin soldiers to the walls of the Hsifan city of Imesh, where Cerebus once lived. Leaving the Conniptins encamped outside, Cerebus scales the walls of the city to assess the enemy troop strength. He finds the entire city population working hurriedly to build a huge altar in the center of the city. Climbing down, he is enticed into a concealed doorway, where he finds himself trapped behind iron bars, on the other side of which is K’Cor, king of Imesh. King K’Cor is a huge fellow with a beard and large, menacing eyes. The king explains that he has enslaved Imesh to build a symbol to scare away invading Venusians. The slaves work for Buz, a nutritive compound blended with an opium derivative, which makes them pliable, but diligent workers. Cerebus, finally getting a word in edgewise, explains that he has come to conquer Imesh. K’Cor offers Cerebus a wager: if Cerebus can defeat K’Cor’s champion, Cerebus can have Imesh, but otherwise, the Conniptins will join the slaves.

Cerebus accepts the wager, and promptly finds himself doing battle with glowing energy globes, which he encountered once before in his youth. Getting past the globes with difficulty, he finds himself confronted with K’Cor’s champion, Lord Koghem, of whom Cerebus makes fast work. Cerebus, near exhaustion, then finds himself faced with an armored K’Cor, but after a short battle, K’Cor halts the fight. He tells Cerebus that he poisoned all the wells in the area surrounding Imesh, and that the Conniptins are all dead by now. Cerebus is left alone, weakened, and without food or shelter.

Merchant of Unshib! Cerebus pushes southward through a blizzard, starving 10 and alone. He is surprised by Red Sophia, who offers him food, shelter, and an opportunity to join her in stealing a magic talisman, the Black Blossom Lotus, in transport nearby from the city of Unshib. Sophia tells Cerebus that her husband, Feras, had died trying to defeat her in combat, Sophia’s pre-consummation requirement. In her tent, Sophia makes sexual advances to Cerebus while explaining how the merchant carrying the Lotus came by it. Cerebus is not aroused. They are interrupted by Sophia’s other partner in crime, Meirgen, who had been out reconnoitering the target. Later, out in the snow, the trio is confronted with a crossbow-wielding advance scout of the caravan. Cerebus offers the scout a share of the loot, but the scout demands to hear the plan before agreeing to join. Cerebus, working up a plan on the spot, tells him that they will destroy the supports of a nearby bridge as the guards cross over, and pick the Lotus out of the rubble. As the group awaits the caravan below the bridge, Cerebus proposes that he should double back and keep an eye on the caravan. He asks the scout for the loan of the scout’s fur to use as a tent. Wearing the scout’s fur, Cerebus runs into the caravan, yelling about Khaiven sorcerers who changed him, and sends the excited guards out looking for the thieves by the bridge. Once alone with the merchant, Cerebus holds him up at swordpoint and walks away with the Lotus.

11 The Merchant Cockroach. In the city of Beduin in Lower Felda, Cerebus sells the Black Blossom Lotus to a sorcerer turned merchant. The merchant throws the Lotus into the river, explaining that he considers mystic objects bad for business when they fall into the wrong hands. As Cerebus leaves with his payment, the merchant takes a nap, but when Cerebus hears the merchant raving, Cerebus returns to investigate. Cerebus finds that the merchant, when asleep, becomes the Cockroach, who wears an antennaed costume and runs amok, killing random strangers whom the Cockroach believes killed his parents. The Cockroach relieves his victims of their gold and throws the gold down a shaft to be forgotten. Cerebus manages to bamboozle the suggestible Cockroach/merchant into moving some of the gold for him in a hay-wagon, but it is all Cerebus can do to keep the merchant from regaining lucidity. As they approach the city guards, the merchant comes to himself and demands that the guards seize Cerebus. Cerebus unbridles a horse and escapes, but not before fingering the merchant as the Cockroach.

12 Beduin by Night. Cerebus thinks about how to recover the Cockroach’s remaining gold now that the Cockroach is in gaol. Cerebus rows a small boat back to Beduin and docks at the Cockroach’s former riverside residence. When Cerebus breaks into the shaft where the gold was once kept, he finds only empty space and a tunnel leading to another room. Climbing through, he finds the Cockroach, recently escaped from prison, standing amidst the gold. The Cockroach threateningly demands an explanation. Cerebus invents the story that he is an orphan whose parents, like the Cockroach’s, were also killed by Red Claw cultists. This instantly wins the Cockroach’s addled sympathies. Taking advantage of this, Cerebus then suggests that the gold should be moved before a villain named Elrod of Melvinbone comes to steal it, and the Cockroach agrees. Placing a quantity of gold aboard the rowboat, the Cockroach attempts to leave Cerebus behind, but Cerebus catches up to him. The swift current forces the boat alongside a landing under a bridge where, to Cerebus’s dismay, Elrod just happens to be fishing. The Cockroach is at a loss to deal with Elrod’s effusive greeting. After a bit of a Laurel and Hardy routine, Elrod accidentally knocks the Cockroach into the river. As Cerebus helps the Cockroach back onto the landing, Elrod comments that the rowboat has a slow leak. Just then, the city guards appear, looking for the fugitive Cockroach. Cerebus makes short work of the guards, then angrily knocks out the Cockroach. Leaving a very confused Elrod on the landing, Cerebus rows away. Cerebus exults over the gold he has in the boat, but the bottom of the boat falls out, and the gold is lost. Grasping a piece of the boat, Cerebus is carried out beyond Beduin’s walls by the current.

Magicking. This is an outtake story from Swords of Cerebus vol. 4. Cerebus, adrift on the Feld River, falls asleep. He awakens to find himself falling in starry space. He slams into a floating square cobblestone floor and recovers on the floor’s surface to meet a white-haired, mustachioed magician. The magician tries empirically to determine what an aardvark is. After magically stretching Cerebus out of shape, he finds the proper reference in “Ye Book of Beasties” and apologetically puts Cerebus back the way he was. The magician tells Cerebus of a cataclysm that left only a few magicians in the world. The remaining magicians are matched off with each other in pairs of good and evil. The magician, who claims to be among the good, also notes that aardvarks have an attribute whereby they concentrate and greatly magnify spells. The magician uses Cerebus to conjure up a weapon with which to fight evil magicians, then returns Cerebus to his plank on the Feld. The magician, however, is left with a huge 20th century military tank about which he knows nothing.

Black Magiking. Cerebus awakens on the bank of the Feld River in Lower 13 Felda. He is confronted by two farmers who believe him to be an evil creature of sorcery, perhaps because his fur is the same color as a mildew blighting the crops. Somehow managing to capture Cerebus, they drag him before a Tarimite priest. Cerebus threatens them with magical metamorphosis. The priest, uncertain whether Cerebus is bluffing, decides to leave Cerebus to the mercy of Necross the Mad, an evil sorcerer. Knocking Cerebus out, the priest and a farmer named Despuess tote Cerebus to Necross’s castle. Despuess takes Cerebus into the castle while the priest waits outside. As Cerebus regains consciousness, Necross’s booming Voice From On High angrily demands to know who has violated his castle. Necross annihilates Despuess with a magical bolt, and the priest runs for help, leaving Cerebus to deal with Necross. Necross appears in a fiery cloud, then falls to coughing from the smoke. After recovering, Necross proudly shows off a huge stone statue named Thrunk, which Necross has been considering bringing to life. Necross tells Cerebus of his plans to destroy the universe. Just then, the priest returns with a mob, and Necross goes to confront them. To Cerebus’s and Necross’s surprise, Necross goes down with an arrow in the chest. A moment later, Thrunk comes alive, Necross’s spirit having entered the thing. Cerebus escapes as Thrunk gets busy crushing the mob and the priest.

Silverspoon. This story is an outtake from Swords of Cerebus vol. 4. The story was originally printed as a series of one-pagers in the Buyer’s Guide for Comic Fandom, now called CBG, in 1980. Silverspoon is a parody of Prince Valiant, with one spoiled young Lord Silverspoon, son of Lord Julius of Palnu, bearing resemblance to the aforementioned prince. The story opens in the Lower Feldan port city of Denieau, where Cerebus takes passage on a trading vessel bound for Palnu. On board the ship, he grumpily makes the acquaintance of Lord Silverspoon, who accidentally knocks Cerebus overboard. Later, Silverspoon “slays” a fake sea serpent that the crew has placed in the water to give Silverspoon a sense of adventure. They do this to keep Silverspoon from gaining a sense of adventure by making them instead engage Onliu pirates. Nevertheless, the ship is eventually attacked, set afire, and sunk by Onliu pirates, leaving Cerebus to swim alone to a desert island. There, he meets again with Silverspoon. Cerebus agrees to gather food for the both of them in exchange for a promise of two hundred gold pieces. However, Cerebus and Silverspoon are soon captured by natives, who tie them between stakes and leave them to die of exposure. The two are rescued by Lord Julius and his troops. Silverspoon tries to convince Julius that Cerebus kidnapped him, but Julius refuses to execute Cerebus summarily. At that, Silverspoon throws a tantrum. Julius has one of his men knock out Silverspoon with a rock. Julius takes Cerebus with him to Palnu, and promises him a reward.

14 The Walls of Palnu. Cerebus is given a hero’s welcome in the wealthy citystate of Palnu. Lord Julius puts Cerebus in charge of the security forces, giving him the official title of Kitchen Staff Supervisor. Obfuscated job titles are part of Julius’s mode of operation, which is to keep everyone in the dark about how the bureaucracy works. Cerebus finds out that the “elected” ruler and Grandlord of Palnu is actually he who has purchased the largest constituency, in this case, Julius. Cerebus faces boredom until one day someone tries to poison Lord Julius. When Cerebus investigates, he finds and breaks up a meeting of revolutionaries. Cerebus pursues one of the revolutionaries through the rain onto a ledge outside Julius’s bedroom window. As they hold discourse and then fall to fighting, Julius leans out of the window and offers Groucho-type wisecracks. Cerebus wins the fight by snout-punching the revolutionary off the ledge.

A Day in the Pits. Cerebus, accompanied by Lord Julius and Baskin, Julius’s 15 stenographer, takes the investigation of the hidden revolutionaries into the Pits of Palnu, a ruined ancient city lying somewhat intact beneath Palnu. There they are trapped by a hooded fellow who introduces himself as leader of the revolutionary group, the “Eye in the Pyramid.” The man introduces Cerebus to the Shadow Crawler, a huge snake, which attacks Cerebus. Using two torches, Cerebus manages to maneuver the beast into knocking itself out against a wall, but the hooded one escapes.

A Night at the Masque. Julius holds a costume party, which gives Cerebus 16 concern regarding proper security. Two thieves, E’lass and Turg (last seen in The Secret!), circulate among the guests in disguise. When disguised assassins close in on Lord Julius, Cerebus creates an amusing diversion. In the confusion, E’lass manages to steal a valuable jewel from one of the guests, and escapes through a hidden doorway temporarily revealed by the hooded leader of the Eye in the Pyramid. Cerebus guides Lord Julius out of the crowd through the only available exit, the same hidden doorway. E’lass, believing that Cerebus is in pursuit of him alone, runs ahead of Cerebus and Julius. Cerebus follows, in pursuit of the hooded one. On a narrow bridge in the Pits, Cerebus and Julius are confronted by the hooded one, who, confident of victory, reveals himself to be Julius’s social secretary. Cerebus tells Julius that he’s sick of the whole mess, and demands gold and a horse out of Palnu as his price for saving Julius’s life. Julius quickly accepts. Cerebus beans the revolutionary with a rock, sending the man over the edge. As Cerebus and Julius turn to leave, E’lass, believing himself nearly caught, throws away the jewel, but then finds himself ignored. Later, Cerebus takes his leave of Palnu with a horse and eight bags of gold. As Julius watches him go, a letter from one of Julius’s nieces is delivered. The letter says that the new Kitchen Staff Supervisor sounds like a friend the niece met a year ago, and that she would look forward to seeing him in a few days. The letter bears the signature of Jaka.

17 Champion. Three weeks later, Cerebus’s horse goes lame, leaving Cerebus to drag his own gold through the rain. He finds his offer to buy a farm hovel accepted. As Cerebus settles down to a warm fire, a fellow with a T’gitan accent (rendered as a German accent) comes to the door and requests to use the hovel for an army headquarters. Cerebus admits the man, whose name is Gudre, and his companion, a tall, silent man named Stromm. Gudre explains that the united T’gitan tribes under his leadership are invading Palnu, and the first target is the city of Fluroc. Cerebus listens to the invasion plan with interest, then with ambition. When Cerebus mentions that Fluroc is guarded by the formidable Commander Krull, Stromm shows signs of getting violent until he is calmed by Gudre. Gudre explains that Krull was responsible for cutting out Stromm’s tongue some years ago. The next day, Cerebus trails after Gudre as he goes to receive a report from a scout. The scout reports that Fluroc is undefended and the roads were deserted but for two priests leaving the city. Julius has moved all troops to the Onliu border.

Cerebus believes that the priests are a disguised Krull and an assistant, since these are the High Holy Days, during which priests are not allowed to leave their places of meditation. Furthermore, Cerebus believes Krull is leaving to bring back some troops to defend against the impending T’gitan invasion. Cerebus then goes to take Krull prisoner, which he succeeds in doing by tricking Krull’s travelling companion into hitting Krull with a rock. A week later, Lord Julius receives the report that Fluroc has fallen into the hands of the T’gitan barbarians.

18 Fluroc. As Cerebus searches private libraries in the looted city for information that would be useful in the invasion of Palnu, he makes the acquaintance of Graus, one of the savage T’gitan tribesmen. Meanwhile, in Palnu, Julius appoints one of his political enemies, Lord Gorce, to command an army against the T’gitans. While Julius hopes to be cheered by news of Gorce’s inevitable demise, Gorce sees the appointment as an opportunity eventually to take control of Palnu.

Back in Fluroc, a party of N’Gamin merchants, guarded by mercenaries, approaches the city walls. Cerebus considers a plan to trick them out of their gold. First he has the T’gitans disguise themselves as Palnans. Then he invites the merchants into the city, but to explain all the corpses left by the T’gitans, he warns the merchants of a mysterious and terrible plague. One fearless merchant enters the city with his guards. Cerebus, disguised as a merchant, greets them and invites the merchant to trade. Cerebus then gets the merchant to leave all his jewels and gold by convincing the merchant that he has contracted the non-existent plague, then selling him the one and only flask of “curative medicine.” Since the transaction was conducted in the T’capmin language, which the T’gitans do not understand, Cerebus later makes up a story about the con, designed to appeal to the less refined sensibilities of the T’gitan tribesmen.

She-Devil in the Shadows. Cerebus, accompanied by Graus, travels to Tog- 19 ith to trade loot for more mercenaries for the invasion of Palnu. Graus enjoys three hours having his fortune told by a black-haired, berobed woman named Perce. From his old acquaintance, the Duke of Agerden (who used to sell pornographic prayer books), Cerebus finds that the price of mercenaries has gone through the roof. The Duke instead offers him a chance to steal the twin of an odd, Maltese Falcon-like owl, in exchange for which pikemen and heavy cavalry might be obtained. The owl is in the possession of one Master Hortne, who has not been seen out of his basement room in three weeks; all who have gone down to investigate have failed to return. Cerebus descends the darkened stairs to confront a sword-wielding woman (a parody of Frank Thorne’s Ghita of Alizarr) whom Cerebus mistakes for Red Sophia. The woman names herself Geet-a and proceeds to give Cerebus a very serious sword duel. Cerebus barely makes it through Hortne’s door, where he finds the wizard Henrot (Hortne is an anagram of Henrot) in the midst of doing wizardly things, surrounded by floating snake-slug-fish creatures called gerbies.

Henrot gives Cerebus a cheery greeting. He explains that he magically cloned Geet-a from traces of Red Sophia and chained Geet-a near his door as an illtempered guard. When Henrot further explains that neither he nor Cerebus can get past Geet-a, Cerebus becomes furious, whereupon the gerbies nearly smother him. Henrot tells Cerebus that the gerbies feed on violent emotions, and has Cerebus relax to get the gerbies off. Cerebus gets an idea. First he has Henrot whip up an ersatz owl. Next, he opens the door and insults Geet-a to anger her, then has Henrot unleash the gerbies. When Geet-a at last lies buried beneath a pile of happy gerbies, Cerebus leaves. Cerebus then goes to collect Graus at the fortune teller’s place. The fortune teller gives Cerebus some drugged wine, rendering him unconscious, then, telling Graus that Cerebus will be staying with her for a while, sends Graus away.

20 Mind Game. Cerebus becomes aware of voices speaking to him. The voices belong to Perce and Mother Wenda, two inner circle Cirinists. They tell Cerebus that he is drugged into a near coma. Cerebus almost immediately finds that he can alternate between two states of consciousness, one in which he can converse with the Cirinists, and one in which he can converse separately with a fellow named Suenteus Po. Po, the 182-year-old founder of Illusionism, spends a lot of time meditating in the Seventh Sphere.^{5} However, each time Cerebus pushes into the Seventh Sphere, his body stops breathing, which makes the Cirinists somewhat nervous. The voices of the Cirinists discuss some sort of ongoing Tarot card reading indicating that Cerebus is a random, but important, disruptive factor.

Po offers to help Cerebus stay alive while in the hands of the Cirinists and goes away briefly to do some research on the matter. The Cirinists tell Cerebus that he is being initiated into the Sacred Order of the Divine Mother, and that he is to gain a revelation in the Seventh Sphere within an hour or forfeit his life. Upon returning, Po has Cerebus tell the Cirinists that “the rebirth is at hand.” This makes the Cirinists so happy that they make plans to keep Cerebus as a permanently drugged object of reverence. Cerebus, understandably annoyed, decides to play Po and the Cirinists against each other. He gives Po hints of a dangerous Cirinist conspiracy, then claims to the Cirinists that he has been possessed by Po, making demonic-sounding utterances. Gradually, Cerebus, working through Wenda and Po, manages to get Togith’s Cirinist and Illusionist cells agitated enough to start fighting each other in the streets of Togith. Cerebus hopes to make the Cirinists now with him leave to battle the Illusionists, so that he will be alone until the drug wears off. However, Wenda, the last to leave the room, gives Cerebus a sleeping potion, which puts him all the way out.

5The panels showing the Seventh Sphere are rendered as odd divisions of black and grey that can be assembled like a puzzle into a picture of Cerebus.

What Happened Between Issues Twenty and Twenty-One. This story is an outtake from Swords of Cerebus vol. 3. G’evan, a thief, steals into Wenda’s apartment after she has left, stuffs Cerebus’s unconscious body into a sack, throws the sack over his shoulder, and leaves. Once outside, G’evan is killed by an aristocratically dressed fellow named Mr. Strangis, who evidently had hired G’evan, but believed G’evan was making away with the aardvark. During the fray, Cerebus, still partly draped in the sack, staggers away. Cerebus finds his way into a tavern and starts drinking, ignorant of the fact that the drugs still in his system are activated by alcohol. He finds his way out of the rear exit and collapses on the street. He is picked up by two New Seprans who run a freak show,^{6} where Cerebus’s comatose form is placed in front of a sign saying “We don’t know what it is, but it’s alive.”

In Harpengate, Cerebus is recognized and purchased by B’ren Kael, a bureaucrat. Kael rushes him to Sir Gerrik, a tall fellow with a receding hairline, thick, upswept eyebrows, and a long, thin mustache. Gerrik has his manservant kill Kael. Cerebus is examined by a woman named Astoria. Sir Gerrik, considering how to keep ahold of Cerebus, worries about the drugs in Cerebus’s system wearing off, but Astoria tells him that it is unsafe to drug Cerebus again. Astoria suggests that she take Cerebus with her to Beduin, where Cerebus will have to stay until the ban on travel is lifted. She mentions that Gudre’s T’gitan troops will soon be decimated by Lord Gorce and that Cerebus will learn a valuable lesson thereby. As Astoria leaves by carriage, Sir Gerrik muses about how little he really knows of Astoria, in spite of all the years he has known her.

Captain Cockroach! Cerebus awakens on a snow-covered public bench in 21 Beduin. He worries about getting back to Fluroc in time to join the T’gitan assault on Palnu, then notices that the streets of Beduin are deserted. He hears the sound of someone knocking on a door and turns to see the Cockroach and Elrod, costumed as Captain Cockroach and his faithful companion, Bunky (parodies of Captain America and his sidekick Bucky), trying to sell United Feldwar States Government-In-Exile War Bonds. As Bunky delivers the sales pitch, Captain Cockroach hisses, “Buy bonds! Beat the Hsiffies!” while bending lamp posts into pretzel shapes. Bunky cheerfully tells Cerebus that the war bonds are part of a plan to wipe out the Mongol-like “Hsiffies,” inhabitants of the Hsifan Khanate.

6A bit of fun is poked at Berni Wrightson, the original artist on Swamp Thing and the artist for Freak Show, which appeared in Heavy Metal in 1982.

The Captain and Bunky offer to take Cerebus to see “the President,” Adam Weisshaupt. When Weisshaupt sees Cerebus, he thanks Cerebus for having tamed the Cockroach the previous spring (referring to events in Beduin by Night). Explaining his plans to control Lower Felda, Weisshaupt tells Cerebus that his anti-Hsifan war bonds scheme has put him at odds with the military tribunal governing Lower Felda, which has declared a twentyfour hour curfew to prevent purchases of the war bonds, and sealed the city to catch the Cockroach. Cerebus accepts a drink from Weisshaupt, which hits Cerebus unusually hard. Weisshaupt gives Cerebus forged papers to get Cerebus out of the city, but as the Cockroach and Elrod conduct Cerebus through the streets, they are nearly hit by a Hsifan Sopai assassin’s throwing weapon. Cerebus, drunk from the wine and lingering drugs in his system, finds it difficult to stay lucid long enough to deal with the assassin, and knocks himself out with a makeshift sling.

22 The Death of Elrod. Elrod is struck by one of the Sopai assassin’s weapons and falls. Just as the Hsifan is about to crush the Cockroach’s skull, Elrod’s spirit takes possession of the assassin. Elrod and the Cockroach carry Cerebus and Elrod’s old body back to Weisshaupt. When Cerebus regains consciousness in Weisshaupt’s office, he refuses to ask what has happened and tries to leave, explaining about the war in Palnu waiting for him. Weisshaupt tells him that not only did the war happen without him, but the T’gitans were completely wiped out by Lord Gorce’s troops several weeks ago; Cerebus was evidently kept drugged by the Cirinists longer than he had thought. Cerebus and Weisshaupt then decide on an elaborate scheme to get the ruling military tribunal to destroy itself by using Elrod’s ability to control the bodies of other persons. The plan fails in action, however, when Elrod finds himself stuck back in his old body, which comes to life again. Cerebus and Weisshaupt are forced to run, but Cerebus leaves Weisshaupt behind.

23 The Beguiling. This story takes off on a Clint Eastwood movie. Miles from Beduin, Cerebus, suffering from a nasty leg wound gotten in an accident, collapses in the snow. He is rescued by three teenage girls, Theresa, Katrina, and Janette. Their teacher, Madame DuFort, offers to put Cerebus up in their boarding school until his leg heals. Madame DuFort explains that she has only the three girls as students this year, and warns Cerebus not to “tamper” with the girls. Cerebus is awakened by a clandestine visit by Theresa, who offers him whiskey and tries to engage him in “adult” conversation. Her monologue is interrupted when a lone New Sepran captain comes to the door downstairs, asking for shelter for his troops. Cerebus leaves his bed, preparing the confront the soldier with only a water jug for a weapon. However, he finds it difficult to do more than observe discretely from the top of the stairs. When the soldier doesn’t take no for an answer, Katrina shoots the soldier with a crossbow. Cerebus deduces that a large number of New Sepran troops are going to pass by the school, and are possibly on the way to conquer Upper Felda. One of the girls comes up with a plan to post a quarantine notice on the farmhouse and to disguise Cerebus as a plague victim. This does not deter a visit by two more New Sepran soldiers, Tiberius and Luis. Tiberius recognizes Cerebus as someone with whom he onced served. Luis then advances with knife drawn, saying, “For wadda y’u did to my brudda I’m a swore i’m a get a y’u someaday, littla gray smarta guy...” Before Luis can finish explaining why he is seeking revenge, both Luis and Tiberius are shot by Katrina and Janette.

Swamp Sounds. Still recovering, Cerebus plays cards and converses with 24 Madame DuFort’s students. He ignores Theresa’s little flirtations. Soon after the girls are called away by Madame DuFort, Cerebus hears a strange sound that he fears is of sorcerous origin. Madame DuFort denies having heard anything. The next day, as Cerebus again plays cards with the girls, Cerebus asserts that he doesn’t believe that true heroism exists. Theresa offers as a counterexample a story that credits Katrina’s uncle, Lord Julius, with personally defeating the leader of the Eye in the Pyramid conspiracy. Cerebus is speechless. Theresa then offers another story that credits Elrod with defeating the Serrean city guards in heroic defense of a small child in a rabbit costume. Although Cerebus is astounded at these distorted versions of his own adventures with Julius and Elrod (as seen in Death’s Dark Tread and A Night at the Masque), he says nothing, and the girls are again called away. Cerebus falls asleep, but then is awakened by the mysterious sound, which is louder than ever. When he finally goes to investigate, he finds the girls conducting a sort of seance, in the center of which are concentric hoops of force. Madame DuFort, caught without her wig, re-introduces himself to Cerebus as Professor Charles X. Claremont, a sorcerer conducting experiments with the girls to raise a hulking Apocalypse Beast, which he calls Woman-Thing (a take-off on Man-Thing). Clarement is a stereotypical mad scientist type, bent on revenge against those who laughed at his research, but he seems otherwise benignly disposed toward Cerebus.

25 This Woman, This Thing. Claremont leaves the girls in a suspended trance state, which he claims to be able to maintain for five to ten years. The Apocalypse Beast carries Claremont and Cerebus away from the school in a sort of sedan. As they travel, Claremont tells Cerebus of the destructive history of Apocalypse Beasts in Estarcion. Claremont exults at the opportunity to use Woman-Thing for evil: anyone who fears the beast will burst into flame—and who would not fear, seeing the Apocalypse Beast? Arriving at a castle, Claremont anticipates destroying the owner and moving in. However, to Claremont’s dismay, the castle dweller is not only delighted and impressed to see Woman-Thing, but offers to buy her. Claremont refuses. Their host fetches some wine and some of his own paintings, which primarily feature the breasts of women. Oblivious to Claremont’s attempts to frighten him, the artist launches into a discussion of the paintings’s deeper meanings. One of the paintings catches Claremont’s eye: a picture of another Apocalypse Beast. Their host happily explains that the monster in the painting is Filbert, who came with the castle. All four go upstairs to see Filbert, the “Sump-Thing” (named for his previous owner, Lord Roth-Sump), who is inanimate until Claremont extends his Apocalypse Beast control spell. Sump-Thing (who strongly resembles Berni Wrightson’s Swamp Thing) awakens to see Woman-Thing, and before Claremont knows what is happening, he is crushed between the two love-struck beasts. The artist/collector, overwhelmed with guilt that Cerebus’s friend has just been killed, offers Cerebus money. Cerebus, shedding a crocodile tear, proceeds to milk that guilt for everything he can get.

6 High Society

High Society chronicles Cerebus’s firsthand introduction to Machiavellian or Byzantine politics—choose your adjective—as opposed to the basic slashand-conquer politics Cerebus has previously known. Cerebus, formerly the Palnan Kitchen Staff Supervisor (actually the head of Lord Julius’s security forces), finds himself the ranking diplomatic representative of Palnu in Iest. That Cerebus holds this position is evidently not what Lord Julius intended. Strangely, the ranking diplomatic representative of Palnu in Iest is formally chosen by election.

The city-state of Iest is run by an entrenched upper class, and by the Eastern Orthodox Tarmite Church, which wields an Inquisition. Iest is divided into many districts, but primarily into Greater Iest, where the seats of power are located, and into surrounding low-lying areas. Iest’s economy, like the economies of its neighboring city-states, is hostage to high-interest loans provided by Lord Julius. Lord Julius therefore holds considerable influence in Iest. Making the political situation slightly twistier is the fact that Julius is in turn greatly in debt to the Sepran Empire, the seat of the Western Church of Tarim. The Eastern and Western Churches are at odds with each other.

A dissident intellectual faction, calling itself the Anarcho-Romantic movement, also wields some influence in Iest. Anarcho-Romantics dress foppishly, wear perfume, and write manifestos demanding construction of more rock gardens and reflecting pools. They yearn for an “intellectual revolution.”

Unexplained, nearly incomprehensible, and conflicting conspiracies are fomented by a number of secret societies. Mention is made of Kevillists, Cirinists, and Illusionists, but their aims and beliefs are not fully fleshed out. Furthermore, one is not necessarily to believe everything one is told about these groups.

The aim of Cirinism was half-jokingly said by Suenteus Po to be “to wipe out fun in our lifetimes.” The Kevillists are said to be a fanatical Cirinist faction (perhaps analogous to the Red Guard during the Communist Chinese Cultural Revolution in the 1960s). There was a Kevillist Rebellion in Lower Felda circa 1408.

Upper Felda suffered a revolution that placed the Cirinists in power there. The Upper Feldan Cirinist government makes its neighboring states very nervous, and therefore has to fend off the occasional border skirmish. Cirinists circulate throughout Estarcion, undermining the Tarimite Church wherever possible.

The Illusionists appear to be harmless, mostly concerned with achieving illumination (enlightenment) through meditation aided by psychoactive substances.

6.1 Cast of High Society

The main characters appearing in High Society are, in order of appearance or mention:

Cerebus, who is seen walking up the stone steps to a hotel in Iest, carrying a sack of loot from his encounter with the artist/collector in the previous issue (This Woman, This Thing, issue #25).

the Regency Desk Clerk, a snobbish fellow.

Holland M. Hadden, a purveyor of gold-plated street lamps.

Ned Greely, another merchant, a friend of the Prime Minister.

Skorz, a sewage engineer with an interesting idea of syntax (“Famous the Aardvark” is how he refers to Cerebus).

The Cockroach, aka the Merely Magnificent Moon Roach,^{7} aka Artemis Strong, aka Blake Jockley,^{8} aka Kevitch,^{9} aka Sergeant Preston Roach. The Cockroach’s civilian identity is that of a nominally sane cab driver, Blake Jockley. The Cockroach is a muscular fellow, a bodyguard and an assassin whose gullibility, mental instability, and malleability are used for various political and propagandistic purposes by those able to gain his trust. The Cockroach still believes himself to be some kind of superhero, but has not settled on a single persona. He lives a rich fantasy life, and often narrates his fantasies to himself as though he were reading a book. To borrow Dave Sim’s use of an old phrase, the Cockroach is a legend in his own mind. He also holds internal conversations with the Kevitch personality. The Moon Roach has developed a new fixation: economic injustice.

7The Moon Roach is a take-off on The Moon Knight.

8Blake Jockley’s name is derived from one of the Moon Knight’s identities, Jake Lockley, a taxi driver.

9The name of Bill Sienkiewicz, the first artist on Moon Knight, is pronounced Sin-Kev’-itch.

The Moon Roach

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Elf

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Captain Hill-Hodges, of waterfront special security.

Fleagle and Drew McGrew, two Onliu miscreants, initially would-be aardvark kidnappers, later Cerebus’s bodyguards. They both have long hair and drooping moustaches. Drew, the blond brother, is somehow irresistable to women. The Moon Roach’s Kevitch personality thinks he’s cute too.

Lord Julius, ruler of Palnu. The Groucho Marx impersonation gets better. Lord Julius has an unending supply of wise remarks, whereby he may or may not be hiding an expertise at convoluted intrigue.

Suenteus Po, still hanging around the higher spheres of consciousness and cultivating paranoia about Cirinist conspiracies.

Elf, aka the Regency Elf, a glowing female sprite in residence at the Regency Hotel. She appears to Cerebus from time to time to talk cheerfully about inconsequential things or to play wickets with him, but occasionally supplies Cerebus with useful information and gossip. No one cares to admit having seen Elf.

Blake Jockley and the brothers McGrew

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

The Prime Minister of Iest, a politician in trouble.

Astoria, Duchess of Parmoc, Julius’s ex-wife. She is a mysterious character, sometimes cold and manipulative, other times warm and manipulative. There are hints that she is an agent of some secret society. She is first seen by Cerebus as the Cockroach’s girlfriend. She has the Cockroach firmly under her thumb, and occasionally sends him off to assassinate various Iestan officials.

Astoria

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Lord Julius and Duke Leonardi

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Luzzo, ranking diplomatic representative of Eshnosopur in Iest and priest of the Church.

Elrod, recently adopted son of Julius, rival of Cerebus for the position of ranking diplomatic representative of Palnu in Iest, and apparently the most charismatic swordsman in Estarcion.

Bran MacMufin, formerly Bran Mak Mufin, the Pigt leader. He reappears as a clean-cut adulant, still spouting Pigtian prophecy as regards the earth pig god-king.

Jaka, Lord Julius’s niece, who is still looking out for Cerebus.

Euling S. Filgate, representative of the Docks district of Iest. He talks like a New Jersey gangster.

the artist/collector chap, another victim of Julius’s cost-cutting measures. The artist/collector was first seen in This Woman, This Thing.

King K’Cor, the anti-Venusian autocrat, making a rare appearance out of his city of Imesh.

Harmon Blakely II, a mostly mild-mannered elderly gentleman, but a powerful figure in Iest’s oligarchy.

Baskin, Julius’s personal secretary.

Suenteus Po, a different fellow than the Illusionist. He is an Anarcho-Romantic who takes a job as Cerebus’s speechwriter and historian, on the condition that he not be required to compromise his principles.

Lord Dunham, an influential leader in the Docks district of Iest.

The Abbess, head of Goodabbey in Iest’s Grace district. It is hinted that she is a Cirinist sympathizer.

Theresa, last seen in Swamp Sounds; now a young aide to Astoria.

Lord Storm’Send, a short, grizzled farmer who despises politics, but knows more about what’s going on in the conspiratorial background than most.

Lord Oldcastle, a friend of Filgate’s from the old neighborhood. Oldcastle reminds one of Rodney Dangerfield somewhat in speech and appearance (if he were drawn by Charles Addams).

6.2 Summary of High Society 45

The commander of the Parmoor district. The way this man speaks may remind one of any of several Monty Python sketches (and indeed, the military outposts in Parmoor have names like Fort Cleese, Fort Palin, and Fort Dead Parrot).

The commander of the Garrison district. The representative of the Iestan Guild of Government Employees, who is

the subject of one of Cerebus’s torture-analogies. Baldy Pate, a fence. Henrot, the wizard. He gazes into his crystal ball and gives Cerebus advice

based on what he sees therein. The commander of Mercenaries, a drinking buddy of Cerebus’s. Commander Pinshaw, commander of Iest’s armed forces. Duke Leonardi, leader of New Sepra, a Chico Marx look-alike.^{10} Arnold the Isshurian, commander of the mercenary cavalry. A very large,

hairy fellow with an adding machine in his brain. The director of the Bank of Iest. Hargraves, representative of the Anarcho-Romantic dissidents. The commander of the invading army.

6.2 Summary of High Society

High Society. Cerebus, after dragging his bag of loot fifty miles through 26 the swamps of Estarcion, unexpectedly finds himself getting the VIP treatment at the Regency Hotel in Iest. He is shown to the Ambassador Suite and given a tie to wear at dinner. Awaiting dinner, Cerebus finds himself pestered by a number of merchants, all offering him sizeable bribes simply to remember their companies’s names. He finally drags the real story out of one of them: the amount of bureaucratic red tape makes it impossible to do business straightforwardly, so highly placed aides, kitchen staff supervisors, and advisors must be approached instead. Cerebus, annoyed by the events of the evening, goes looking for someone to beat up. Just as Cerebus is about to enjoy a serious brawl at a tavern, the dock police put a halt to the proceedings, telling Cerebus that as a diplomatic representative, he is under their protection. Dejected, Cerebus walks back to his hotel.

10Chico was born Leonard Marx.

27 The Kidnapping of an Aardvark. Two hammer-wielding ruffians, announcing themselves as the McGrew brothers, Dirty Fleagle and Dirty Drew, attempt to grab Cerebus as he climbs up the steps to the Regency. Cerebus easily subdues them. When they explain their less-than-perfect plan to kidnap Cerebus for one hundred gold pieces, Cerebus takes them to an inn owned by an old acquaintance of Cerebus’s. Settling into a room with the McGrews, some writing materials, a crossbow, and two Diamondback card decks, Cerebus pens his own kidnap note demanding twelve thousand crowns in ransom from the Iestan government, signs it “the Eye in the Pyramid,” and delivers it by crossbow bolt. Cerebus then suggests a game of Diamondback to the McGrews. A narrative digression on the subject of Diamondback analyzes two variants of the game as they relate to socioeconomic conditions in the social strata and geographical regions where the variants originated. Meanwhile, the government, concerned about maintaining good relations with Lord Julius, goes through various upheavals in arranging to pay the ransom. A week later, after a very long game of Diamondback in which tempers get short and the McGrews inflict a bit of damage on the inn, Cerebus is notified of the signal indicating that the ransom has been paid. As he prepares to pick up the ransom, Fleagle catches Cerebus with his back turned and knocks Cerebus unconscious.

28 Mind Game II. Cerebus, while unconscious, finds himself drifting through a sort of cosmic cavern. He encounters once again the voice of Suenteus Po, who is meditating in the Illusionists’ Eighth Sphere. Cerebus and Po discuss Cerebus’s training in the meditative and sorcerous arts, the nature of the Seventh and Eighth Spheres, the interest in Cerebus possibly held by the Cirinists, the Kevillists, or the Illusionists, the relationship of the Orthodox Tarimites to the current political situation in Estarcion, and the state of the economy. Po warns Cerebus that the Western Church of Tarim might try to assassinate Cerebus as an Eastern Reformer, and that the Cirinists have become not only ruthless and brutal, but also unstable and unpredictable. As Cerebus finds himself drifting out of the Eighth Sphere, he malevolently plants the idea with Po that some conspiracy was involved in Cerebus’s “kidnapping.”

Repercussions. Cerebus awakens back in his room at the Regency. He is 29 delivered a note from the Iestan Prime Minister demanding reimbursement for the ransom paid for Cerebus’s return. Cerebus gives an equivocal reply. As Cerebus prepares to take a bath, a disembodied voice notes that no aardvarks have appeared in Iest in 800 years. When Cerebus investigates the source of the voice, he finds the Regency Elf floating in midair. She introduces herself to Cerebus as Elf, and informs Cerebus that the McGrews have been thrown into prison, but that the McGrews didn’t have any money on them when they were caught. Cerebus arranges a private conversation with the McGrews at the prison. After Cerebus casually beats them with a stick, they reveal that they had received nothing but a “dam statchoo of a dam duck,” which they had sold. Back at the Regency, Cerebus fends off another request for restitution from the PM, and converses with Elf. Cerebus writes a note asking for help from a Mr. Ned Greely, one of the merchants who offered Cerebus bribes at the start of High Society. Elf reveals that Greely is having an affair with his secretary and regularly has assignations at the Regency. Cerebus revises his note appropriately. The next day, as Cerebus plays wickets with the Elf, Cerebus receives an invitation to join the PM at the PM’s weekly pheasant hunt.

Debts. Cerebus and the Elf ponder the location of the ransom money and 30 the value of the duck statue. Later, at the pheasant hunt, the PM speaks candidly about why he is hounding Cerebus for an otherwise insignificant amount of money: a papal audit of the government’s books is looming, and the PM is short 750,000 crowns. He needs Cerebus’s “leverage,” such as that which Cerebus applied to Ned Greely, to help recover the money. The PM reveals that the Church’s lack of cash reserves and a few bad wheat crops are to blame for the audit. The PM suggests that Greely cannot help Cerebus, but that Holland M. Hadden, who first approached Cerebus with a serious bribe to consider gold-plated street lamps, can: if Cerebus can convince Hadden to cancel a debt owed Hadden by the Iestan government, Cerebus is off the hook. However, Cerebus finds that Elf has no juicy gossip for him this time. Cerebus instead gets Hadden drunk and convinces him to sign an agreement cancelling the debt in exchange for considering a proposal for solid gold street lamps and an implied promise of political influence. As Cerebus conducts Hadden to a carriage, Hadden is smashed underneath a heavy stone crescent dropped by the Merely Magnificent Moon Roach, who yells, “Unorthodox economic revenge!”

31 Chasing Cootie. Cerebus chases the Bug into the apartment of a woman who seems unexcited by the arrival of either one. When Cerebus asks if she is the Bug’s keeper, she assents. She tells Cerebus about how she met the Moon Roach, and that she fell madly in love with him. She found that the Moon Roach had an ungovernable separate personality, Kevitch, which drove the Moon Roach, so she built another, more malleable personality for him: Artemis Strong. The woman desperately plays on Cerebus’s sympathies to let the Moon Roach go, finally threatening suicide, and Cerebus relents. As Cerebus returns to the Regency, he is confronted by agents of the Inquisition who want to know about Cerebus’s involvement with Hadden’s death. Cerebus denies having received a signed and dated contract from the man. The agents tell Cerebus that his behavior is not viewed with favor, and that they will not be so nice next time. Then the agents are crushed beneath a stone crescent. Cerebus again goes in pursuit of the Moon Roach, but is unable to catch up to him. Instead, Cerebus yells at the Moon Roach that his girlfriend will be dragged to prison. As the Moon Roach flees, his thoughts name her Astoria, and reveal that she has already been moved to a new hideout. When Cerebus goes back to Astoria’s apartment, he finds it empty. Cerebus returns to his hotel room and pours his frustration out to the Elf.

32 Alliance. On the other side of the door to the Ambassador Suite, a messenger from the PM tells Cerebus that he must cough up 12,000 crowns or be thrown to the Inquisition. Cerebus, without scapegoat, money, or alibi, starts to sweat, but then he finds the Moon Roach and Astoria waiting in the parlor. Astoria offers him a “business proposition.” Cerebus snarls and attacks the Moon Roach. Just as Cerebus is about to brain the Moon Roach with a candle stick, Astoria tells Cerebus what to say to the PM: Cerebus is to authorize an increase in tariffs on Palnan goods entering Iest. Cerebus leaves off on the Roach, yells the suggestion through the door, and finds it accepted. Once things are calm, Astoria sends the Roach away and moves to assert control over Cerebus. She explains that if Cerebus obeys her, he will become wealthy as a means to her organization’s ends; if he resist, he will probably wind up dead. She retires behind a previously hidden doorway.

Later, on the wind-swept terrace of the Regency, Cerebus questions Astoria about Julius’s motives in appointing him Kitchen Staff Supervisor. Astoria speculates that Julius may have been grooming Cerebus as his successor. In the course of conversation, Cerebus tells her that all he wants is more money than anyone else has. Over nine days, Cerebus wonderingly watches Astoria demonstrate an easy familiarity with political machinery. Elf expresses her misgivings about Astoria to Cerebus. Elf’s lack of specifics frustrates and enrages Cerebus. Because he knows from past experience that getting mad at an elf is a bad idea, he struggles to contain his anger, but a strange psychic manifestation briefly leaps out of his head, snarling at Elf, frightening her. Elf indignantly disappears. Later, a package arrives, which Cerebus immediately throws into the fireplace, suspecting a trap laid by Elf. Just then, the Inquisition arrives to search Cerebus’s apartment for narcotics. When the Inquisition leaves, Astoria, having been stirring the ashes, pretends (or exaggerates) intoxication from the fumes and makes sexual advances toward Cerebus, which he rejects. Before he can disentangle himself, the Moon Roach returns.

Friction. Astoria smooths the Moon Roach’s ruffled feathers in the face 33 of some goading by Cerebus. Cerebus comments that Astoria seems to have sobered quickly, and Astoria retires, pleading wooziness. Later, Cerebus finds that Elf has scribbled “Cerebus wee-wees in the sink” in magically indelible glowing letters both in the hotel and on the gate to the city. The Roach, sitting on a building ledge in the rain, fantasizes that Astoria will come crawling back to him. In his jealousy, he decides to take some ill-conceived initiative. Soon afterward, as Cerebus tries to think of how to make up with Elf, he intercepts a message arriving for Astoria. The message says that the Governor of the Legislature has been crushed beneath a giant stone moon. When Cerebus gives the message to Astoria, she rushes out to deal with the crisis, asking Cerebus to cancel all her appointments for the day.

34 Three Days Before. Elrod comes to visit Cerebus, much to Cerebus’s displeasure. Cerebus knocks him out and arranges to have him put on a boat going anywhere far away. Before he can return to his room, Cerebus is greeted by a nattily dressed Bran Mak Mufin (or Bran MacMufin as it is spelled now), who quotes Pigtian prophecy to Cerebus and places himself at Cerebus’s disposal. When Cerebus returns to his room, he finds Elrod waiting, dripping water. Elrod suggests that he and Cerebus will be roomies. Before Cerebus can defenestrate Elrod, the Moon Roach appears in the window and threatens to kill anyone who tries to help Cerebus. When the Roach leaves, Elrod points out that Cerebus has many more supporters than the Roach can kill. Astoria returns, and Elrod chivalrously takes his leave, believing that Astoria is Cerebus’s mistress. Astoria asks Cerebus how he cultivated the friendship of “the most charismatic swordsman in Estarcion.” Astoria then finds out that Cerebus forgot to cancel her appointments, and that some very important persons felt snubbed as a result. Astoria claims that they are low on cash, although she has a large hidden supply. Cerebus leaves to get drunk and Elrod returns with an invitation for Cerebus to attend Lord Julius’s “Petuniacon” at which the guest of honor is His Highness King Elrod of Melvinbone, Lord Julius’s nominee to be the next ranking diplomatic representative of Palnu in Iest.

35 Two Days Before. Cerebus returns, and Astoria gives him pills to sober him, then grooms him and fits him for a new suit of clothes. Astoria and Cerebus meet with the PM to attempt to convince him that Cerebus is a better choice than Elrod. The PM explains that Iest must bow to Julius’s will, since Iest owes Julius so much money. As Astoria and the PM trade threats, Julius, disguised as a waiter, enters and engages Cerebus in Marxian dialogue on the side. Astoria drags Cerebus away. Later, as Cerebus changes into his new jacket and tie to dine with the PM, he is told that Jaka is there to see him.

36 The Night Before. Cerebus, seeing Jaka, puts on a snob act, boasts of how well-off he is, and then backhandedly invites Jaka to move in with him and Astoria. Jaka is clearly hurt by the way Cerebus is treating her, but Cerebus appears oblivious. Jaka makes a reference to her first meeting with Cerebus, and asks him when he finally remembered her. Cerebus contritely tells her that he had flashes of recollection when he was drunk, and once when he saw a dancer in Beduin. Cerebus temporarily lightens the mood by telling Jaka a joke, but the conversation takes a turn for the worse when Cerebus archly implies that Jaka is there for a handout, which Jaka denies. Cerebus tells her that from his experience as a mercenary, a tax collector, and professional gambler, he knows when someone is broke. Jaka admits that she spent her last hundred crowns on an antique. When Jaka explains the origin of the antique, Cerebus tries to make her feel foolish for buying it. Jaka makes her exit, rejecting Cerebus’s offer of money, but leaving the wrapped antique for Cerebus. Cerebus opens the wrapping to find his long-lost sword (last seen in Beduin by Night).

It’s Showtime. As Astoria chastises Cerebus for failing to make a better 37 effort in his own behalf, Cerebus blankly puts away his sword. Later, downstairs at the Petuniacon, while Astoria makes arrangements to have food served and for Cerebus to meet delegates in the Oak Room, Bran asks Cerebus why he is bothering with any political post lower than Prime Minister. Although Cerebus does not take the suggestion seriously, Astoria seizes on it and rushes away to investigate the possibilities. Filgate, a representative from “da docks” enters and solicits a bribe from Cerebus, but leaves unsatisfied. As Cerebus signs autographs for delegates, one delegate casually asks what it’s like living with Lord Julius’s ex-wife. Cerebus snaps his quillpoint. Recovering quickly, and using someone else’s quill, he turns the smear of his signature into a crude sketch of a tree. Seeing this, all the delegates go mad for “Cerebus originals.” Later, after the delegates have all left, the artist/collector from This Woman, This Thing confronts Cerebus. The artist complains that Cerebus inspired Elrod to draw hundreds of bunny sketches, thus cutting the artist out of a 5500-crown commission from Lord Julius to make serious sketches for the delegates.

Petuniacon Day Two. Cerebus sits on the “Us and Them: Eradicating 38 Them” panel with Julius, Elrod, and King K’Cor. Spies transcribe the proceedings for their various masters. Elrod proves himself a bit more adept at slinging the bull than Cerebus. Cerebus, infuriated at being one-upped and at hearing distorted versions of his earlier adventures with Elrod, challenges Elrod to a sword battle. Before Cerebus can get started on Elrod, Astoria drags Cerebus outside for a quick chat. She informs Cerebus that the government is about to collapse and that an opposition coalition wishes Cerebus to form a new party and run for Prime Minister. She also fills him in on all the Prime Ministerial perks. Cerebus acquiesces, but also arranges another private tavern brawl to work out his frustrations.

39 Petuniacon Day Three: The Ambassador Suite. Harmon Blakely II pays a visit to Cerebus to judge for himself the suitability of the Prime Ministerial candidate-to-be. As Blakely holds forth on the decline of the Church and the city-state, Lord Julius walks into the room wearing a towel and a bathing cap, and asks Cerebus if he can use the tub. Cerebus merely points the way to the tub, much to Blakely’s surprise. Blakely tries to puzzle out what ploy Julius may or may not be signalling. Julius, thus having succeeded in confusing everyone, takes his leave. A carriage driver enters, looking for the person who asked for a taxi. The McGrew brothers enter through the window. Fleagle angrily complains that he was unjustly incarcerated. The carriage driver, suffering a bad reaction to the McGrews’s presence, tears off his jacket to reveal himself as the Moon Roach. The McGrews, seeing a large, dangerous-looking fellow in a strange costume, decide that the Moon Roach is a “funny boy,” and a melee ensues. As soon as the McGrews lie in a heap, Cerebus gets rid of the Roach:

cerebus: Bug! Bug!! Mad financiers! Selling orphans to meet their stock margins!! moon roach: Where? Where? cerebus: Lower Felda!!

When Elrod appears, offering his hand to the losing diplomatic candidate, Cerebus slams Elrod around the room, and Elrod staggers away. Blakely, ignoring the interruption, speculates on the nature of any skeletons in Cerebus’s closet, particularly as they might be known to the mysterious anti-Cerebus magical graffito artist. The Elf appears and claims responsibility, but, having forgiven Cerebus, she cheers the idea of an aardvark Prime Minister.

40 Campaign. With Astoria and Bran as campaign advisors, and the Moon roach and the McGrews as bodyguards, Cerebus sets out on the campaign trail as the champion of “New Republicanism.” As Astoria negotiates votes from the Dock district, Bran fills Cerebus’s head with military strategies by which he could conquer Estarcion. Cerebus contemplates an “Aardvarkian Empire.” Later, the campaigners visit the Abbess of Goodabbey in the Grace district. The Abbess takes Cerebus aside and interviews him with a technique like a cross between free association and a Tarot reading. The Abbess offers her representatives on the condition that Cerebus dismiss Astoria. Crossing the river by ferry to Harbourside, Cerebus addresses the Suburban Merchants Guild, exhorting them to yell “Lower interest rates or death!” It is revealed that Julius is running a goat in opposition to Cerebus’s candidacy for Prime Minister.

Goat. Julius and his goat-candidate visit the Abbess, with whom Julius 40 conducts an amusingly flippant Marxian dialog, which is transcribed by spies.

Heroes. The Moon Roach, dragging himself about in a depression, delivers 41 to Cerebus an edition of The Times, the front page of which claims that Cerebus is trailing the goat. When Cerebus asks the Moon Roach what’s got him down, the Roach sobs that Astoria loves Cerebus, not him. Drew McGrew fetches Cerebus to confer with the campaign directors about how to counter the Times’s article. Meanwhile, the Roach and the Elf consume Cerebus’s whiskey, and the Roach pours out his sorrows to Elf. On the campaign trail, Cerebus plays cards with his staff; Bran reads a note from the Abbess in which she relents somewhat in the face of negotiations on the topic of Astoria; Cerebus demonstrates sophistication in political speech patterns. Back at the hotel, Elf tells Cerebus that the Roach received instructions from Astoria to search for an antique albatross—the duck statue. Cerebus, already exhausted from campaigning, stays up another night and day sending out letters inquiring as to the possible whereabouts of the albatross and receiving reports back. Just as Cerebus is about to give up and go to bed, Blakely notifies Cerebus that they have to mount a last-minute campaign in the northerly farming districts.

Campaign’s End. The campaign staff visits Lord Storm’Send, a short, rustic 42 old farmer who dislikes politics and voting. Filgate visits Lord Oldcastle, a friend of his from the old neighborhood. Cerebus offers money and soldiers to the frontier district-garrisons. A headline war breaks out between Lord Julius’s partisan paper, The Times, and Cerebus’s paper, The Sun.

Election Night. The campaigners await the election results behind the cur- 43 tains of a balcony seat in a public hall. Cerebus, having stage fright, refuses to address the hall. When Astoria fails to persuade him, Bran takes Cerebus aside and strokes his ego. This too fails to get Cerebus out on the balcony. Astoria invites Elrod to put in an appearance, but Cerebus throws Elrod over the balcony. When initial election returns are announced in favor of the goat, there is speculation that Julius fixed the election. Cerebus’s campaign staff considers leaving in case a victorious Julius decides to make arrests. Cerebus angrily announces to all that Cerebus will beat Julius in a war, if not in the election. Cerebus dispatches Fleagle and Bran to make arrangements for a coup. Soon afterward, Julius, hearing alarming reports of Cerebus’s men surrounding the hall, finds that he has no forces of his own with which to counter. The situation is saved when Astoria figures out that somehow the votes are tied, with one vote undecided.

44 The Deciding Vote. Cerebus, Astoria, and the Moon Roach take a horsedrawn sleigh back to Storm’Send. The Moon Roach fantasizes that he is Sergeant Preston of the Northern Iestan Mounted Police as he lashes the horses forward. Lord Storm’Send, after locking Astoria and the Moon Roach in his storage hut, tells Cerebus to don snowshoes and follow him to mend fences. Astoria and the Moon Roach discover Storm’Send’s homebrewed liquor supply, and proceed to get plastered and silly, running through a series of comic book character impressions. Storm’Send explains his opposition to the electoral process as Cerebus clumsily follows behind. Then it appears that Storm’Send, for someone who puts on a rustic act, is remarkably well informed: Storm’Send tells Cerebus that the albatross picked up by the McGrews as Cerebus’s ransom was a fake sent by the Kevillists to trip up Cerebus. Cerebus asks what the albatross is for. Storm’Send, surprised that Cerebus doesn’t know, asks Cerebus why he wants to be Prime Minister. Cerebus tells him, “For the money.” Cerebus frankly answers all of Storm’Send’s questions about reasons for the makeup of his campaign staff. Storm’Send, reaching a cliff with two torches, lights one of them to indicate how the deciding vote should be cast, but doesn’t tell Cerebus his decision. Cerebus punches Storm’Send, collects Astoria and the Moon Roach, and heads back to the Regency, only to find all the bridges out and the roads blocked by avalanches. Cerebus finds out he’s won the election from the manager of a hotel.

45 The Bureaucratic Rebellion. This story initiates a sequence called “Cerebus’ Six Crises,” the last of which ends High Society. Cerebus, once in office, quickly brings the representative of the Guild of Government Employees to heel by grasping the man’s hand and applying a simple analogy:

cerebus: When the Prime Minister—your thumb—is united with the bureaucracy—your fingers—great works can be accomplished. The lifting of spirits, the grasping of ideas. We hold the future between us...If, however, the Prime Minister and his bureaucracy are not united...like...so! representative: Gahh! cerebus: We have what? representative: Disunity! cerebus: What else? representative: Disharmony! cerebus: What else? representative: Agony!

Having obtained the everlasting loyalty of fingers to thumb, Cerebus then sends his old friend Baldy Pate out to fence the inauguration jewels to raise an army of mercenaries. Cerebus introduces to Blakely a violently raving Red Claw cultist as the new Minister for External Affairs. Cerebus consults Henrot on the matter of civil unrest. Cerebus tells Henrot that he plans to invade and plunder Lower Felda. The Moon Roach nearly scuttles Cerebus’s plans by apprehending Baldy, but Cerebus is able to send Baldy out again with the jewels. Cerebus finds an absurd casus belli, pumps up Commander Pinshaw, and turns loose his invasion. The invasion succeeds, but Lower Felda turns out to be broke and also in debt to Lord Julius to the tune of 689 million crowns. Lord Julius sends a 38 million crown bill to Cerebus to cover Lower Felda’s next interest payment, due within thirty days, along with an exploding cigar and and invitation to a summit.

“A Night in Iest” or “Summit Enchanted Evening.” Lord Julius, Duke 46 Leonardi, and Cerebus meet to divide Lower Felda in exchange for having Julius reduce Cerebus’s debts. However, Cerebus’s forces are being driven out of Lower Felda, which puts him in a poor negotiating position. Julius, speaking in pig latin, informs Cerebus that Leonardi’s New Sepra is vulnerable and very rich. Cerebus meets with his cabinet to tell them to expect economic recovery the day after next. Cerebus then meets with two bishops in the lower city, gets them drunk, and finds out from them the significance of the albatross: its holder wields control over both the Eastern and Western Churches. When Cerebus’s forces begin the invasion and looting of New Sepra, Cerebus collects almost four million paper crowns.

47 Balances. Suenteus Po, the Anarcho-Romantic, is hired as Cerebus’s speechwriter and historian. As Astoria and Po casually cross intellectual swords, Cerebus dozes. Later, when a manifesto of demands from a group of Anarcho-Romantic dissidents is delivered to Cerebus, and Cerebus utters the words “Something should be done about them,” the Cockroach, skulking outside Cerebus’s window, takes some initiative. The Cockroach locates the dissidents, announces himself as Sergeant Preston, and beats them up.

Meanwhile, Cerebus’s mercenaries capture the New Sepran treasury to find it evacuated and empty of money. Lord Julius notifies Cerebus that he is devaluating the New Sepran crown to a hundredth of its current value. The devaluation leaves Cerebus no money with which to have the mercenaries pursue the New Sepran treasury into the Red Marches. Cerebus finds himself futilely trying to figure out Julius’s strategy. He instructs Bran to take the Pigts into battle against the New Sepran outpost in the Red Marches, in the hope that the New Sepran treasury holds something of intrinsic value.

Filgate tells Cerebus that the Bank of Iest was forced to okay the New Sepran devaluation notice for fear that Julius would also devalue the Iestan currency. The Moon Roach, overhearing Cerebus complain that everyone is against him, forces the Bank’s director to sign a loyalty oath to Cerebus. This action angers the Bank into threatening devaluation of the Iestan crown unless they are given the power to oversee the PM’s spending. A battered Commander Pinshaw informs Cerebus that the New Seprans have driven out the Iestan forces. Cerebus signs a treaty in New Sepra’s favor. As the Anarcho-Romantics riot in the lower city, their representative demands that Cerebus accept an Anarcho-Romantic advisor and fire “Sergeant Preston.” Cerebus accepts the demands. Cerebus is informed that an unknown army is approaching. Cerebus initially believes that they are Bran’s forces, but two days later he is told that the army is a Hsifan invasion force.

Upstairs Downstairs. Cerebus directs the defense effort as best he can in the 48 face of absurdly trivial demands from the Anarcho-Romantics. Fed up with the Anarcho-Romantics and having more serious problems to resolve, Cerebus offers the Anarcho-Romantic representative, Hargraves, a choice among signing a revised agreement, leading an Anarcho-Romantic defense force, or having a hundred Anarcho-Romantics buy their way out at ten thousand crowns each. Hargraves signs the agreement. When Cerebus is left only with Blakely to decode and relay dispatches from the front received in the war room, Blakely decides that having the war room floor polished is more important than having the dispatches delivered in a timely manner. In the middle of dealing with riots and threatened defections, Cerebus is informed that Jaka is there to see him.

Jaka tries to get Cerebus to leave with her, telling him that “they” are writing off Cerebus and are going to throw Cerebus to the wolves. She asks Cerebus if he discusses his plans with Astoria. When Cerebus seems disinclined to escape, Jaka tells Cerebus that she loves him. She weepingly starts to paint a scenario of what will happen to Cerebus when his doors are finally broken down, but Cerebus, unwilling to listen, slaps her, and Jaka flees.

The Last Stand. Astoria is forced to cover for Cerebus when she finds him 49 slumbering under the influence of four bottles of whiskey. Julius arrives in the war room and ponders the situation board. Cerebus awakens and staggers back into the war room. The Moon Roach loudly announces that the PM has been drugged, and splashes Cerebus with water, thus causing the war room to be filled with the odor of Cerebus’s wet fur. Astoria tells Cerebus that they should take asylum in Lower Felda, but Cerebus thinks better of the plan. The McGrews meanwhile practice their Hsifan phrases. Cerebus, much to Astoria’s displeasure, puts on a brave face and decides to stay until the Hsifans arrive. Sword in hand, Cerebus runs out to confront the Hsifans alone. The point man of the invading force recognizes Cerebus, saying “You’re not an imposter!” The man leads a Conniptin cheer for Cerebus. The invaders turn out to be Conniptins friendly to Cerebus; the soldiers were once under Cerebus’s personal command (as seen in Day of the Earth Pig! and Swords Against Imesh).

Denoument. Cerebus awakens the next day to find that the Western Church 50 has taken over, that his staff has deserted to the Church, and the Anarcho-Romantics are under arrest. Cerebus considers having the Conniptins put up a fight, but Astoria points out that the Church has superior forces arriving soon from the Sepran Empire, and the rest of the city-states would unite against Cerebus. Astoria again proposes exile in Lower Felda and rebuilding. When Cerebus asks her why, she tells him that she feels rebuilding republicanism may lead to suffrage for women. Cerebus suffers a bit of rightous male wrath at that and tries to argue with Astoria. Finally, Cerebus tells Astoria he is leaving, not to Lower Felda, but to find Jaka. Astoria tries discourage him:

astoria: There are a few things you don’t know about her... cerebus: You mean like the fact that she’s Lord Julius’ niece? You’re not the only one who knows a few secrets around here, Astoria... astoria: I was Lord Julius’s “niece” once...he’s had quite a number of “nieces” over the years...the only difference was that he married \underline{me}!

Cerebus flinches, but continues on his way. Astoria roars that no one walks out on her, but Cerebus ignores her and returns to his room to prepare for his journey. Elf tries to convince Cerebus to stay, but Cerebus tearfully tells her no. They embrace. Cerebus dons his old jacket and his necklace, picks up his sword, then kicks the sack with which he originally arrived at the Regency. Amongst the junk that spills out is the statuette of an albatross. Cerebus throws the now worthless albatross over the Regency’s parapet and goes on his way. Elsewhere, Cerebus’s speechwriter, having refused to recant his heretical Anarcho-Romantic beliefs, writes his history/manifesto on the walls of his dungeon cell.

7 Exodus

51 Exodus is a one-issue epilogue to High Society. It is not included with the phone books. The reason for this is that Dave Sim felt including it in High Society would ruin the serious ending, and Exodus was too much an epilogue to High Society to include at the beginning of the subsequent novel, Church & State.

Cerebus takes passage in the hold of a river barge going to Togith, but it turns out that he isn’t alone: Lord Julius and Elrod appear out of two barrels in the hold, and Duke Leonardi drops in, falling squarely onto Cerebus. The unexpected presences of his former political rivals make Cerebus’s blood pressure rise. The situation is temporarily saved when the Purely Platonic Priest Roach (The Cockroach, trying on yet another persona) puts in an appearance, looking for fugitives and heretics on behalf of the Church. Almost everyone scrambles into the barrels before the Priest Roach enters the hold. With no room left in the barrels, Leonardi is left alone to contend with the Priest Roach, but the Priest Roach proves gullible enough even for Leonardi to outtalk him.

When the Priest Roach leaves, Cerebus offers to slay everyone, but then Lord Oldcastle drops into the hold, landing on Cerebus and squashing him flat. It takes a few moments, but Cerebus recovers.

elrod: Be–I say–be reasonable, son...

cerebus: Reasonable? Cerebus is tired of being reasonable...Cerebus is going to try homicidal instead! Cerebus is going to chop all of you up and make bureaurat stew...

Before he can carry out his threat, everyone is buried in a load of potatoes, coincidentally containing the McGrews. Cerebus crawls out and takes advantage of the ship’s docking in Port Seprania to make his exit.

8 Church & State

Although the series never completely loses its light touch, by the time of Church & State, Cerebus graduates to a kind of Swiftian satire of the darkest sort. The later chapters in Church & State mix in more surrealism (this may seem strange to say—as if a funny animal isn’t surreal enough), as Cerebus’s odd dreams are prominently featured, often serving to foreshadow events to come.

Church & State, a 1219 page work in two volumes, is a satire on—guess what subject? The theme seems to be “You can get what you want and still be unhappy,” a little lesson delivered by Cerebus himself to one of his constituents. Some parts of the story are incomprehensible without reading previous works, but it is possible in general to read and enjoy Church & State on its own.

The story is set in United Feldwar States, a loose federation of city states arranged and held together by President Adam Weisshaupt. The religious and political establishments are at odds, but maneuverings by President Weisshaupt tend to keep things under control. The capital of the UFS is the city-state of Iest, which has remained under nominal occupation by forces of the Sepran Empire since the end of Cerebus’s first administration. Greater Iest, where the ruling class lives, is built partly on the plateau of an ominous, black, truncated tower whose sides appear to be composed of huge demon heads and skulls. A road leading from Greater Iest to Lower Iest spirals around the tower. Various mini-communities along the road are built against the hillside; almost no one comments on the staring, gaping, demon faces. The other states in the UFS are Lower Felda, New Sepra, Palnu, and Togith. Iest is the largest of the city-states.

The religious establishment, the Tarimite Church, is divided into two sects, Eastern and Western. The Eastern sect, headquartered in Iest, has a formal hierarchy including a pope, archbishops, and priests. The Western sect is run by The Lion of Serrea. In spite of the division between the Churches, there is an arrangement that gives the Lion of Serrea, or his representative in Iest, a high degree of control over the Eastern Church. The deity of the Churches is named Tarim; heaven is called Vanaheim. There otherwise appear to be many significant similarities to Christianity as it was practiced in our Middle Ages. In particular, the infallibility of the Pope is not to be challenged, heresy is punishable by death, and fornication is forbidden. In other respects, of course, there are many satirically presented similarities to Christianity even (and, perhaps, particularly) as it is practiced by various sects today.

The heresies of most concern in Iest are Cirinism and Kevillism. Cirinism is a political and religious philosophy in which the chief deity is the female Terim rather than the male Tarim, and in which society is dominated by women, with only mothers holding franchise, and in which men may be property. Cirin is the name or honorary title of the leader of Upper Felda, a powerful Cirinist state that poses a military threat to the UFS. Kevillism is a philosophy alleged to have been devised by Astoria. Kevillism encompasses Cirinism, but deviates from Cirinism (unacceptably to Cirin’s government) in that Kevillism also advocates state-owned prostitution, guaranteed minimum incomes for women, a mother’s right to determine life or death for her children without interference from the state, and the right to self-determination for women.

Initially, the conflict and competition in Church & State appears to be solely for wealth and political power. The underlying conflict that eventually emerges is that of the struggle between the dualities of the universe, yin and yang, male and female. Depiction of this struggle, even to the point of inventing a new creation myth, is not a new concept in fantasy (or religion); the trick, as always, is to give it a new framework. In the case of Cerebus, the threads of the conflict can be seen early on, in the conspiracies of Cirinism (the advocates of a female deity, Terim) against the Orthodox Tarimites (advocates of a male deity).

Images of the Tarot (also used as Diamondback cards), such as the priestess, the queen, the sword, and the scepter, appear from time to time in Church & State. In particular, each major character is auditioning for a role of priest, priestess, magician, king, or queen, in whichever religious faction emerges supreme, or possibly in whatever results from a merger of both.

8.1 Cast of Church & State

The main characters appearing in Church & State are (in order of appearance or mention):

Cerebus, recently resigned as prime minister of Iest. He has taken to writing a book, On Governing.

Lord Julius, still the leader of Palnu. He has acquired a duck-walk, a round cap, and a large sword whose scabbard is emblazoned with the letters “U.F.S.” Julius is the primary target of wrath in Cerebus’s book.

Duke Leonardi, somehow still the leader of New Sepra.

Countess Michelle Detin, a wealthy widow (or so she says) living alone in her mansion in New Sepra. She knows many important persons, but nevertheless has tax problems with her holdings in Palnu. She is a gracious hostess, but Cerebus tries her patience, primarily because Cerebus can be almost completely blind to social cues.

Countess Michelle Detin

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

The Secret Sacred Wars Roach

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Vichy Dupris, a persistent, unwanted suitor of Michelle. Vichy is an Onliu Anarcho-Romantic.

Artemis Roach, aka Wolveroach, aka Secret Sacred Wars Roach, now possessed and sometimes controlled by an incorporeal manifestation of Professor Claremont.

The Wolveroach

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Astoria, now thought to be the architect of Kevillism, Cirin’s one-time chosen successor, and an exile from Upper Felda. Although she played a major role as an advisor in Cerebus’s first administration (Cerebus would not make a move without consulting her), Astoria stays in the background for the first half of Church & State.

Adam Weisshaupt, the bewigged schemer last seen in The Death of Elrod. Weisshaupt is still President of the UFS, but the UFS’s government is no longer “in exile.” Weisshaupt engineered the formation of the UFS by promising to make the leaders of the component city states wealthy beyond their wildest dreams. It is Weisshaupt who finds a way to force Cerebus to take the position of prime minister once again, but this time as Weisshaupt’s puppet.

Mrs. Henrot-Gutch

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Adam Weisshaupt

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

The Abbess, Astoria’s conversational foil while Astoria takes refuge in Goodabbey.

Red Sophia, that sword-wielding barbarienne from Cerebus’s past. Sophia is used as the trap that serves to get Cerebus’s political career going again. Sophia would be a throw-away character if it weren’t for her bedtime chats with Cerebus, in which she gently points out Cerebus’s hypocrisies. She is allowed to grow up a little bit in Church & State, although she never loses the plate-mail bikini.

Mrs. Henrot-Gutch, Sophia’s mother, a vicious troll of a woman who serves as a comic distraction to all. She is somehow able to mock and even to strike Cerebus with impunity. To know her is to loath her.

Harmon Blakely II, formerly of Cerebus’s cabinet. He is introduced to Mrs. Henrot-Gutch at a private dinner with Cerebus. He becomes a tentative suitor of Mrs. Henrot-Gutch, but comes to know her and flees from her.

Pope Harmony IV, soon to be the late Pope Harmony IV.

Elf, one of Cerebus’s few friends, but Cerebus is usually in no mood to talk to Elf. Sometimes Elf appears in Cerebus’s dreams.

Bishop Powers, representative in Iest of the Lion of Serrea. Bishop Powers is given the task of selecting a pontiff of the Eastern Church following the execution of Pope Harmony IV, and of attempting to bring the power of the Church to supremacy over Weisshaupt’s government. Weisshaupt and Powers fear and perhaps respect each other, but things get chilly and dangerous when they meet.

Theresa, an informant to Weisshaupt, formerly an aide to Astoria.

Archbishop Posey, aide to Bishop Powers, later an unwilling aide to Cerebus. Cerebus describes Posey to his face as a limp-wristed pucker-faced sack of suet.

Bear, a large, bearded, muscular fellow who serves as bodyguard and manservant to Cerebus. Bear is somewhat hesitant of speech. Bear also made an appearance in Cerebus Jam.

66 8 CHURCH & STATE

Red Sophia

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Bishop Powers

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Boobah, Cerebus’s relatively skinny bodyguard, scribe, and manservant. Boobah is always thinking of food, which, in combination with his clumsiness, tends to get him into hot water with both Cerebus and Mrs. Henrot-Gutch.

The big round glowing white strange thing, one of a series of magical, mystical things that happen to Cerebus.

Bran MacMufin, the Pigt leader, reappearing yet again as a clean-cut adulant. He believes absolutely in Cerebus’s holy perfection. But Bran is also a spy of a sort for his people.

Jaka, making a brief, but powerful, reappearance in Church & State. Although it has been revealed that she is one of Lord Julius’s nieces, she is determined to stay independent; she remains a tavern dancer.

Thrunk, aka Necross the Mad, is a giant, animated, stone creature.

Fleagle and Drew McGrew, now sidekicks to Artemis Roach. They wear white costumes each with a big black cockroach painted on the front, and sweat a lot around Artemis, because Artemis treats them like boot camp recruits.

Henrot, the magician, appears briefly as a “talking head” at the Floating Floor in Space in Church & State volume 2.

Fred, Ethel, and the little fellow with the hair, aka the Trinity of Tarim, are a huge creature with one body and three heads that first appears on the Floating Floor in Space in Church & State volume 2. All three component pieces of the creature, Filbert, Woman-Thing, and the artist/collector chap, first appeared in This Woman, This Thing, issue #25.

Cirin, leader of Upper Felda, a Cirinist state that is the chief military threat to the UFS. She is a sorceress of sorts. Weisshaupt describes her as “the cow that walks like a man.”

68 8 CHURCH & STATE

Boobah

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Theresa

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

8.1 Cast of Church & State 69 Bran MacMufin

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

70 8 CHURCH & STATE

Jaka as an adult and as a child in Cerebus’s dream

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

8.1 Cast of Church & State 71

Bear

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Fred, Ethel, and the little fellow with the hair

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

The Judge, aka The Man in the Moon, aka George, a paunchy humanoid being in black judge’s robes who resides on the moon. He has been watching Earth since before humans appeared.

The Judge

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

8.2 Summary of Church & State

8.2.1 Volume 1

52 Writing. The story opens in a tavern, where Cerebus is writing his book, in which he analyzes what went wrong with his administration and recommends various harsh methods for avoiding similar problems in the future. Cerebus, by and large, is completely disgusted with the machineries and trappings of government. Lord Julius, leader of Palnu, is given prominent mention in the book. While writing, Cerebus gets into a fight with a lout who spills Cerebus’s drink on Cerebus’s notebook.

The Insecure Sinecure.^{11} This chapter parodies Prince Valiant. Cerebus is 52 visited in a tavern by Lord Julius’s wastrel son, Lord Silverspoon, and by two sycophantic friends of Silverspoon’s, Lords Gwain and Trystrim. Silverspoon dictates a letter inviting himself and Cerebus to the estate of Countess Detin, claiming to be able to help her with her Palnan tax problems. When the Countess sends funds (payable to Cerebus) for the journey, Cerebus cleverly manages to ditch Silverspoon.

The Countess & the Aardvark. Countess Michelle Detin receives Cerebus 53 on a rainy evening. Over drinks, the subjects of Cirinism and Kevillism, Astoria, Weisshaupt, wealth, and Lord Julius are mentioned, but are not discussed in detail. The countess asks Cerebus how much money he would want if he could have any amount at all. “All of it,” replies Cerebus. The conversation is interrupted by the arrival of a rain-soaked suitor of Michelle’s, Vichy. While Michelle argues with Vichy, Cerebus goes upstairs to work some more on his book. The next day, Cerebus offers to fix Michelle’s tax problems by suggesting a bribe for certain officials at the Palnan embassy. Later that night, Michelle’s “Uncle Artemis,” aka the Cockroach, wearing a new “Wolveroach” costume, crashes his way into the house and collapses.

The Origin of the Wolveroach. After putting Artemis to bed, Michelle puts 54 Cerebus’s head in her lap and strokes his forelock. She tells Cerebus how she came to know Artemis and how Artemis was used by Astoria and Weisshaupt. The conversation is interrupted by the arrival of a smug Weisshaupt, who asks Cerebus if he’d like to be Prime Minister again. Weisshaupt tells Cerebus that he has gained the cooperation of several city-states to form the United Feldwar States. He further tells Cerebus that the UFS was formed for the express purpose of making those in power wealthy. Cerebus elicits an explanation from Weisshaupt as to how Artemis fits into the plans: Artemis, while appearing to be a champion of the oppressed, will conveniently assassinate any member of the oligarchy whose excessive greed poses the danger of arousing the populace. Cerebus declines the job of Prime Minister in detail, and Weisshaupt leaves. Suddenly, Michelle realizes that she put Artemis in Vichy’s room. On cue, there is the sound of something nasty happening upstairs.

11Issue 52 also contains the story “Elfguest,” described in section 4.4.4 on page 13.

55 The Why & the Are. Vichy is discovered to have been beaten up by Artemis. Michelle asks Cerebus to talk to Artemis. Cerebus ill-temperedly agrees. Finding Artemis, Cerebus tries to warn him that he is being used by Weisshaupt. Artemis smokes, loudly sucking and exhaling smoke through his nostrils.

cerebus: Weisshaupt’s doing it to you again, Bug...only this time he’s making you think it’s your idea...You can’t trust anyone that you think is your friend...they’re all secretly his agents... Do you know what Cerebus is trying to say...Artemis? artemis: In a world of chaos—against overwhelming odds—the Wolveroach stands...alone! cerebus: What the hell are you smoking? artemis: A small bundle of toothpicks...they serve to increase the strength and stamina of my wondrous Wolverlungs to their full Wolvercapacity...

Cerebus gives up. Cerebus’s refusal to try harder with Artemis annoys Michelle. Later, when Cerebus is writing alone in his room, Michelle enters and angrily asks Cerebus what he has done with Artemis. Cerebus yells back that he has done nothing with Artemis. The ugly little scene is interrupted by a knock at the door. A soldier at the door warns Michelle to stay inside, as they are looking for Artemis, who apparently earlier tried to kill the Iestan ambassador.

56 The Mystic We. Cerebus comforts Michelle as they wait for Artemis or the soldiers to prevail. Artemis proceeds to wipe out the soldiers. A disembodied voice speaks to Cerebus, directing him to go to the basement, where Artemis is waiting and smoking another bundle of toothpicks. The voice turns out to be a spirit manifestation of Professor Charles X. Claremont (seen crushed to death by Apocalypse Beasts in This Woman, This Thing), who is inhabiting Artemis’s head. Claremont asks Cerebus to take his leave of Michelle, explaining accurately (and, he claims, all-knowingly) that not only does Michelle not love Cerebus, but Cerebus was planning to run preemptively from Michelle anyway. Claremont manages to trick Cerebus into alienating Michelle completely by convincing Cerebus to try to exert control over Michelle. Michelle gives Cerebus money, and he leaves.

Suddenly, Sophia. Cerebus awakens in bed with a hangover, next to Red 57 Sophia. Weisshaupt enters the room and informs Cerebus that he is now married to Sophia, which fact Sophia verifies by displaying her wedding ring. Cerebus accuses Weisshaupt of drugging him, which Weisshaupt admits having done. Before Cerebus can leave, Weisshaupt tells him that wife abandonment is a capital offense, and that only Weisshaupt can grant Cerebus a divorce. Weisshaupt offers Cerebus a choice: he can accept the position of Prime Minister for a few years, occasionally help Weisshaupt to write anti-Cirinist romance fiction, and then be granted a divorce, or he can stay forever trapped in Iest, powerless, and married to Sophia. Cerebus unhappily acquiesces to Weisshaupt’s terms. Once dressed, bewigged, and waiting for his carriage, Cerebus receives a nasty introduction to Mrs. Henrot-Gutch, Sophia’s mother.

All Lined Up. During a formal, private dinner with Harmon Blakely at the 58 Regency, Cerebus and Mrs. Henrot-Gutch fight childishly. Later that night, in bed, Cerebus and Sophia have a tiff about Sophia’s mother. A servant fetches Cerebus out of bed to meet with Pope Harmony IV and a representative of the Lion of Serrea. The Pope, obviously world-weary, cryptically warns Cerebus that Weisshaupt plans to take control of the Church, and thus wield absolute political and religious power. The Pope reveals himself as a Cirinist sympathizer, and entreats Cerebus to stop Weisshaupt before a military confrontation with Cirin develops. The Pope’s admission of heresy causes the representative of the Lion of Serrea to execute the Pope on the spot. Cerebus, shaken, leaves the audience hall, and ignoring an attempt by Elf to talk to him, he returns to bed.

Carroll E. King Reads is a humorous interlude indicating that Weisshaupt’s 59 romance fiction may be finding the wrong audience. One of a set of vignettes poking fun at the pretensions of the upper class.

59 At the Club One Afternoon is another humorous interlude reminiscent of Being There: Cerebus’s usual saturnine converse is taken for an admirable brevity of speech.

59 Memoirs has Elf reading the fifth chapter in the partial draft of Cerebus’s book and making the same amusing comment made by everyone else who has read the book: “I liked the part about the butler.”

59 Note is a one-page interlude, consisting of the torn scraps of a letter to Cerebus from Jaka, writing on behalf of Lord Julius, formally congratulating Cerebus on his marriage and reappointment as Prime Minister.

59 First Impression introduces Mrs. Tynsdale-Clyde, who faints when Cerebus’s tail most improperly rips out of his ill-fitting trousers.

59 Tree Planting. Euling S. Filgate meets his match in Mrs. Henrot-Gutch at a minor ceremony.

59 Approved By shows Cerebus signing documents like a good puppet Prime Minister, without wanting to know what he is signing.

59 Rough Pope is a confrontation between Bishop Powers and Weisshaupt concerning the separation of power between church and state.

60 Sophia. Cerebus and Sophia have a bedtime conversation in which Sophia gently skewers Cerebus’s bad temper, complaints, and apparent hypocrisy in staying Prime Minister. Cerebus says nothing.

60 Henrot-Gutch reveals another loathsome facet of Sophia’s mother, frightening away Harmon Blakely, who was otherwise enjoying tea and a quiet chat with her.

60 Astoria. Astoria, conversing with the Abbess, ridicules both Weisshaupt’s political maneuverings and Cirin’s reactions to Weisshaupt.

60 Elf has Elf trying to amuse Cerebus as Cerebus is retiring for the evening by telling Cerebus a joke she heard about him. She is not successful.

60 Theresa. Theresa, now a mature woman, pitches her services as an informant to Weisshaupt.

Stormy Weather is a humorous interlude in which Sophia, that brawling 61 barberienne, is shown huddling and shivering against Cerebus in bed during a thunder storm. A particularly loud peal sends Mrs. Henrot-Gutch running into their bedroom for comfort.

Cerebus Was and Something Cerebus Was have Cerebus working on his book 61 and trying to figure out how to stop Weisshaupt without joining forces with Julius.

Mrs. Tynsdale-Clyde’s Tea. Mrs. Tynsdale-Clyde has an amusing interpre- 61 tation of events in First Impression.

Powers. Bishop Powers contemplates whom to choose as the successor to 61 Harmony IV. After thinking about it for a while, he decides he’s figured out how to get Weisshaupt by the short and curlies.

Boom. Weisshaupt and Cerebus have an evening drink together. Weisshaupt 61 compliments Cerebus about going along so well with the whole UFS scheme, and offers to answer any question that Cerebus cares to ask. Cerebus asks Weisshaupt how he managed to get Lord Julius to join the UFS. By way of an answer, Weisshaupt takes Cerebus up to the roof. They approach a black, bulky object, barely visible in the darkness. Weisshaupt explains that he found a way to guarantee the safety of Palnu, and demonstrates it by setting off a huge cannon—a terrifying weapon of destruction previously unknown in Estarcion.

Appointments. Theresa briefs Weisshaupt on the events in Upper Felda 62 that led to Astoria’s exile. Theresa tells Weisshaupt that Astoria was being groomed as Cirin’s successor. However, Astoria found herself at odds with the bureaucracy, and had developed a stubborn streak of independence. Astoria eloped with one Sir Gerrik, Cirin’s adopted son, and quickly became pregnant. However, during her preganancy, Astoria disappeared for a time. When she reappeared, she was no longer pregnant. Although proof of a live birth would grant Astoria certain rights of citizenship, Astoria refused to answer questions about whether the child was alive, asserting that it was not the State’s business. She created a scandal by asserting that the state had no business interfering in a mother’s relationship with her child.

The briefing is interrupted by Lord Julius and Duke Leonardi, who express a desire to withdraw their city-states from the UFS. That conversation is in turn interrupted by Skorz (first seen trying to catch Cerebus’s attention at the beginning of High Society), resulting in a brief Marxian routine:

scorz: I am here for to talking at you “sewage, sewage, sewage.”

julius: Well, you’ve come to the right place.

scorz: All over the states of the coming together Feldwar is nothing but sewage, sewage, sewage. leonardi: Who is dissa guy...? julius: I don’t know, but so far he hasn’t said anything I’ve disagreed with...

Before the argument over maintaining the federation can resume, Weisshaupt is informed that Bishop Powers has selected Cerebus to succeed Harmony IV as Pontiff of the Eastern Church.

63 Mind Game IV. Cerebus sits alone, drinking heavily and letting his mind wander. On the last page, Cerebus is informed that he has been nominated to be Pontiff of the Eastern Church. Note: The artwork includes twenty panels that must be rearranged like pieces of a puzzle to form a single picture. The picture shows Cerebus, dressed in papal vestment, standing contemplatively in shadow with his back to two tall stone icons, one a simple cross, the other similar to an ankh, or perhaps the symbol used to mean “female.^{12}” Discontiguous lettering in the background spells “Mind Game III.”

64 Never Pray for Change shows a relatively civil conversation between Bishop Powers and Weisshaupt about the choice of Pontiff. The conversation is interrupted by Mrs. Henrot-Gutch’s homicidally-intentioned search for Harmon Blakely, who has managed to enrage her. A subsequent fight between Cerebus and Mrs. Henrot-Gutch is interrupted by Weisshaupt, who attempts to sweet-talk Cerebus, who as Pontiff is now beyond Weisshaupt’s control.

12Various cross-like and ankh-like symbols appear in the backgrounds and adorn panels throughout Cerebus, occurring most frequently in the first half of Church & State. There are two such symbols that appear to be affiliated with the Tarimite Church; both vary in their representations.

Weisshaupt leaves most displeased. Cerebus’s papal robes are delivered by Archbishop Posey, a timid soul who gets the blunt end of Cerebus’s now unleashed frustration.

Anything Done for the First Time Unleashes a Demon. Bishop Powers sus- 65 pects that he has made a dreadful mistake, and sends Archbishop Posey back to deliver a message to Cerebus. Cerebus has moved himself, his family, and two guards, Bear and Boobah, into a hotel halfway down the eastern slope of Iest. A crowd has gathered in front of the hotel. Posey gains an audience with Pope Cerebus, who finds an amusing way to force Posey’s allegiance. Cerebus, now “Most Holy”, addresses the crowd. Most Holy tells them how it really is with Tarim:

If you are poor or weak or sick or crippled, then, there is not much hope...the best Most Holy can do is to put in a good word for you with Tarim. However, it is going to cost you...Tarim’s mercy does not come cheap! You must give Most Holy all of your gold! No jewelry, no paper money! Gold!

The Thrill of Agony and Victory of Defeat. Weisshaupt is aghast at the news 66 of Most Holy’s demands. Meanwhile, Most Holy brutally delivers the lesson “You can get what you want and still be unhappy:” he blesses a child at the request of its mother, then throws the child over the heads of the crowd. As Most Holy’s gold reserves grow, an economic crisis looms, but Most Holy does not relent. In fact, to motivate his followers to cough up more gold, Most Holy rehearses a speech in which he plans defenestrations of persons old or infirm, babies, women, and mothers. Posey, listening to this, get sick in the bushes. As Most Holy and Posey converse, a glowing white light approaches and briefly forms a halo over Most Holy, then vanishes. Bran MacMufin arrives at the hotel.

Day of Greatness, Age of Consent reintroduces Bran to Most Holy. Bran 67 manages to excuse himself for failing to muster the Pigt army to save Cerebus’s previous administration. Most Holy has another tiff with Sophia. Astoria argues with the Abbess about the problems created by Weisshaupt’s federation and Cerebus’s papacy. The economic crisis deepens. Most Holy delivers a speech in which he claims that Tarim will destroy the world in ten

days unless all the gold is delivered to Most Holy.

68 Another Thing Coming. Most Holy delivers another one of his brutal speeches, reiterating Tarim’s threat. Weisshaupt blows a gasket. Most Holy develops a head cold, and is pulled from his sick bed to be addressed by Weisshaupt from the roof of the building across the street. Upon the roof are set all of Weisshaupt’s cannons, pointing directly at Most Holy. Weisshaupt demands all the gold.

69 Ignore It, It’s Just Another Reality. Most Holy delivers a threat that causes Weisshaupt’s supporters to desert Weisshaupt in fear of their mortal souls. Weisshaupt becomes so furious that he collapses of a heart attack. He is helped back to his carriage in defeat. Later, Most Holy is visited again by the glowing white light. Most Holy strikes it, only to be zapped, after which the light disappears, leaving Cerebus still sizzling and popping. Most Holy then sneezes a ferocious stream of fire, burning through a supporting beam. Pleased with this new ability, he decides to try it out on Mrs. Henrot-Gutch. It doesn’t work; he sneezes on Mrs. Henrot-Gutch and is clobbered. Most Holy then receives a summons from Elf.

70 Sane As It Ever Was. Following Elf’s instructions, Cerebus traverses a secret passage filled with traps from his hotel to the Regency. He arrives bruised and dirtied. In a darkened room, Elf conducts a conversation with Most Holy. Elf tells Most Holy that Most Holy must stop what he’s doing. However, the floating image of Elf isn’t really Elf, and Most Holy apparently isn’t fooled.

71 Hovering Below the Fray. Weisshaupt’s illness leaves a power vacuum in the UFS. Bishop Powers and Hedges, Weishhaupt’s aide, discuss the succession. Bran and Most Holy converse about the nature of Most Holy’s perfection and about Most Holy’s plans. Lord Julius confers with Bishop Powers about the Presidency, suggesting that Bishop Powers assume the job.

72 Time Passes and Life Manoeuvres. Lord Julius apparently confuses Bishop Powers long enough to get himself nominated for the job of President. Most Holy announces that Tarim is going to destroy the world no matter what, and the populace’s one chance to save itself is to conquer, utterly without weapons, the Red Marches—quicksand, black flies, and all—in forty-eight hours. The populace complies; Bran sends word to his people by carrier pigeon. Sophia and Most Holy have a final tiff. In the lobby, where Most Holy’s gold is stored, gold coins fly into Most Holy’s hand and fuse, forming a glowing gold hemisphere, and frightening Most Holy.

Fascination & Fruition. Sophia packs up and leaves Most Holy. When Cere- 73 bus goes down to the cellar for whiskey with which to drown his frustration, the glowing white sphere reappears and for the first time speaks to Most Holy, reminding Most Holy about Jaka. Most Holy sends Bear to get Jaka, who is dancing in a tavern in the lower city. Upon his return, Bear surprises Most Holy with the information that Jaka is now married, and she had wanted to bring her husband along.

Acquired Tastes. Jaka is brought into Most Holy’s presence. Cerebus rants 74 and raves about Sophia. Forcing Jaka to her knees, he asks Jaka why she didn’t wait for him. Jaka cries that she had promised to wait for Cerebus to remember—and that Cerebus remembered, but never came back. She reminds Cerebus of how they met, and explains her feelings at the time. Cerebus asks Jaka to stay with him, but she gently refuses by offering an analogy about how her Uncle Julius’s power changed her relationship with her family. Cerebus considers, then dons his old jacket and sword, and tells Jaka, “Let’s Go.”

Terrible Analogies. Jaka refuses to leave her husband. Cerebus offers to dis- 75 solve her marriage. Finally, Jaka tells him that she is pregnant, and she hides her face in her hands. Cerebus dejectedly drops his sword and gives up. Bran enters to inform Most Holy of Most Holy’s victory in the Red Marches. Bran sees Jaka and retreats apologetically. Cerebus and Jaka embrace tenderly, and Cerebus lets Jaka go away, back to her own life. Cerebus sits down to sulk.

Varying Reasons of Assorted Depths. Most Holy makes an ill-tempered visit 76 to Weisshaupt’s deathbed at Weisshaupt’s request. Most Holy taunts Weisshaupt cruelly. Weisshaupt tells him that there are two other aardvarks in Estarcion, but does not reveal who or where the other aardvarks are. Weisshaupt also mentions that Suenteus Po (the Illusionist, one must presume) is Weisshaupt’s uncle. When Weisshaupt is close to death, he suddenly sees some metaphysical answer or some quality of divinity in Cerebus. Cerebus tries to force an answer from Weisshaupt about the other aardvarks.

Weisshaupt asks Cerebus for absolution. Cerebus, unable to get the dying Weisshaupt to answer his question, leaves in anger after slapping Weisshaupt and telling Weisshaupt to go to Hell.

77 Odd Transformations details one of Most Holy’s dreams. As he wades through water, a puffy-eyed human peasant woman approaches, yelling, “How many times do your father and I have to tell you...nobody wants to look at those ears. Put y’damn hat on...” Escaping, he flies through the air, eventually colliding with a tree. He visits the moon and makes love to Jaka. He awakens, and the reader is then treated to several pages of Most Holy visiting the jakes.

78 Odd Tranformations Part Two continues Most Holy’s night of dreams, this time on an empty bladder. This dream has Lord Julius mentioning that “Astoria’s busy changing the baby...she’s trying to change it into another aardvark.” Most Holy meets Woman-Thing and stabs her with a short sword. Most Holy walks on water, only to be thrown over a waterfall. Becoming a child, he sees Astoria and Michelle arguing. Michelle says that she is Lord Julius’s nice niece. Older again, he sees Jaka both as an adult and as a child. Cerebus is awakened by a short, old fellow in a farmer’s cap (Lord Storm’Send, who cast a deciding vote for Cerebus in High Society) who now has some things to tell Most Holy.

79 Spinning Straw into Gold. Most Holy instructs Bear to get rid of Storm’Send and returns to bed. But Lord Julius invades his bedroom. Bear, woken again by Most Holy’s bellow, is subjected to some slapstick pranks by Lord Julius. Bear quits in disgust, leaving Most Holy alone to contend with Lord Julius and the persistent Lord Storm’Send.

80 Causality Casualty. Lord Storm’Send hints that Cerebus has set something big and uncontrollable in motion. Just as Storm’Send prepares to tell Cerebus about what is happening to the mountain on which Greater Iest sits, Sophia returns and interrupts the conversation. Sophia tells Cerebus that she found marriage and sexual relations with the Pope of the Eastern Church to be perverse. Sophia in turn is interrupted by Lord Julius, who reintroduces Most Holy to Thrunk (last seen in Black Magiking), a giant stone creature in papal robes. Thrunk smashes his massive fists through the hotel wall to get at Most Holy. Most Holy runs to Bran’s room. Bran, in despair, stabs himself in the chest and dies. Thrunk finally seizes Most Holy and flings him over the rooftops into the lower city. Cerebus takes several bad jolts against the cliff’s demonic faces. Managing to slow his fall a bit, he eventually slams into the ground.

8.2.2 Volume 2

Becoming Synonymous with Something Indescribable. Cerebus regains con- 81 sciousness to see Artemis, who has a new, all black costume with a roach logo on the front. Artemis spouts nonsense about Secret Sacred Wars. Claremont, who has lost control of Artemis, tries to yell a warning to Cerebus, but is pushed back by Artemis. Artemis introduces Cerebus to his Secret Sacred Wars apprentices, Drew and Fleagle McGrew, who live in fear of Artemis. Artemis, carried away while delivering a lecture on how conflict builds character, runs amok, and Cerebus flings him into a packing case of whiskey, the contents of which Cerebus and the McGrews proceed to consume. Cerebus is suddenly visited by the glowing white sphere, which envelops him and disappears with him.

Talking Heads. Cerebus finds himself falling in starry space, and lands on a 82 flagstone floor floating therein. He is greeted by the floating, disembodied heads of Henrot and the unnamed magician who first appeared in Magicking. Meanwhile, back in Iest, Thrunk starts issuing papal demands. Astoria and her ex-husband, Lord Julius, observing events from a window, remind each other why they divorced. Back at the floating floor, the three-headed creature called “Fred, Ethel, and the little fellow with the hair” makes an appearance, buried up to its waist in the wall. Mention is made of the Final Ascension. The three-headed creature claims to be the Trinity of Tarim. The creature rises out of the floor and walks away. Henrot apologizes for not making it to Cerebus’s wedding, and Cerebus (still with a stuffy nose from his cold, which first afflicted him in Another Thing Coming) tells Henrot that Sophia left. The glowing sphere returns, frightening Henrot. The sphere envelops Cerebus and returns him to the storeroom from which he was taken.

Touch Not the Priestess. Cerebus is greeted by Michelle, who offers to feed 83 him. As she cleans up and prepares dinner, she conducts a running monologue, to which Cerebus listens quietly with his accustomed glowering expression.

For some reason you all apparently want to be Tarim. Not a pope or a prophet or even a messiah...you want to be the creator of all things...Well you can’t be can you? I mean what would you do? Create the heavens and the earth again? You don’t have to...they’re already here. It killed Weisshaupt. \underline{Killed} him...and for what?

Then Michelle tells Cerebus a little bit about herself and Weisshaupt. She was once the only female to attend Weisshaupt’s university, and Weisshaupt became her mentor when she graduated. Weisshaupt unsuccessfully tried to interest her in economics. When Lord Julius apparently out of the blue gave Michelle a huge endowment, Weisshaupt became paranoid, accusing Michelle of conspiring with Julius and Astoria, but Michelle’s guilelessness eventually convinced Weisshaupt to trust her. Finally, Michelle tells Cerebus that everything she told Cerebus when he stayed with her was a lie, and that she wants him to know what really happened.

84 Chariot of the Queen, Chariot of the Lovers. Michelle feeds dinner to Artemis and “the boys.” After dinner, Michelle tells Cerebus that she came to Iest to help Weisshaupt put his last affairs in order. She says that Weisshaupt asked her to take care of Artemis; that she misses Weisshaupt; that she neither wants nor ever has wanted anything to do with his schemes, but that Weisshaupt had given her something to give to Cerebus. She pulls a sealed envelope out of a drawer, explaining that the envelope will enable Cerebus to regain the papacy. Before she gives Cerebus the envelope from Weisshaupt, she demands that Cerebus understand that she won’t fulfill Weisshaupt’s request to stand with Cerebus at “the final watchamacallit.” Michelle tells Cerebus that Weisshaupt wanted someone “mundane” like Cerebus to be declared Tarim On Earth, so that Weisshaupt would be remembered for his political reforms. Michelle hands Cerebus the envelope and asks him to go. Cerebus opens the envelope to find a drawing of a cannon. Michelle, looking out the window, is amazed to see that it has started snowing in the middle of summer.

Missing It for the World. Cerebus walks through the snow on the road that 85 spirals around the black Tower toward the upper city. On the way, Cerebus is halted by a meeting with two drugged-out princes (caricatures of Mick Jagger and Keith Richards of the Rolling Stones) out seeking their fortunes. Cerebus accepts a drink of whiskey, which turns out to be laced with codeine. Cerebus, intoxicated, temporarily becomes very affectionate and generous.

Flying Off the Handle at Oblique Angles. Elrod makes an appearance in 86 a six-legged roach costume. He had heard about Artemis’s Secret Sacred Wars and, having mistaken the rumor for some significant religious event, wanted to take part. Artemis approaches, walking straight up the cliff face, soliloquizing to himself that the events predicted by Weisshaupt have come to pass, and that he will recover Astoria once Epop, the Anti-Pope, is defeated. Artemis shanghais Elrod by slamming a sucker-gripped hand onto the top of Elrod’s costume. With Elrod in tow, Artemis continues his climb. Cerebus mostly snaps out of his altered state of consciousness at an agonized shout from Claremont. Afraid that Artemis will reach the papal hoard of gold, Cerebus pursues Artemis up the cliff face, which has started growing upward, forming a hollow Tower around the upper city. Artemis, clutching an acrophobic Elrod, with Cerebus not far behind, muses that he plans to gain all the gold, become Tarim On Earth, and assume the power that escaped Weisshaupt. Claremont, rising out of the top of Artemis’s head, tries to yell warnings to Cerebus to stop Artemis, but is shoved back into Artemis’s head. Artemis, who has almost completely turned the tables on Claremont, is somehow deriving power from Claremont.

Towers Analogous. Artemis’s mutterings parody Batman’s soliloquies in Bat- 87 man: The Dark Knight Returns, frightening Elrod. Astoria waits above in the snow for Artemis to complete his climb. As Cerebus tries to catch up with Artemis, Cerebus hallucinates that the tower is falling over, but Cerebus nevertheless continues to climb. The growth of the Tower is not going unnoticed in the upper city; a stone demon’s head in the new Tower face crashes through a church window near Bishop Powers. When Artemis finally arrives at the top, he is so intent on confronting Thrunk that he pushes Astoria out of the way.

Out With the In Crowd. Astoria is helpless to prevent Artemis from attacking 88 Thrunk. Thrunk kicks Artemis, still clutching Elrod, over the rooftops and down to the lower city. Just as Astoria decides to try to reach the gold herself, Cerebus completes his climb. As Cerebus runs through the streets, Astoria yells at Cerebus to stop. Cerebus, without turning around, tells her to go away, and Astoria, to her horror, finds herself dissolving away in sparkles of light. Cerebus does not notice. He reaches the cannons, finding a box of matches on which is inscribed “Good Luck. W.” Just before Thrunk can smash him, Cerebus ignites one of the cannons and blows Thrunk to smithereens. The always-present crowd welcomes back Cerebus as Pope. Cerebus returns to his hotel and demands that Boobah fetch him food.

89 Odd Transformations No. 3: Dead Friends. Cerebus, asleep, dreams of a confrontation between Bran and Weisshaupt. Weisshaupt turns into a mansized Thrunk; Bran and Thrunk fight, leaving Cerebus alone. Elf appears and gives Cerebus a glowing gold coin that grows into a translucent glowing sphere of gold. Cerebus awakens, and goes downstairs to find the hemisphere of gold that had formed in his hand back in Time Passes and Life Manoeuvres in Volume 1.

90 Five. Boobah creeps out into the hallway for a midnight snack, only to find Most Holy frozen in place, his head enveloped by the glowing white sphere. In the dimness of the hallway, Lord Julius, Duke Leonardi, and Boobah put up a curtain in front of Most Holy and Boobah holds a sign saying “See the Living Tarim” with a list of admission prices. Sophia enters and tries to talk to Most Holy, who is unable to perceive her while Most Holy is in the grip of the sphere. Sophia yells at Most Holy for having been seen with Jaka, slaps him, and cries that she is trying to save their marriage. When Most Holy doesn’t respond, Sophia leaves, bawling. Bishop Powers enters, carrying a heavy statuette of an albatross, and tells those present that the Lion of Serrea has been assassinated. Outside, the Tower continues to grow.

91 Talking to Tarim.^{13} Oblivious to the external events in the preceding issue, Cerebus and the glowing white sphere converse back on the floating floor in space. The sphere talks about the Final Ascension. Whoever makes the Ascension must carry a perfect sphere of purest gold to the top of the Tower. The glowing white sphere says that in seven previous incarnations, it had itself tried to make the Ascension of the Tower, but each time, the Tower had fallen over. The object of the Ascension is to gain entry to Heaven, to talk to Tarim. Cerebus and the sphere speculate on the nature of Heaven. Cerebus decides that Cerebus will make the Ascension, now that the Tower is growing again. The glowing white sphere delivers Cerebus back to the hotel, leaving Cerebus with a perfect gold sphere. Cerebus looks up to see Julius, Powers, and Boobah.

13Issue 91 also contains a Jam story; see section 4.4.3 on page 12.

Audacious Tenacity, Tenacious Audacity. On bended knee, Powers tells Most 92 Holy about the assassination of the Lion of Serrea, which has left Most Holy as the sole and absolute Pontiff of the united Churches of the Living Tarim. Most Holy demands that every able-bodied man over the age of eight is to be drafted into a weaponless army of conquest. Just then, Cerebus remembers the Ascension and the sphere of gold in his hands, and starts to run out the door. He is halted by an artist and his agent^{14} who offer to immortalize Most Holy in a painting. As Most Holy poses for the painting, the gold sphere, resting on the floor, dissolves into gold coins. The artists leave, and Most Holy discovers what happened to his sphere. Powers tells Most Holy that if judgement is not rendered quickly on the imprisoned assassin of the Lion of Serrea, that the Lion’s legions will attack Iest. The prisoner is Astoria.

The Prisoner. Most Holy immediately journeys to the prison where Astoria 93 is being held. The carriage bearing Most Holy knocks Artemis, who has managed to climb back up the tower, back over the edge. In the dungeon, Most Holy and Astoria are left alone to talk.

So. Astoria, in chains before Most Holy, admits killing the Lion of Serrea. 94 Astoria and Most Holy argue, with Astoria trying to convince Most Holy that he owes her for helping to put Most Holy where he is. Most Holy denies that he owes Astoria anything. Finally, Astoria proposes an exchange of favors. Pulling down her undergarments, she offers to help the Most Holy aardvark with one of his “acquired tastes.” Kicking her underpants onto Most Holy’s head, she points out that by Church law, Most Holy cannot take advantage of her offer outside marriage. Most Holy thinks for a moment, then gags and blindfolds Astoria. Quickly declaring Cerebus and Sophia divorced and

14The artist and agent are caricatures of Bill Marks and Seth, affiliated with Vortex Comics.

Cerebus and Astoria married, Cerebus rapes her.

95 Odd Transformation 4. Asleep between Astoria’s legs, Cerebus dreams that he is being crushed and drowned beneath the weight of giant iron chains. He is awakened by the guard, who came to investigate when things got very quiet. Cerebus sends the guard away, and sits away from Astoria, where he declares himself and Astoria to be divorced, and falls asleep again. Cerebus dreams that he has thrown Astoria over the edge of a high building and is watching her fall, while behind him, the mouth of a cannon looms.

96 An Anchor That’s Going Places. Cerebus is awakened by Bishop Powers, who tells Most Holy that there are technical problems in recreating the perfect gold sphere. Powers hands Most Holy a scroll written in a language that Most Holy only pretends that he can read. Powers leaves, saying “’til midnight, then.” Most Holy removes Astoria’s gag and blindfold. Astoria, recognizing the significance of the gold sphere, angrily tells Most Holy that the Ascension is a sacred trust between Terim and her adherents, not for male Tarimites to attempt. Astoria questions Cerebus about his reasons for wanting to make the Ascension, and warns him not to try it. Astoria offers to read the scroll to Cerebus. It is a demand for Astoria’s trial to start at midnight, and for her execution by fire. Astoria and Cerebus then fall into a bitter religious argument.

97 The Unknown Given. Astoria tells Cerebus that Cerebus is caught between a rock and a hard place: if he, not Cirin, assumes the privilege of executing Astoria, Cirin will destroy Iest. If he keeps Astoria imprisoned, the Churches will fear that she is running things behind the scenes. If he lets her go, it will be perceived that he is “soft on Kevillism.” Astoria, whispering in Cerebus’s ear, tells Cerebus that she should be allowed to escape, then spits in his ear as a “little wedding present.” Nevertheless, Cerebus tries to pull Astoria’s chains free, but Bishop Powers returns. Powers tells Cerebus that the new gold sphere also has a flaw in it and must be recast. Then it is too late for Cerebus to free Astoria; he is forced to leave with Powers to prepare for Astoria’s trial.

98 The Trial. With full ceremony, Most Holy begins Astoria’s trial. A Cirinist spy, observing this, releases a carrier pigeon. Astoria gives her view of events. She claims that when Cerebus told her to go away, she disappeared to find herself in the presence of the Lion of Serrea, and was somehow compelled, without knowing why, to kill him. The Lion of Serrea had simply sat there and let her stab him. Powers calls Astoria a harlot, a slut, and a sorceress who must be destroyed before she destroys the Church. Astoria, weeping angrily, replies that she doesn’t need to destroy the Church; that the Church will destroy itself. Most Holy angrily asks Astoria if she sees herself as a prophet, and demands, “What is truth?”

Accurately Inexplicable. Astoria offers an aphorism that Powers calls blas- 99 phemous. Powers demands that Most Holy pass sentence. Astoria sarcastically agrees, then seems to become very calm, asking Cerebus to listen to faint echos; she says that she and Cerebus are the echos. Suddenly, Cerebus and Astoria seem to swap aspects: Astoria becomes a tall, male aardvark, its hands tied with rope; Cerebus is surprised to find himself in a woman’s body, dressed in the robes of a priestess. The tall aardvark invites Cerebus to “remember how it ends...how it always ends,” and begins to describe the process of burning a heretic. The priestess, not wishing to listen, yells for the guard to “Kill him!” which breaks the spell. Cerebus is again himself, a short aardvark in papal garb, and he falls to his knees to be sick. Powers quietly points out that to delay sentence any longer would be dangerous.

Endgame. Astoria continues her quiet description of the burning until Cere- 100 bus has her gagged. Cerebus washes his hands in a bowl of water brought by Posey and tells Powers that if Powers wants her dead then Powers must sentence her. Powers declines, directing Posey to explain. Posey timidly tells Cerebus that the killer of a Pontiff of the Western Church becomes the new Pontiff; Astoria is now technically Pontiff of the Western Church, and for Powers to sentence Astoria to death in turn would make Powers the new Pontiff of the Western Church. Powers points out that Cerebus need only pass sentence, and Cerebus will become undisputed ruler of the united Churches. Meanwhile, Cirin, notified by carrier pigeon of Astoria’s trial, musters all her mercenary forces against Iest. There is only one panel showing Cirin; underneath veils and bandages, she appears to be an aardvark.

Ave Avid. Events would seem to have a forgone conclusion, but Cerebus dis- 101 covers a small wooden chest that has mysteriously appeared before him without anyone noticing. Astoria, her appearance vacillating between aardvark and human, thinks “Something’s...different.” Opening the chest, Cerebus discovers a perfect gold sphere and a note saying “Good Luck. W.” Cerebus quickly sheds his robes, and carrying the sphere, runs. He crashes through a window. Astoria manages to pull down her gag and futilely yells, “Stop him!”

102 The Final Ascension. As the glass fragments of the window explode outward with Cerebus, and Cerebus reaches for the nearby inner surface of the still growing Tower, the reader is given glimpses of what other characters are doing and how they are affected by this extended moment: the Abbess views a projected image of the Tower from afar; Astoria struggles with her chains and shouts; the corpse of Bran stirs slightly where it fell; seated by a campfire in a forest, Bear mysteriously finds a glass splinter in his hand; Jaka steadily watches the Tower grow from a window in the lower city; in the upper city, Harmon Blakely stands stricken in the street. Just as Cerebus grasps the inner edge of the Tower, the Tower launches itself upward, separating from the rest of the wall surrounding the upper city, rising into the sky faster and faster, rotating.

103 Mass With Substance. The Tower’s rotation accelerates, making it difficult for Cerebus to keep his grip, but Cerebus eventually finds that he able to stand on the inner surface of the cylinder. Frightened when he looks down at the receding ground, he runs, and almost loses his gold sphere. Recovering, he proceeds toward the top of the Tower. Among the interlocking stone demon heads and skulls, he walks past a stone head resembling Weisshaupt. But as he tries to pass the head, more stone Weisshaupt heads fall into his path.

104 Earnest Nonsense. This episode is a cross-over with Bob Burden’s Flaming Carrot. A strange biped creature with a long carrot for a head and a flame rising from its top leaps to Cerebus’s rescue. It offers to be Cerebus’s guide. The Flaming Carrot takes Cerebus up, answering Cerebus’s questions almost, but not quite, entirely nonsensically. The two then encounter a Gumby-like creature named Sponge Boy. Suddenly the Tower stops rotating altogether, and gravity on the inner surface is lost. The Flaming Carrot and Sponge Boy head back down, leaving Cerebus clinging to the wall a few feet from the top of the Tower. The moon looms hugely overhead in black space.

105 Couple Capable. Finally reaching the top, Cerebus encounters Fred, Ethel, and the little fellow with the hair, as well as the return of some gravity. The three-headed creature is also carrying a gold sphere. The creature, displeased at finding competition for the Ascension, kicks Cerebus’s gold sphere out of Cerebus’s grasp, and the sphere is lost. Cerebus runs around the circumference of the rapidly narrowing top of the Tower, but is unable to escape capture.

Greatness Among Flies. Grasping Cerebus tightly in its fist, the little fellow 106 with the hair tries to convince Cerebus that Cerebus is so unworthy of the Ascension that he should leap into the void, and sets Cerebus down to give him the chance to do so. Instead, Cerebus climbs part of the way back down inside the Tower, out of reach of the creature. The end of the Ascension approaches, and the creature thinks that it has nevertheless “won,” but then the top of the Tower, having narrowed to the width of the creature’s feet, simply breaks off, hurling the creature into the void instead. Cerebus climbs up to the new top of the Tower as the Tower approaches the Moon.

Walking on the Moon. This chapter bears the note “with apologies to Jules 107 Feiffer and Lou Jacobi.” Falling to the moon, but not quite touching the surface, Cerebus is greeted by a middle-aged humanoid dressed in the robes of a judge, who also hovers a few inches above the surface. The Judge says that he is a little disappointed, because he was pulling for Weisshaupt to make the Ascension. The Judge anticipates all of Cerebus’s questions, so Cerebus says nothing. First, the Judge says, he is not Tarim, he is simply there to observe and to judge. He sees everything happening on Earth, hears everything, and remembers everything. The Judge begins to recite a summary of the Earth’s entire history, past, present, and future, but makes a digression:

In the interest of maintaining your limited span of attention, I’ll answer your questions now...afterward...I’ll return to the meaningful issues at hand. First...the question you had in mind when you began your Final Ascension...you don’t conquer the known world when you get back...so the question of ‘how’ becomes immaterial...you came very close, but that’s as close as you come.

Extrusion, Intrusion. The Judge continues with a history of Estarcion. The 108 Judge walks a little too far ahead of Cerebus, and is temporarily out of earshot. Cerebus discovers that he can shout, but cannot make a sound. When Cerebus catches up, the Judge, who hasn’t noticed Cerebus’s absence, is at the end of a story about a happening in Jerusalem (the first mention of a familiar city anywhere in Cerebus). The Judge mentions the Big Bang as it will be theorized in six thousand years, and opines that the theory is a non-sequitur. “In the beginning,” he says, “there was the void, and the void was Tarim.”

109 Abhorring Vacuums. The Judge tells Cerebus the story of the creation of the universe by Tarim’s investigation/rape of a bright light, Terim, appearing in the void.

110 All the Suns are Daughters. The stars resulting from the rape and white explosion of Terim are all that remain of Terim. The Judge analyzes the ironic aspects of creation:

And what do you suppose was Tarim’s reaction? —As the white light began to fade and the enormity of the wrong he had done became evident. Regret? Astonishment verging on catatonia? Overwhelming sorrow? Not on your sweet life. Anger! Anger was Tarim’s first reaction.

111 Base Tranquillity. The Judge disgustedly describes the future exploration of the moon six thousand years hence, soon thereafter followed by the destruction of all life in the solar system by a man-made doomsday weapon. The Judge, hearing Cerebus’s unspoken question, advises Cerebus that all Cerebus can do is to accept suffering and be redeemed by it. He tells Cerebus what has transpired on Earth in Cerebus’s absence: Tarim did not bring about the end of the world as Most Holy had predicted, and all Cerebus’s followers have deserted him; Cirin has conquered and sacked Iest, seizing all of Most Holy’s gold. As the glowing white sphere appears and begins to envelop Cerebus to take him back to Earth, the Judge tells Cerebus that Cerebus will live only a few more years, and die alone, unmourned, and unloved; and that if he ever considers his suffering unjustified, then he is to remember his “second marriage.” The sphere takes Cerebus away, depositing him back in the deserted ruins of his hotel on the slope of Greater Iest.

9 Square One

The story picks up again in a “double issue” with Cerebus alone next to his 112, 113 hotel, the street partially destroyed by a rockslide of demon heads and skulls from the cliff face. Square One is an extended coda to Church & State. There is no dialogue until the very end.

Some small print in issue #111 states that Square One, like Exodus, will not be reprinted.

Cerebus wanders through the ruins in silence and dejection. Entering his hotel, Cerebus finds broken windows and a farewell note from Boobah. Going upstairs, Cerebus recalls Sophia packing to leave. He finds one of her platemail bikini bottoms in a drawer, and appears to think about how much he misses her. Then he unsuccessfully tries to rend the article, and leaves it on the floor. He goes to find his old jacket, necklace, and sword, in the same room where Bran’s corpse still lies. Cerebus dons his gear, keeping a wary eye on the decaying and fly-ridden corpse as though he fears it might return to life. He cuts his thumb slightly when he reaches for his sword without looking. Dressed and armed, he turns away for a moment, then is startled anew by the corpse. He retreats from the room. Out in the courtyard, he remembers Boobah waiting to transcribe his every word.

Walking down the steps leading from the hotel, he finds that the steps end in an abyss where a rockslide demolished that part of the hillside. He sits down and dejectedly recalls the Judge’s prophecy that he would die alone, unmourned, and unloved. He tosses his sword backward over his shoulder, removes his gear, and tries to commit suicide by leaping from the edge, but finds himself unable to make the leap. Covering his eyes, he tries to make himself fall, but catches himself. He pulls himself back up and falls asleep in the rain. Cerebus dreams that he is walking alone in a blizzard, then dreams that he is on a suspended rope bridge that unravels underneath him. He awakens with a stiff neck. Picking up his sword, jacket, and necklace, he returns to the hotel and falls asleep face down on a rolled-up carpet.

Awakening near sunset, Cerebus leaves the hotel by another route, having found a single gold coin—all that remains of his former papal fortune. He walks into a tavern farther down the spiral road leading from Greater Iest. The only persons in the tavern are a middle-aged man and the hulking barkeep. The customer quietly rants about the Cirinists, hawking and spitting occasionally in the general direction of the barkeep’s shirt. “No women allowed in taverns. Citizenship for mothers only. No music. No dancing,” he laments. When the customer rhetorically asks where it all ends, Cerebus, standing in the doorway, says “Someone pushes a button and blows up the Sun.” He turns and leaves.

10 Jaka’s Story

Jaka’s Story kicks off a new Cerebus novel. The components of Jaka’s Story are named Prologue (issue #114), Pogrom’s Progress (issues #115– #119), The Poet (issues #120–#130), and Mystery Achievement (issues #131–#136). This section, which covers only Prologue and Pogrom’s Progress, concludes the Cerebus story summaries for this edition of the Companion. Sim writes that the last issue of Jaka’s Story, #136, is scheduled for July, 1990.

Jaka’s Story cross-cuts between a prose story of Jaka’s childhood and the tale of a visit by Cerebus, the ex-Pope, to Jaka’s current surroundings. The story concentrates primarily on subtle interactions between the characters and on a lopsided love triangle. Jaka now lives with her husband in one of the small, now mostly empty, Iestan towerside communities.

10.1 Cast of Jaka’s Story

In order of appearance:

Nurse, Jaka’s childhood nurse and guardian. Nurse is a large woman, barren, full of worries, conservative attitudes, and interesting northern ideas about health. Nurse’s face is not seen, except faintly in deep shadow; often, a postage-stamp panel insertion of Jaka’s rag doll and only friend, Missy, is shown. Nurse appears only in the flashbacks to Jaka’s childhood.

Jaka, quiet as a child, raised almost entirely by Nurse in Lord Julius’s household in Palnu. She is full of imagination, in spite of her dull tutelage; she seems to endure her childhood rather than live it. As an adult, she has a stern work ethic and little patience with foolishness, particularly from her husband. It is revealed that she was (and may still be) one of Lord Julius’s principal heirs.

Rick, Jaka’s husband, is a mild fellow, sort of lazy. He is skinny, blond, and wears a thin, short, goatee. He loves Jaka with a desperate sort of yearning. He isn’t wild about having to chase work, but does what he must to keep Jaka happy with him.

Cerebus, understandably still in the grips of depression, follows the road down from Greater Iest, warily looking for people.

Pud Withers, Jaka’s quiet and unfailingly polite employer, a paunchy, middle-aged fellow with close-cropped hair just at the top of his head and a thin moustache. He owns a modest tavern in which customers fail to appear night after night. He is a widower; in spite of his outward reserve, he is terribly lonely, and entertains a few little fantasies about Jaka. He and Jaka have a formal conversational ritual every day when Jaka comes to Withers’s grocery; a less formal one when Jaka comes to work at night.

10.2 Summary of Jaka’s Story

10.2.1 In the Present

Jaka awakens and tries to encourage Rick, her husband, to get up, get out, 114 and look for work. A little bit of tension develops over Rick’s listless efforts to do all that and his unenthusiastic response to Jaka’s breakfast porridge.

Cerebus enters Withers’s tavern and demands ale. Cerebus pays with his one 115 96 10 JAKA’S STORY

Rick

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

Pud Withers

\begin{matrix}c \\ ◯ \end{matrix} 1989 Dave Sim

gold coin, which causes an unexpectedly fearful reaction from Mr. Withers. The door bells tinkle in the adjoining grocery and Withers, still in a bad state, goes to see his customer, muffing his usual pleasantries. Cerebus, hearing Jaka’s voice coming from the grocery, goes to investigate. When Jaka sees him, she goes to embrace him, rejoicing that Cerebus is still alive. Jaka invites Cerebus over to her house for tea. At Jaka’s, Cerebus nearly bolts when the subject of Jaka’s husband, Rick, comes up. Jaka tells Cerebus briefly about the madness in Iest when Cirin’s female army arrived; Jaka moved to the mountainside to escape it. Cerebus asks what happened to Jaka’s baby. Jaka hesitates, then tells him that she had a miscarriage.

Cerebus insists that he has to go, but Jaka asks him to stay to meet Rick. 116 When Rick arrives, Cerebus seems surprised at his friendly greeting. Jaka tells Rick that she has asked Cerebus to stay for a few days. Rick guides Cerebus to an empty baby’s room next to the master bedroom. Cerebus tells Rick that he is in love with Jaka. He braces for a confrontation, but Rick does not seem to take the statement seriously. Later, Rick gets amorous with Jaka, but Jaka puts him off to get ready for work.

As Jaka waits at Withers’s tavern to begin dancing, Rick, waiting at home for 117 her to return, plays a little game with a ball and bucket. Cerebus, lying awake in bed, is annoyed by the bumping noise of the ball. Finally, Rick retires and starts snoring loudly. Jaka takes her night’s wages of three copper bits and leaves the tavern, no customers having arrived, no performance given. When Jaka returns home and checks on Cerebus, Cerebus feigns sleep. Rick reawakens and suggests to Jaka that Cerebus’s name, being the same as the ex-Pope’s, is not a safe one to have with Cirinists about. Cerebus lies awake, listening to the conversation. Jaka again rejects Rick’s amorous advances, much to Cerebus’s relief.

Cerebus sleeps for two days, then is gently awakened by Jaka. Cerebus offers 118 some serious apologies to Jaka: that he remembered her, and didn’t come back; that he treated her so badly when she came to return his sword; that he had hit Jaka when she tried to save him from an invading army; that he tried to get Jaka to leave with him when she was pregnant with another man’s baby. Jaka listens quietly, seated on the bed. Then Cerebus tells Jaka that he isn’t leaving without her. Jaka responds by giving Cerebus back his gold coin; Mr. Withers could not handle the legal ramifications of such a huge amount of money. She tells Cerebus that he is welcome to stay, but that she loves her husband—not Cerebus. She explains that the city is not safe for Cerebus, and that although she has lived most of her life without fear, she does so no longer:

I don’t know if you understand what I’m saying... You don’t frighten me, but it’s because of you that I’m always afraid... Having said all that, I’m hoping you’ll stay... I’ll feel better knowing you’re somewhere safe.

Rick, eavesdropping, becomes ill with the realization that Cerebus is the ex-Pope.

119 Marital tensions rise in Jaka and Rick’s household; Rick briefly suffers from superstitious awe of Cerebus, and Jaka remonstrates him for asking Cerebus if he knows how to make a pillar of fire. As the three enjoy a “night out” at Withers’s tavern, Rick, much to Cerebus’s surprise, gets tipsy after only one ale, then suggests that he and “Fred” the aardvark return home. At night Cerebus listens to the couple’s bedtime conversation, savoring the possibility that they might break up. Rick and Jaka soon make up, however, and Cerebus weeps listening to the sound of lovemaking.

10.2.2 Jaka’s Childhood

114 The adult Jaka’s morning routine is cross-cut with flashbacks to her childhood, when Nurse orchestrated the morning awakening, bathing, dressing, and breakfast. The young Jaka dotes on Missy, her rag doll, given to her by one of the household staff when Jaka’s mother died.

115 Jaka’s imaginary play world at her playground begins to fail her. Instead, 116 she spins around and around until it returns to her; a tiny glimpse of her 117 future appears; she yells “Jaka!” Nurse suddenly looms over Jaka to see what

has possessed her charge. Jaka gets an urge to flee, and, dizzy from her spinning, instead falls and strikes her head against the base of the playground’s

118 wooden horse. Young Jaka is unconscious for five days and suffers fever and hallucinations for days after that, with only Missy to protect her from huge nightmare insects.

Young Jaka recovers from the blow to her head, but exhibits a certain dull- 119 ness. The remainder of her childhood indeed becomes an exercise in dullness when she becomes old enough to receive parochial instruction. Jaka cultivates dullness as a sort of defense against her guardians, which makes Nurse think of her as somewhat retarded. At age ten, Nurse gives Jaka warnings of the evils, humiliations, and tortures of human sexuality.

This concludes the story summaries. Jaka’s Story continues in the next book, “The Poet,” issue #120.

11 Further Reading

Various comics have been mentioned throughout the Companion; these must be obtained by searches for back issues at comic book stores and other sources. Graphic novels, such as the collected Batman: The Dark Knight Returns, may be found at comic book stores, although some general book stores carry them as well. Examples of swords-and-sorcery and fantasy books by Robert E. Howard (Conan), Fritz Leiber (Fafhrd and the Grey Mouser), and Michael Moorcock (Elric) may be found in the Fantasy and Science Fiction sections of most general bookstores.

The Illuminatus! Trilogy, by Robert Shea and Robert Anton Wilson (prose; published by Dell), is a sort of anarchist’s primer disguised as a humorous paranoid fantasy for adults. Warning: it’s hard to follow. Cerebus’s Adam Weisshaupt is patterned after an 18th century Adam Weishaupt, of whom a bit is said in the Illuminatus! Trilogy. In the Illuminatus! Trilogy, there is much discussion of the role in history of secret societies and conspiracies, both real and imagined. Adam Weishaupt (1748–1830) was a Bavarian student of Jesuits, a professor of canon law at the University of Ingolstadt, a Freemason, and founder of the Order of the Illuminati.^{15} It is alleged that the 18th century Weishaupt usurped the identity of George Washington. Allusions to this allegation, and to other aspects of the Illuminatus! Trilogy are made in Cerebus; for example, Cerebus wrote of the revolutionary conspiracy by “The Eye in the Pyramid,” speculating that the conspiracy was actually engineered by Lord Julius. The name of the revolutionary group is taken from the first book in Illuminatus! Trilogy.

15This much about Weishaupt is from New England and the Bavarian Illuminati by Vernon Stauffer, published in 1918 and reissued in 1967 by Russell & Russel division of Atheneum House, Inc.

The Historical Illuminatus Chronicles, volumes 1 and 2 (The Earth Will Shake and The Widow’s Son), by Robert Anton Wilson (prose; older editions published by Bluejay Books; newer editions by Lynx Books), expand on the themes in the Illuminatus! Trilogy using slightly more conventional storytelling techniques. The stories are nominally historical dramas, following the exploits of an 18th century Neapolitan lad named Sigismundo Celine, who is inducted into and victimized by various European secret societies.

12 Aardvark-Vanaheim Ordering Information

These are the prices of the Cerebus phone books available at the time of this writing. Prices, of course, are subject to change, so should be confirmed by looking at current issues of Cerebus or by contacting the publisher.

Title  Price (Canadian)
Cerebus  $27.00 + $2.50 P&H
High Society  $25.00 + $2.50
Church & State volume 1 $30.00 + $2.50
Church & State volume 2 $30.00 + $2.50

Contact information:

Aardvark-Vanaheim, Inc. Box 1674, Station ‘C’ Kitchener, Ontario N2G 4R2 Canada phone: (519) 576-7820, or (519) 576-0610; 9:00 to 18:00 EST A-V accepts Visa, MasterCard, certified checks, and money orders. US prices are converted to equivalent funds, and are a few dollars lower at the time of this writing.

The author’s experience with ordering has been that credit card phone orders are sometimes shipped right away, other times after a few days, via regular postal service. Mail delivery takes between twelve and twenty-seven calendar days across North America.

A few comic book shops have been known to carry the phone books. In the past, A-V’s quantity/pricing policies have not been to the liking of many shops, but this is starting to change, so the phone books may start showing up in shops more often.

13 Acknowledgements

Antonio Romero, soon to be reknowned Melville expert, spent a great amount of time poring over a draft of the Companion, exchanging exegetical electronic mail messages, and making helpful suggestions. His efforts are much appreciated.

Barry Short and employees Catherine and Jeff of 21st Century Comics and Toys in Orange, California, were very helpful in tracking down Cerebus material. Mr. Short provided many specifics on the parodies appearing in Cerebus, from characters and authors, right down to issue numbers and page numbers of Conan and Moon Knight.

The author’s first idea to investigate the Cerebus series came from messages appearing in a public forum in the USENET computer network, the “net news” group rec.arts.comics, also known to its readers as rac. Rac supports a multi-threaded, up-to-the-minute, dynamic, ongoing discussion of anything and everything related to comics. Cerebus appeals to a number of rac’s frequent contributors, and is the subject of much discussion and speculation. The Companion owes something to these discussions. In particular, thanks go to these persons for filling in some bibliographical details and for providing interesting discussions in the first half of 1989:

   Jerry Boyajian, boyajian@ruby.dec.com
   Roger Crew, crew@polya.stanford.edu

Dan’l DanehyOakes, djo@pacbell.com

Justin du Coeur, inmet!justin

   Tom Galloway, tyg@caen.engin.umich.edu
   D. W. James, vnend@phoenix.princeton.edu
   Ron Rejmaniak, golgo@buengc.bu.edu
   Antonio Romero, romero@confidence.princeton.edu

The author’s e-mail addresses are:

   dolphin@lurker.irvine.ca.us (Internet)
   ucivax!lurker!dolphin  (UUCP)

355-1437 (MCI Mail)

Comments and queries about the Cerebus Companion are welcome.

Thanks also go to Harry Kremer of Now & Then Books in Kitchener, Ontario (a former employer of Dave Sim), for confirming some bibliographic data. Richard and Lisa of Comic Quest in El Toro, California, provided some additional information and encouragement.

Some bibliographical details of swords-and-sorcery stories and comics were confirmed by The International Book of Comics by Denis Gifford, Exploring Fantasy Worlds, edited by Darrell Schweitzer, The Dark Barbarian, edited by Don Herron, and The Great Comic Book Artists by Ron Goulart.

Jerry Sweet Irvine, California April–July, 1989

14 The Cerebus Glossary

104 GLOSSARY

aardvark—Cerebus’s species Aardvarkian Age—notation on a map of Estarcion Agerden, Duke of—an old acquaintance of Cerebus living in Togith Akshun Crossing—a Palnan river crossing Akshun, Duke of—a legendary hero Alanne—a milk maid whose pregnancy is attributed to Necross the Mad Albatross, the—a token of power that can unite the Eastern and Western Tarimite Churches Albino, the—reference to Elrod, who is an albino ale—a fermented, malted beverage, i.e. beer; Cerebus’s favorite drink Alfred—Suenteus Po the Second, son of the conqueror Suenteus Po the First Alucase—an imaginary substance believed by the Cockroach to impart superhuman strength Ambassador Suite—Cerebus’s suite of rooms in the Regency Hotel Amstanat—region in the southeast Anarcho-Romanticism—a New Republican splinter party formed by the self-styled intelligentsia of Iest Andros—a legendary hero-king Apocalypse Beast—any of several legendary destructive creatures, including Fred and Ethel apricot brandy—a liqueur of which Cerebus is fond Arnold the Isshurian—a mercenary cavalry commander and adding machine Artemis—one of the Cockroach’s civilian personalities; full name: Artemis Strong Ascension—see Final Ascension Ashtoth—a god by which oaths are sworn Astoria—Duchess of Parmoc, possibly the founder of Kevillism, an advisor to Cerebus AUFO—the Anti-United Feldwar Organization, a terrorist group secretly controlled by Weisshaupt Avers—a satellite city of Palnu Baldy—full name Baldy Pate; a fence, an old acquantance of Cerebus’s Bandit Moons—unknown entities that once kidnapped Sponge Boy Barmaid and the Watchdog, the—an obscene song Baskin—The Palnan Minister for Executive Planning, Lord Julius’s stenographer Bay of Sunshee—a port in Palnu Bear—an old acquaintance of Cerebus’s; later a mercenary guard in Cerebus’s employ Beasties, Ye Book of—a book consulted by magicians Beduin—a city in Lower Felda with many waterways Bennie—a Sergeant-Major in the Iestan military Big Garn—an Iestan fence Black Blossom Lotus—a magic talisman stolen by Cerebus Black Death—the usual plague, attributed to Necross the Mad Black Sun cult—a ruthless, death-worshipping cult Black Tower—the ominous tower on which Greater Iest is built Black Tower Empire, the—an empire toppled by the Pigts Blake—an informant of Astoria’s Blakely—an electoral district of Iest Blakely, Cyrus N.—one of Harmon Blakely’s ancestors Blakely, Harmon II—member of the Iestan legislature, later a member of Cerebus’s cabinet Blinky Boar and the Strawberry Patch—a children’s fable Blood Wars—an ongoing conflict among northern barbarian tribes Blue Duck, the—a tavern in Port Seprania Boobah—a mercenary and Cerebus’s secretary Bootsie—a prostitute in Maggie’s cathouse Boreala—a Northwestern region, home to marauders B’orn—a town in the Docks district of Iest Brahpur—a northwestern city Brandur—a city in Boreala Brugel Arms—a tavern in Togith in which Henrot has taken residence Bug—one of Cerebus’s names for the Cockroach Bunky—the Cockroach’s name for Elrod Bureaucratic Rebellion—a power grab by the Iestan Bureaucrats just after the Election of 1413 Buz—a nutritious, but narcotic substance used by K’Cor to feed and enslave the population of Imesh cannons—the weapons with which Weisshaupt guarantees the safety of the UFS against outside invasion Captain Arnold and His Armpit Avengers—a mercenary force Captain Cockroach—a persona of the Cockroach manufactured by Weisshaupt Carroll E. King Reads—a bookstore in Iest Carron—an arrogant Iestan aristocrat, partner in Carron, McKiel and Benny Cerebus—an aardvark adventurer Cernebus and Sophina—one of Weisshaupt’s planned romance novels Chaos Gem, the—a magical gem coveted by Death, thrown down a well by Cerebus Chateau Dehrsion ’26—a favorite vintage of the Estarcion aristrocracies Chateau Donte ’21—a favorite vintage of Weisshaupt Cheen—a general; member of the military tribunal ruling Lower Felda Chesmi River—a river flowing through Iest, joining with the Feld River Children’s Fables—stories collected by Suenteus Po Cihnu—a satellite city of Palnu Cinnamon Crisis—a shortage in Palnu of cinnamon from the Eshtannin Wilderness caused by Onliu naval predation Ci’nuffinn—a province of Palnu Circe—leader of the Cirinist military forces Cirin—name or honorary title of the leader of Upper Felda Cirinism—a female-supremacist philosophy, a conspiracy Claremont, Charles X.—an incorporeal being inhabiting the Cockroach’s head; formerly a mad sorcerer Clovis—one by whose beard, beads, blood, teeth, and other attributes oaths are sworn; the name taken by several Eastern popes Cock and Bull—a tavern in Beduin Cockroach, the—the base persona of a mad avenger Colin-Hampton—the late President of the Bank of Iest comne ye tama stet fegria—a Paranian insult Concordance Eve—a midwinter holiday Conniptin Tribes—militant northern tribes Cootie—one of Cerebus’s names for the Cockroach Corn-King rituals—an collection of ancient religious rituals Crown—a unit of currency Cutter, the—a Palnan trade ship, sunk by Onliu pirates, on which Cerebus and Lord Silverspoon took passage Dabord, Lucien—a Diamondback player defeated by Cerebus Darnier Prison—a prison in Lower Felda Dead Albino—a name taken by Elrod during his out of body experience Death—a malevolent supernatural being Deguin—a city guard in Beduin Dehrsion—a southwestern city on the Sofim River, known for wine making Deidre—a city guard in Beduin Denieau—a Lower Feldan port city; site of a temple partly financed by Leopold the Gambling Priest’s Diamondback winnings Derkass—the Anti-Sacred Wars Roach: Cerebus. (Say it backward for amusement.) Dernier, Gaston—a Diamondback player defeated by Cerebus Desert of Neshfalnin—eastern region near the Sea of Tanchya Deshen—Second General, member of the military tribunal ruling Lower Felda Despuess—a Lower Feldan farmer annihilated by Necross the Mad Detin, Michelle—a former student of Weisshaupt’s living in New Sepra as a widowed countess D’Gan—leader of Togith Diamondback—a card game of which there are many variants dinosaurs—a race of creatures that ruled Earth until they halved the oxygen content of the atmosphere with a hairspray Directory of Palnan Executives, Study Groups, and Department Heads—a much sought-after, possibly non-existent document for which exorbitant fees are charged Divine Mother—name given to Terim by the Cirinists D’Mitri—a citizen of Theyr crushed by Thrunk Docks, the—an electoral district of Iest, at the confluence of the Feld and Chesmi Rivers Dog and Duck—a tavern in Iest Dog and Raven—a tavern in Pedron in which Cerebus plays Diamondback Doran—a Magus with whom Cerebus once studied the sorcerous arts Draser—an Iestan aristocrat Drew—see McGrew, Dirty Drew Dualism as Fallacy—a Cirinist tract duck—the mistaken identity of the Albatross DuFort—see Madame DuFort dung worms—smelly beasts Dunham—a town in the Docks district of Iest earth pig—alternative name for an aardvark East Chapel—a town in Iest and Suburbs Eastern Monoliths—an empire toppled by the Pigts Eastern Orthodox Church of Tarim—the official name of the Eastern Tarimite Church Edmonds, Lord Francis—name on a package believed by Cerebus to have been sent as a trap by Elf Eighth Sphere—an altered state of consciousness E’lass—a sleazy character who introduced Cerebus to Jaka Election of 1413—the election in which Cerebus first becomes Prime Minister of Iest Elf—a friendly extranormal being in residence at the Regency Hotel Elmo—Lord Storm’Send’s cousin; the Northbell district’s deciding vote Elrod—a tall albino swordsman, wandering King of Melvinbone; aka “Bunky,” the Cockroach’s sidekick; later, Lord Julius’s adopted son and ranking diplomatic representative of Palnu in Iest Emperor God of the Shen—a title given to the sitting ruler of the Shen Energy Globes of Imesh—floating globes that fire energy bolts; Cerebus has run their gauntlet twice Eniath—a satellite city of Palnu Enothas—a southwestern city-state Éoaghan—region northwest of Boreala Epop—the Cockroach’s name for the anti-Pope Eric—one of Cerebus’s uncles Eshnosopur—a southern city-state known for its warm climate and assassinations by poison Eshtahnin Wilderness—the southeasternmost region of the continent Estarcion—the continent on which most of Cerebus takes place Ethel—the female head of the Trinity of Tarim Exodus Inward—a ritual cloistering of the upper echelons of the Eastern Tarimite Church, invoked in 1413 prior to the election and previously during Alfred’s invasion of the East Eye in the Pyramid—a revolutionary conspiracy in Palnu Eye of Terim—most precious of the five Spheres of the Gods, from which wizards may draw power Feld River—a river flowing through the Tcapmin Kingdoms and the Feldwar States Feldwar States—a federation of city-states whose capital is Iest Feras, Tanes—Red Sophia’s late first husband fermented goat’s blood—a favorite drink of the Pigts Festival of Virgins—a Lower Feldan tradition Filbert—the mysterious artist/collector’s Apocalypse Beast; later, Fred Filgate, Euling S.—representative of the Docks district in Iest; later a member of Cerebus’s cabinet Final Ascension, the—a mystical journey up the black tower to converse with Tarim, or perhaps to become Tarim Five-Day Rebellion—a relatively recent occurrence in Eshnosopur Flame Jewel—hot property of a wizard, target of Cerebus’s first adventure Flaming Carrot—a creature that comes to the aid of Cerebus during the Final Ascension Fleagle—see McGrew, Dirty Fleagle flour—a useful substance with which Cerebus starts tavern brawls Fluroc—a Palnan city defended by Krull Formation Act—an agreement forming the UFS that also defines the separation of powers between the Church and the government Fort Chesmi—an outpost in Garrison district Fort Chicolini—a New Sepran outpost Fort Cleese—an outpost in Parmoor district Fort Columbia—a Tcapmin outpost just north of the Wall of Tsi Fort Dead Parrot—an outpost in Parmoor district Fort Felicity—a New Sepran outpost Fort Frot—an outpost in Parmoor district Fort Garrison—an outpost in Garrison district Fort Gort—an outpost in Garrison district Fort Grace—an outpost in Grace district Fort Grit—an outpost in Parmoor district Fort Hartin—an outpost in Grace district Fort Hermes—an outpost in Garrison district Fort Iron—an outpost in Garrison district Fort Palin—an outpost in Parmoor district Fort Redemption—an outpost in Grace district Fort Torte—an outpost in Parmoor district Fort Tory—an outpost in Parmoor district Fort Trof—an outpost in Parmoor district Fort Waldemor—a New Sepran outpost Fred—(1) a Swamp-Thing-like creature, formerly the Apocalypse Beast named Filbert; the male head in the Trinity of Tarim; (2) an alias given by Rick to Cerebus for Cerebus’s protection Frog and Duck—a tavern in Palnu Frostonite—very cold fragments of the Moon Roach’s exploded home planet, Froston; in liquid form, whiskey Galfann—leader of an army that once assailed Fluroc Gan-Fonil—a southeastern city near Amstanat Garrison—an electoral district of Iest Garsk—once captain of the guards in Imesh Gatson—a corrupt former Prime Minister of Iest Geet-a—an out-of-control clone of Sophia used as a watchdog by Henrot Genn, Turan—a mercenary captain, former employer of Cerebus gerbies—wizards’ pets that feed on violent emotions Gerrik—Astoria’s first husband; thought to be Cirin’s adopted son Gesolde—a priest, emissary from the Sepran Empire G’evan—a thief in Togith who steals Cerebus from the Cirinists Ghetty’s Point—site of an army encampment commanded by Krull Giepie—a Palnan importer of goods Gilash—by whose third eye oaths are sworn Ginshu root—a non-existent mind-control drug Gishan, Jacin—a Diamondback player defeated by Cerebus Gisse—a prince who once achieved mastery over an Apocalypse Beast Gizzard Stomper—Fleagle McGrew’s war hammer Goat, Julius—Lord Julius’s goat, the candidate in opposition to Cerebus in the Election of 1413 Gods’ Fence, the—mountain range between Isshurian Lands and Boreala gold sphere—a talisman necessary for the Final Ascension Goodabbey—the Abbess’s demesne Gorce—a bureaucrat in Vessariat; political opponent of Lord Julius; commander of the forces that crushed the invading T’Gitans Gort—an Illusionist of great meditative skill living in Togith Grace—a town in the Docks district of Iest Graus—a T’gitan warrior, a friend of Stromm, and briefly one of Cerebus’s travelling companions Gray Frostonite—a substance of which, according to the Cockroach, Cerebus is made, and which can take away super-powers Great Fire of Beduin—another catastrophe attributed to Necross the Mad Great Flood, the—a flood occurring every twelve thousand years that covers the surface of the Earth; the result of fallout from the redwoods’s Sodium Chloride Bomb Greater Iest—the part of Iest built on the demonic tower; the governmental and high-rent district of Iest Greely, Ned—partner in Greely, Bleaker and Spleen Grimes—an assistant to Krull Griselda—a gerbie Gudre—leader of the T’gitans, an un-T’gitanly short, pudgy fellow who wears an eyepatch Gums—a village idiot in a Lower Feldan farming community Gurann—a city in Boreala reknowned for freak shows and kissing techniques Gwane—a friend of Lord Silverspoon’s Hadden, Holland M.—partner in Hadden, Hadden and Dipp; purveyor of gold-plated street lamps Hammel—a spice merchant in Fluroc, deceased Hampton, Scott and Bo—Tcapmin soldiers Hanek—one of Cerebus’s servants Harbourside—a town in Iest and Suburbs Hargraves—spokesman for the Anarcho-Romantic Steering Committee Harker—an aide to Lord Gorce Harmony—the name taken by four Eastern pontiffs Harne—borough of Lower Iest Harpengate—a town near Iest Hedges—Weisshaupt’s aide Henrot—a powerful wizard, father of Red Sophia Henrot-Gutch, Mrs.—Red Sophia’s mother Hent—a Corporal in the Iestan military Herces—a wizard, creator of the Black Blossom Lotus Heskins—a tailor Hezzreth—king of the Conniptins Hill-Hodges—captain of the Iestan Dock Police Hobbsgate Prison—a gaol in Iest Hortne—an alias used by Henrot Howe, M. T.—Director of the Bank of Iest Hsifan Khanate—large empire east of the Red Marches inhabited by a nomadic asiatiform ethnic group Hsiffy—derogatory term for Hsifan Hyanns—an Iestan military commander Ianag—a region neighboring Palnu Iest—a central city-state, capital of the United Feldwar States Iest and Suburbs—an electoral district of Iest Illusionism—a meditative philosophy, possibly a conspiracy as well Imesh—a Hsifan city where Cerebus once lived, now enslaved and ruled by K’Cor indenture law—an Iestan law, instituted after the Cirinist invasion, regulating ownership of gold Inquisition, the—enforcement arm of the Orthodox Tarimite Church Invasion of 1414—the “Hsifan” (actually Conniptin) invasion of Iest just prior to the end of Cerebus’s first administration Ishtar—one by whom Pigtian oaths are sworn Isshurian Lands—northernmost western region Jaka—a tavern dancer, Lord Julius’s niece Janette—one of Madame DuFort’s students Jockley, Blake—the Cockroach’s taxi driver persona Johns—an Iestan military commander Judge, the—a quasi-omniscient being who watches Earth from the Moon Julius—Lord of Palnu Jylcew—a satellite city of Palnu Kael, B’ren—an inspector; executive assistant to the Reeve of Harpengate Karet—Cerebus’s favorite variation of Diamondback Katrina—one of Madame DuFort’s students; a niece of Lord Julius’s K’Cor—king of Imesh Keef—a prince strangely resembling Keith Richards of the Rolling Stones Kevillism—a philosophy derived from Cirinism and Shen devotional beliefs; a conspiracy Kevillist Rebellion—an occurrence in Lower Felda ca 1408 Kevitch—one of the Cockroach’s personalities; possibly his real name Khaidoge the Ungovernable—map notation on an eastern region bounded by the Pimina Mountains and the Wall of Tsi Khaiven—group known for sorcerers; possibly associated with Khaidoge the Ungovernable Khem—the Dark Demon, with many faces and a crimson claw; presumably the object of worship by Red Claw cultists Kiel—borough of Lower Iest Kitchen Staff Supervisor—an important cabinet position in Palnu once held by Cerebus; the director of Julius’s security forces Klog—a Borealan marauder defeated by Cerebus in single combat Kóin Tribes—eastern tribes Korem—king of Panrovy Koshem—a Panrovian lord, K’Cor’s champion Kota Mountains—mountain range between the Sepran Empire and the southern city-states Krauftig—a T’gitan warrior Krull—a Palnan military commander, a very strong, brutal, capable fellow with a streak of egotism Lafort—Executive Commander of Lower Felda; one of the ruling military tribunal LeChamp—a Lower Feldan general L’Echeau—site of an orphanage planned by Leopold the Gambling Priest Leonardi—Duke of New Sepra Leopold—(1) organizer of a legendary Festival of Groundhogs; (2) the Gambling Priest, a onetime Diamondback opponent of Cerebus’s Lesla—an Iestan military commander Liebschaft—a T’gitan warrior beaten up by Cerebus Lion of Serrea, the—the head of the Western Church Liquid Amber Frostonite—a substance that can rob the Cockroach of his Sacred Wars Powers; aka whiskey Lishyan—a region south of Khaidoge the Ungovernable and east of Onliu little fellow with the hair—a mysterious artist and collector; later the middle head in the Trinity of Tarim little people—magical woods-dwelling creatures Living Tarim, the—an honorific given to the sitting Pope of the Eastern Church Living Tarim, the Sacred Amulet of—a golden talisman of power Lohi—brother of the Black Sun cult Lower Felda—a member state of the United Feldwar States Lower Iest—an electoral district of Iest Lucio—a New Sepran who runs a freak show that exhibited an unconscious Cerebus Luis—(1) a New Sepran sergeant, killed at Madame DuFort’s school; (2) a Diamondback player who cheats Luzzo—ranking diplomatic representative of Eshnosopur in Iest; a priest of the Inquisition Lyran the Death-Dealer—sergeant-at-arms of the Black Sun cult Mac Mufin, Bran—aka Star of the East, leader of the Pigts; believes Cerebus is a reincarnated Pigtian god-king Mac Mury, Fret—a Pigtian warrior, left in charge of the Pigts in Bran Mak Mufin’s absence Macs Milc, Cromag—a Pigt warrior Madame DuFort—an elderly woman who runs a small private school for girls; later revealed to be Professor Charles X. Claremont Maggie’s—a house of ill repute in Iest Magic—young Jaka’s name for Thunder, a wooden horse Maines—a Lower Feldan outpost Mann—borough of Lower Iest Match Diamondback—an Onliu variant of Diamondback McGrew, Dirty Drew—an Onliu miscreant, Dirty Fleagle’s baby brother McGrew, Dirty Fleagle—another Onliu miscreant; a third grade drop-out Mealc—island in the Sea of the South Mealca—city on the island of Mealc Meirgen of Ress—a Tcapmin, one of Sophia’s partners in crime Melvinbone—Elrod’s kingdom Mercer—borough of Lower Iest Merchants’ Association—an influential group in Iest Michelle—a former girlfriend of Cerebus’s; also the first name of Countess Detin Minon—a general; member of the military tribunal ruling Lower Felda M’isly—a god who runs afoul of Cerebus on Demonhorn mountain Missy—young Jaka’s rag doll and only friend Mit—a short, put-upon priest of the Black Sun cult Moon Roach, the—a persona of the Cockroach manufactured by Astoria Moreau—a legendary Diamondback player compared unfavorably to Cerebus Moresh—a city just south of the Wall of Tsi on the Feld River Most Holy—the title assumed by Cerebus upon his ascension to the papacy Mox, Bill—agent for the artist Set Mystic Cataclysm—an event caused by the Pigts that ended the Black Tower Empire and wiped out most known magic Nama Lotus—a flower with hallucinogenic perfume Nameless God of the Black Sun—an aardvark-like god of the Black Sun cult Necross the Mad—an evil sorcerer who became Thrunk, a giant stone creature Nell—a possibly fictitious niece of Cerebus’s Nelson—a town in Iest and Suburbs New Iest—former name of New Sepra New Matriarchy, the—(1) a Cirinist tract; (2) a name for Cirinist government New Republicanism—one of the platforms on which Cerebus runs in the Election of 1413 Newcastle—a manor in Oldcastle district of Iest N’Gamin—a region in the Tcapmin kingdoms known for merchants Nice girls don’t explode—Tarim’s comment on the creation of the universe No Man’s Sea—a northern body of water Northbell—an electoral district of Iest Nurse—a northern woman, Jaka’s nanny when Jaka was a young girl Nymin Barbarians, the—dwellers in the far southeast Ocean Sea—bounds the western part of the continent Oesis River—a river flowing through Iest, joining with the Feld River Old Boreala—brand used on either whiskey or wine Old Goober Bear and the Nasty Hedgehogs—a children’s fable Oldcastle—an electoral district of Iest, the lord of which was a member of Cerebus’s cabinet Oldham—a town in Iest and Suburbs On Governing—Cerebus’s book Onliu—(1) a region in the southeast known for bloodthirsty soldiers; (2) someone from Onliu Oscar—a friend of Rick’s Osiris River—a westerly river in Estarcion Oval Hideout—President Weisshaupt’s office in Beduin Pageant—a thrice-yearly Palnan entertainment begun by Lord Julius Palnu—a southern city-state ruled by Lord Julius Panrovy—a southern city-state known for dueling instructors and mad monks Parania—a region in Boreala Parmoc—a satellite city of Palnu Parmoor—electoral district of Iest Pedron—a town in Lower Felda Pentazin—an exotic poison from Eshnosopur Perce—an inner-circle Cirinist, thought by Cerebus to be a prostitute Peren—a Lower Feldan outpost Petuniacon—a political conference at the Festival of Petunias Petunias, Festival of—a festival commissioned by Lord Julius Pieureu—a Lower Feldan farmer Pigts—a fierce warrior race, small in number, living below the Red Marches; they speak a Conniptin dialect Pimina Mountains—mountain range east of Khaidoge the Ungovernable Pinshaw—Iestan military commander of the attacks on Lower Felda and New Sepra Pits, the—a city, buried by an earthquake, over which modern-day Palnu is built Po, Suenteus—a common name in Estarcion, given to (1) a ruthless conqueror; (2) Alfred, son of the conqueror; (3) the founder of Illusionism;

(4) Cerebus’s speech writer, an

Anarcho-Romantic

porridge—boiled ground oat grain, eaten for breakfast Port Seprania—a town in Seprania province of New Sepra Posey—an archbishop of the Western Church, later Cerebus’s aide potato—if raw and unpeeled, one of Cerebus’s favorite snacks Powers—a bishop of the Western Church, representative in Iest of the Lion of Serrea Preston, Sergeant—one of the Moon Roach’s personas Priestess—a more fashionable card game than Diamondback Prime Minister and the Hussey, The—an anti-Cirinist romance novel penned by Cerebus under Weisshaupt’s guidance Pud—see Withers, Pud Punch the Wonder Bunny—a name Elrod uses to refer to Cerebus Quarbet—a \frac{1}{4} copper-bit ante in Karet Rachinne—a city in which Claremont once studied Ram and Peackock Tavern—a former tavern in Imesh Ram’s Lords Tavern—Cerebus’s favorite place in which to pick fights in Iest Rashazar—an Eshnosopurian prince, the “Lion of the Coast” Red Claw cult—another killer cult Red Crescent Quarter—an formerly unsavory section of Imesh Red Horn—a tavern in Togith Red Marches—a region of quicksand and black flies, home to the Pigts Red Sophia—daughter of the wizard Henrot; later Cerebus’s wife Redeemer Dynasty, the—an empire toppled by the Pigts redwoods—a race of creatures that ruled Earth until they exploded a Sodium Chloride Bomb Regency, the—the oldest and nicest hotel in Iest; constructed with Pigt slave labor Reisenshein—a fictitious genius who invented Alucase Repredigressionals—a New Republican splinter party Ress—a city east of the Tcapmin Kingdoms Rick—Jaka’s husband Riguini—a corporal in the New Sepran military Rivershire—an electoral district of Iest Roach-Arang—one of the Cockroach’s weapons Rose—one of Cerebus’s aunts Roth-Sump—one of the masters of an Apocalypse Beast Sacred Wars—aka Secret Sacred Wars; the Cockroach’s war against the False Pope, or possibly the Cockroach’s struggle to maintain control over Claremont Salice, Elwyn—a guard at Hobbsgate Prison San Tremain—a city in Lower Felda known for a recent avalanche attributed to Necross the Mad Sarte—head of Seaf’s military government Schwagerin—a T’gitan warrior Scorz—partner in Scorz, Scorz and Sons, a sewage engineering firm Sea of Tanchya—easternmost body of water Sea of the South—bounds the southern part of the continent, including the southern city-states, and Onliu, Lishyan, and Amstanat Seaf—a region north of Palnu Secret Sacred Wars—see Sacred Wars Sedra—mistress of K’Cor Seersucker—Elrod’s black blade, destroyed in battle with Cerebus Sepran Empire, the—large empire in the west, south of Boreala Sepran lamp oil payoff scandal—a scandal haunting the ruling military tribunal of Lower Felda Seret—former Executive Commander; member of the military tribunal ruling Lower Felda Serrea—important city in the Sepran Empire Set—(1) a deity, perhaps the same as the ancient Egyptians’s; (2) an artist who temporarily distracts Cerebus from the Final Ascension Seventh Sphere—an altered state of consciousness Seymour—one of Lord Julius’s uncles Shadow Crawler—a terrifying snake-like creature that dwells in the Pits of Palnu Shea—an Illusionist involved in the Togith street conflict between the Illusionists and Cirinists Shen Steppes—northernmost eastern region Shomber—prince of Ianag Silverspoon—Lord Julius’s wastrel son Six Crises, the—a document by Suenteus Po, Cerebus’s speechwriter Skullpulper—Drew McGrew’s war hammer skunks—ceremonially burned by plains dwellers Smina—a city in the far northeast with an important bank Sofim River—a river flowing through the central part of the Sepran Empire Sopai—Hsifan assassins Sophia—see Red Sophia Spleen, Hagbad—inventor of armour Sponge Boy—a friend of the Flaming Carrot Squinteye—a sailor Starym, Innec—a one-time Grandlord of Palnu; kidnaped from the Regency ca 1263 Stemson—an exporter in Fluroc, deceased Stories for Young Girls—a very old collection of instructional fables attributed to Cirin Storm’Send—a farm in Northbell district of Iest; the lord thereof Strangis—an aristocrat who hired G’evan to steal Cerebus from the Cirinists Stromm—Gudre’s son, revered as a T’gitan god of thunder; a tall, blond T’gitan warrior who lost his tongue to Krull Suburban Merchants’ Guild—an influential group in Iest and Suburbs succubus—a supernatural being with the power to enslave warriors’s souls Suenteus Po—see Po, Suenteus Sump Thing—another name for Filbert, Lord Roth-Sump’s Apocalypse Beast Sun, The—a partisan paper of Cerebus’s T’agath—borough of Lower Iest T’ana—electoral district of Iest Tanes—see Feras, Tanes Tansubal—a city in Boreala; home of Henrot; possibly a province of the Sepran Empire Tarim—male deity; the Void Tarim, the Chronicles of—a holy book of the Orthodox Tarimite Church tavern—an old-time bar/cafe Tavers, House of—Jaka’s and Julius’s family House Tcapmin Kingdoms—northeastern region Tchens—brother of the Black Sun cult Tebet—a ten copper-bit ante in Karet Temza—a port on the western coast of Boreala Tepin—an acquaintance of G’evan Terim—female deity; that which was once the non-Void Terms of Penance—a Tarimite article by which heretics may recant T’gitans—warlike tribes united by Gudre Theocrats—Cerebus’s support group in the Election of 1413 Theopheles—son of a farmer in Parmoc Theresa—an aide to Astoria, later an informant to Weisshaupt; a former student of DuFort/Claremont Theyr—a Lower Feldan farming community Thomas—a man with whom one of Cerebus’s old girlfriends ran off Throgo—a Tcapmin border guard Thrunk—a giant, animated, stone creature that temporarily dislodges Cerebus from the Papacy; formerly a magician named Necross the Mad Thugg the Unseemly—an eyepatch-wearing admirer of Sophia’s Thunder—a wooden horse on springs and hinges on the Tavers playground Tiberius—a New Sepran soldier, killed at Madame DuFort’s school Times, the—an Iestan newspaper allied with Lord Julius Togith—one of the city-states in the United Feldwar States Tom—a Lower Iestian street urchin toothpicks—the Wolveroach’s favorite smoking mixture Trebu—brother of the Black Sun cult Trinity of Tarim—Fred, Ethel, and the little artist fellow with the hair True History of the 1413 Election—a document by Suenteus Po, Cerebus’s speechwriter Trystrim—a friend of Lord Silverspoon’s Tucker—an Iestan bank official Turg the Unduly Tall—a lout who hangs out with E’lass Turl—a Conniptin captain Two-Step Diamondback—see Karet Tynsdale-Clyde, Mrs.—head of an anti-Cirinist Presidential Commission UFS—abbreviation for the United Feldwar States unicorn—the emblem of the House of Tavers United Feldwar States—a federation formed by Weisshaupt in 1414, consisting of Lower Felda, Iest, New Sepra, Togith, and Palnu unorthodox economic revenge—The Moon Roach’s motto Unshib—a city in the Tcapmin Kingdoms Upper Felda—a Cirinist state Vanaheim—heaven Vant—an archbishop, a Cirinist sympathizer and spy vapours—in northern health practice, mists produced by heated water; thought to be dangerous Venusians—extraterrestrials whose arrival K’Cor fears Vessariat—a satellite city of Palnu Vichy—full name Vichy Dupris, an unwanted suitor of Michelle Detin’s; an Onliu Anarco-Romantic Vinter’s Gate—a Cirinist convent Von Croyder—an old Iestan adversary of Cerebus’s Wall of Tsi—a great wall south of the Tcapmin Kingdoms running from the Feld River almost to the Pimina Mountains, built to keep out Khaidoge the Ungovernable Warmtho—a special anti-Frostonite sheilding Weisshaupt—President of the UFS; formerly an official in Lower Felda Wenda—an inner-circle Cirinist priestess whiskey—another favorite beverage of Cerebus’s wickets—croquet, a favorite game of Elf’s Wineman Hotel—a hotel in Port Seprania where Cerebus finds himself in bed with Sophia Withers, Pud—a widower, owner of a tavern and grocery on a slope of the Iestan tower Wolf’s Gate—an area of Togith Wolveroach—a persona of the Cockroach manufactured by Weisshaupt Woman Thing—Claremont’s Apocalypse Beast; later, Ethel wuffa wuffa—the sound snowshoes make Wyndmel-Smith, Lucius and Aristonnia—wealthy citizens of Palnu X-Persons—a group of AUFO agents hand-picked by Weisshaupt yak—one of Cerebus’s favorite meat animals Yan-Su—a brass disc used by Sopai assassins as a throwing weapon You-Know-Who—nom de grafito of Elf Zerbutz—T’gitan pronunciation of Cerebus’s name Z’hin Z’homlie—one by whose beard wizardly oaths are sworn

15 Index

aardvark others, 85, 93 Pigtian statue of, 26 uses in magic, 30

Aardvark-Vanaheim

contact information, 103 Aardvarkian Empire, 55 Abbess, the, 46, 55, 67, 83 Abhorring Vacuums, 95 accents, 3

Northern, 4 Accurately Inexplicable, 92 acknowledgements, 104 Acquired Tastes, 84 Adams, Cory, 15 Adams, Neal, 15 Add One Mummified Bat, 16 address of publisher, 103 Agerden, Duke of, 34 Ages of Cerebus, The, 15 albatross, 55, 90

recovery of, 60 significance of, 58

All Lined Up, 78 All the Suns are Daughters, 95 Alliance, 50 Amazing Heroes, 15 Ambassador Suite, 47, 50 Anarcho-Romantics, 40

manifesto from, 58

riot of, 58 Anchor That’s Going Places, An, 91 Anderson, Murphy, 12 Animated Cerebus, the, 16 Another Thing Coming, 83 Anything Done for the First Time Un-

leashes a Demon, 82 Anything Goes!, 15 Apocalypse Beast, 39, 86 Applicant, The, 12 Appointments, 80 Approved By, 79 apricot brandy, 25 Aragonés, Sergio, 15 armour, 12 Arnold the Isshurian, 47 Artemis, 42, 50, 64, 76 artist/collector chap, 39, 46, 53, 71, 86,

94 Astoria, 80

alliance with Cerebus, 50 argument with Abbess, 83 background of, 80 description of, 44, 64 drunk with Moon Roach, 56 first appearance of, 50 imprisonment of, 90 magical disappearance of, 89 rape of, 91 revealed as Julius’s ex-wife, 53 trial of, 92 with Sir Gerrik, 36 At the Club One Afternoon, 79 Audacious Tenacity, Tenacious Audac-

ity, 90 audit, 49 Austin, Terry, 12 AV in 3-D, 11 Ave Avid, 93

back issues, 9 backup features, 9 Balances, 58 Baldy Pate, 47 Base Tranquillity, 95 Baskin, 32, 46 Batman, 88 Bear, 12, 82

description of, 68

quitting employment, 85 Beasties, Ye Book of, 30 Beatty, Terry, 11 Becoming Synonymous with Some-

thing Indescribable, 86 Beduin by Night, 29 Beduin, 29, 36, 37, 38 Beguiling, The, 38 bi-weekly edition, 8, 10 Big Bang, 95 Black Blossom Lotus, 28, 29 Black Magiking, 30 Black Sun Rising, 27 Black Sun, Temple of, 27 Blakely, Harmon, 68

delivering dispatches, 59 description of, 46 first appearance of, 54

Bodē, Mark, 13 Boobah, 82

description of, 68 Boom, 80 Boreala

location of, 3 marauders, 24 press gangs, 24

Breaking Up Is Hard To Do, 15 Bucky, 37 Bug, 21 Bunky, 37 bunny sketches, 53 bunny suit, 19 Burden, Bob, 11, 13 Bureaucratic Rebellion, The, 57 Buz, 28 Campaign’s End, 55 Campaign, 54 cannons, 80, 88, 89 Captain America, 37 Captain Cockroach, 37 Captive in Boreala, 24 Carroll E. King Reads, 79 Causality Casualty, 85 Cerberus, 7 Cerebus comic artwork, 8 authors of, 6 availability of, 8 bi-weekly edition, 8, 10 continuing, 97 date of introduction, 6 early stories, 17 out-of-print material, 9 phone books, 9 radio show, 17 Cerebus Dreams II, 11 Cerebus Dreams, 10 Cerebus Jam, 12 Cerebus versus the Spirit, 12 Cerebus Was, 80 Cerebus

appearance of white light over, 83 appointment as Kitchen Staff Su-

pervisor, 32 attempted suicide of, 97 candidacy for Prime Minister, 53 conquest of Red Marches, 84 demanding gold for Tarim, 82 description of, 4 desire for money, 51, 76 elevation to papacy, 82 fall from papacy, 86 future of, 95, 96 nomination to papacy, 81 puzzle posters of, 35, 81 re-elevation as Prime Minister, 78 resignation as Prime Minister, 60 takes honorific Most Holy, 82 threat to destroy world, 83 writing a book, 76

Champion, 33 Chariot of the Queen, Chariot of the

Lovers, 87 Chasing Cootie, 50 Church & State

cast of, 64 price of phone book, 104 section, 62 summary of, 76

Cirin appearance of, 93 description of, 71 relation to Astoria, 80

Cirinism

description of, 63 Cirinists, 35, 38, 41, 48 city-states

location of, 3 Claremont, Chris, 22 Claremont, Professor Charles X., 39,

64, 86, 88 description of, 22 first appearance of, 39 reappearance of, 78

Cockney accents, 3 Cockroach, the, 29, 37, 64

Artemis, 50, 76 Captain Cockroach, 37 description of, 19, 42 first appearance of, 29 grab for papacy, 88 jealousy of Cerebus, 51 Kevitch, 50 Moon Roach, 50 Priest Roach, 61 Secret Sacred Wars, 86 Sergeant Preston, 56, 58 use in UFS, 77 with Bunky, 37, 88 Wolveroach, 76

codeine, 88 Collins, Max, 11 Comic Informer, The, 15 Comics Journal, The, 15 Conan the Barbarian, 18, 24 Conniptins

first appearance of, 27 invading Iest, 60

conspiracies, 41 Cooke, Alistair, 3 Cootie, 21 Countess & the Aardvark, The, 76 coup, preparation for, 56 Couple Capable, 94 Covenant, Thomas, 4 creation of the universe, 95 cross-overs, 13

Dalkey, Kara, 17 Dark Horse Comics, 13 Dark Knight Returns, The, 88 date, relative to present, 1 Day in the Pits, A, 32 Day of Greatness, Age of Consent, 83 Day of the Earth-Pig, 27 Deadman, 19 Death of Elrod, the, 37 Death’s Dark Tread, 25 Debts, 49 Deciding Vote, The, 56 Defense of Fort Columbia, The, 12 Demonhorn, 10 Denieau, 31 Denoument, 60 desert island, 31 Despuess, 31 Detin, Countess Michelle, 76, 87

description of, 64

with Weisshaupt, 77, 87 Diamondback, 10, 48 Distant Soil, A, 12 doomsday weapon, 96 Doran, 16 Doran, Colleen, 12 Drew, 42 duck statue, 49, 55 DuFort, Madame, 38 Dunham, 46

E’lass, 26, 32 Earnest Nonsense, 94 Earth Will Shake, The, 103 Eastern Church, 40 Eastman, Kevin, 13 echos, 92 Eighth Sphere, 48 Eisner, Will, 12 Election Night, 56 Elf, 68, 80

comments on Cerebus’s book, 79 description of, 42 first appearance of, 49 mad at Cerebus, 51

Elfguest, 13 Elfquest, 13 Elric of Melnibone, 19 Elrod, 27, 30, 37, 39, 88

delivering invitation, 52 description of, 19, 46 first appearance of, 25 temporary death of, 37 Endgame, 92 energy globes, 28 English accents, 3 Epic Illustrated, 14 Estarcion

description of, 1 history of, 95 technology of, 3

Ethel, 71, 86, 94

exile, 60 Exodus, 61 Extrusion, Intrusion, 95 Eye in the Pyramid, 32, 39, 48 Eye in the Pyramid, The, 103 Eye of Terim, 24

F.O.O.G., 15

Fafhrd and the Grey Mouser, 7

Famous the Aardvark, 41 Fantagraphics Books, 15 Fascination & Fruition, 84 Feiffer, Jules, 94 Feld River, 30 Feras, Tanes, 25, 28 Filbert, 39 Filgate, Euling S.

description of, 46 first appearance of, 53 meeting with Mrs. Henrot-Gutch,

79 Final Ascension, The, 11, 93 first mention of, 86 history of, 90 First Fifth, the, 17 First Impression, 79 First Invention of Armour: 1404, The,

12 First Publications, 13 Five, 89 Flame Jewel, The, 22 Flaming Carrot, 11, 13, 94 Fleagle, 42 floating floor in space, 30, 86, 90 Fluroc, 33, 34, 37 Flying Off the Handle at Oblique An-

gles, 88 fnord, 102 Foghorn Leghorn, 19 fortune teller, 34 freak show, 36 Fred, 71, 86, 94 French accents, 3 Friction, 51 Friendly Reminder, A, 14

G’evan, 36 Geet-a, 34 gerbies, 35 Gerhard, 6, 14 Gerrik, 36 Girl Next Door, The, 14 Goat, 55

candidate for Prime Minister, 55 Goodabbey, 55 Gorce, 34, 36, 37 Grace District, 55 graffito, magical, 51 Grandlord of Palnu, 32 Graus, 34 Greater Iest, 40

accents, 3 tower of, 62

Greatness Among Flies, 94 Greely, Ned, 41, 49 Groo the Wanderer, 15 Gudre, 33, 36 Guild of Government Employees, 57 Gwain, 76 Hadden, Holland M., 41, 49, 50

death of, 50 Hampton, Bo, 12 Hampton, Scott, 12 hangover cures, 10 Harbourside, 55 Hargraves, 47, 59 Harmony IV, 68, 78 Harpengate, 36 Hedges, 83 Henrot, 35, 47, 57, 71, 86 description of, 18 first appearance of, 24

Henrot-Gutch, 79 description of, 68 introduction to Cerebus, 78 search for Blakely, 82

Heroes, 55

High Holy Days, 33

High Society cast of, 41 price of phone book, 104 section, 40 summary of, 47

Hill-Hodges, 42 His First Fifth, 14 His First Sword, 16 Hortne, 34 Hovering Below the Fray, 83 Howard, Robert E., 18, 19 Hsifan Khanate

location of, 3 Hsifans, 37

Idol, The, 25

Iest, 40 economy of, 40 growth of tower, 88 invasion by Cirin, 96 invasion by Conniptins, 60 invasion by Hsifans, 59 location of, 3 weather in, 3

Ignore It, It’s Just Another Reality, 83 Illuminatus! Trilogy, The, 102 Illusionism, 35 Illusionists, 41, 48 Imesh, 28 inauguration jewels, 57 Inquisition, 40, 50, 51 Insecure Sinecure, The, 76 issue numbers, 9 It’s Showtime, 53 Italian accents, 3

Jacobi, Lou, 94 Jagger, Mick, 88 Jaka’s Story, 97

cast of, 98 summary of, 100

Jaka, 46, 52, 71, 84 audience with Most Holy, 84 description of, 19, 98 first appearance of, 26 letter from, 33, 79 returning Cerebus’s sword, 52 trying to get Cerebus to leave, 59

Janette, 38 Jockley, Blake, 42 Journey, 11 Judge, The, 71, 94 Julius, 32, 39, 64

bathing in Cerebus’s rooms, 54 description of, 22, 42 disguised as waiter, 52 first appearance of, 32 nomination to Presidency, 84

K’Cor, 46

first appearance of, 28 Kael, B’ren, 36 Karet, 10 Katrina, 38, 39 Keef, 88 Kevillism

description of, 63 Kevillists, 41, 48 Kevitch, 42, 50 Kidnapping of an Aardvark, The, 48 Kitchen Staff Supervisor

Cerebus’s appointment to, 32 Koghem, 28 Krull, 33

Laird, Peter, 13 Last Stand, The, 59 Leiber, Fritz, 7 Leonardi, 47, 57 Leopold, 10 Lion of Serrea, 62, 78

assassination of, 90 loans to city-states, 40 Loubert, Denise, 7, 9 Loubert, Michael, 7 Lower Felda, 29, 30, 31

division of, 57

invasion by Cerebus, 57 military tribunal of, 37

Lower Iest, 40

accents, 3 Luis, 38 Luth, Tom, 15 Luzzo

description of, 44

M’isly, 10 magical creatures, 3 Magicking, 30 mail delivery times, 104 Mak Morn, Bran, 19 Mak Mufin, Bran, 52, 71

change in spelling, 46 description of, 19, 46 first appearance of, 25 reappearance of, 83 sent for Pigtian armies, 58 suicide of, 86

Marvel, 14 Marx, Chico, 3, 47 Marx, Groucho, 22 Mass With Substance, 93 McGrew brothers, 49, 71, 86

description of, 42 first appearance of, 48

Meirgen, 28 Memoirs, 79 Merchant Cockroach, the, 29 Merchant of Unshib, 28 Messner-Loebs, Bill, 11 Miami Mice, 13 Michelle, 15, 76 mildew, 31 Mind Game II, 48 Mind Game III, 82 Mind Game IV, 81 Mind Game, 35 Minnecon, 17 Missing It for the World, 88 Mit, 27 Moon Roach, the, 42, 50 moon

exploration of, 96 moonshine, 56 Moorcock, Michael, 19 Morning After, The, 10 Mrs. Tynsdale-Clyde’s Tea, 80 Ms. Tree, 11 Mystery Achievement, 97 Mystic We, The, 78

N’Gamin merchants, 34

Nameless God, 27 Necross the Mad, 21

first appearance of, 31 Neil the Horse, 11 Never Pray for Change, 82 New Republicanism, 54 New Sepra

accents, 3

devaluation of currency, 58 invasion by Cerebus, 58 location of, 3 soldiers, 38

Night at the Masque, A, 32

Night Before, The, 52 Night In Iest, A, 57 Night on the Town, A, 11 normalman, 11 Note, 79 Nurse

description of, 98

Odd Transformation 4, 91

Odd Transformations No. 3: Dead

Friends, 89 Odd Transformations Part Two, 85 Odd Transformations, 85 Oldcastle, 55

description of, 46 Omaha, The Cat Dancer, 8 Onliu, 31, 33, 42 accents, 3

Origin of the Wolveroach, The, 76 Out With the In Crowd, 89 outtakes

summary of, 10 owl, 34 Palnu, 33, 34, 37, 40

diplomatic representative of, 40 location of, 3 pits of, 32 ruler of, 32 T’gitan invasion of, 37

panel discussions, 53 Pate, Baldy, 57 Perce, 34, 35 Petuniacon Day Three, 54 Petuniacon Day Two, 53 Petuniacon, 52 pheasant hunt, 49 phone book collections, 9 Picts, 19 Pigts, 19

first appearance of, 25 Pinshaw, 47, 57 plague, 34, 38 Po, Suenteus

Anarcho-Romantic, 58, 60 description of, 42, 46 in Eighth Sphere, 48 in Seventh Sphere, 35 Weisshaupt’s uncle, 85

Poet, The, 97 Pogrom’s Progress, 97 Popeye, 12 portfolios, 16 Posey

description of, 68 Powers, 80

confrontation with Weisshaupt, 79 description of, 68

Preston, Sergeant, 42

attacks on dissidents, 58 Priest Roach, 61 Prime Minister of Iest, 49

election for, 53

meeting with Cerebus, 52 Prince Valiant, 31, 76 Prisoner, The, 90 Prologue, 97 publisher, address of, 103

Red Claw cultists, 30, 57 Red Marches

location of, 3 Red Sonja, 18 Red Sophia, 34, 67

Black Blossom Lotus affair, 28 confrontation with Cerebus, 89 description of, 18 divorce from Cerebus, 91 first appearance of, 24 marriage to Cerebus, 78 tiff with Most Holy, 83, 84

Regency Hotel, 47 religious establishment, 40, 62 Repercussions, 49 Richards, Keith, 88 Rick

description of, 98 Rip Off Press, 13 Rolling Stones, 88 Roth-Sump, 39 Rough Pope, 79

Saba, Arn, 11 Sacred Wars, 86 Sane As It Ever Was, 83 Savage Tales, 24 Secret Sacred Wars Roach, 64 Secret Sacred Wars, 86 secret societies, 41 Secret, The, 26 Selling Insurance, 14 Sepran Empire, 40

location of, 3 Serrea, 39 Seventh Sphere, 35 Shadow Crawler, 32 She-Devil in the Shadows, 34 Shea, Robert, 102 Shockwave, 17 Silverspoon, 31, 76 Sim, Dave, 6

commentaries by, 10 interviews of, 15

Single Page, The, 9 Six Deadly Sins, The, 17 sketches by candidates, 53 Skorz, 41, 81 So, 91 Something Cerebus Was, 80 Something’s different, 93 Song of Red Sophia, 24 Sopai assassin, 37 Sophia, 79 Southwestern accents, 3 spider-beast, 27 Spinning Straw into Gold, 85 Spirit, the, 12 Spleen, Hagbard, 12 Sponge Boy, 94 Square One, 96 Squinteye the Sailor, 12 Storm’Send, 55, 56, 85

description of, 46 Stormy Weather, 80 Strangis, 36 Stromm, 33 Strong, Artemis, 42, 50 Suburban Merchants Guild, 55 succubus, 24 Suddenly, Sophia, 78 Summit Enchanted Evening, 57 Sump-Thing, 39, 86 Sun, The, 55 Swamp Sounds, 38 Swords Against Imesh, 28 Swords Against Wizardry, 7 Swords of Cerebus, 10 Swords of Lankhmar, The, 7

T’gitans, 33, 34, 36, 37 Talking Heads, 86 Talking to Tarim, 90 tank, 30 Tarim

Church of, 41, 62 priests of, 31 Trinity of, 87

Tarot, 55, 63

Tcapmin Kingdoms location of, 3

technology, 3

Teenage Mutant Ninja Turtles, 13 telephone number of publisher, 104 Terrible Analogies, 84 Theresa, 22, 38, 46, 68, 80 This Woman, This Thing, 10, 39 Thomas, Roy, 18, 24 Thorne, Frank, 18 Three Days Before, 52 Three-D, 11 Thrill of Agony and Victory of Defeat,

The, 82 Thrunk

description of, 71 destruction of, 89 first appearance of, 31 usurpation of papacy, 86

Tiberius, 38

Time Passes and Life Manoeuvres, 84 Times, The, 55 Todd, Larry, 13 Togith, 34 Touch Not the Priestess, 87 tower

growth of, 88, 90, 93

launch of, 93 Towers Analogous, 88 Tree Planting, 79 Trial, The, 92 Trinity of Tarim, 71 True North, 15 Trystrim, 76 Turg the Unduly-Tall, 26, 32 Two Days Before, 52 Tynsdale-Clyde, Mrs., 79, 80

Unique Story, A, 9 United Feldwar States

description of, 62 formation of, 77

universe, creation of, 95 Unknown Given, The, 91 Upper Felda, 38, 80 Upstairs Downstairs, 59

Valentino, 11 Varying Reasons of Assorted Depths,

85 Venusians, 28 Vichy, 64, 76

Walking on the Moon, 94 Walls of Palnu, The, 32 war bonds, 37 WaRP Graphics, 13 Weisshaupt, 37

collapse of, 83

confrontation with Cerebus, 83 description of, 22, 64 first appearance of, 37 gift to Cerebus, 87, 89, 93 historical references to, 103 on death bed, 85 possible motivation, 78 reappearance of, 77 stone sculptures of, 94 Well-Equipped Bar, A, 16 Wenda, 35, 36 Western Church, 40

take-over of Iest, 60 What Happened Between Issues

Twenty and Twenty-One, 36 Why & the Are, The, 77 wickets, 49 Widow’s Son, The, 103 Wilson, Robert Anton, 102 Windsor-Smith, Barry, 11, 18, 19, 24 Withers, Pud

description of, 98 Wolveroach, 64, 76 Woman-Thing, 39, 85, 86 world, description of, 1 Wrightson, Berni, 36 Writing, 76

X-Men, 22

Xavier, Charles, 22

Yosemite Sam, 3 Young Cerebus, 14

Zone, Ray, 11